A Complete Bibliography of Publications in the International Journal on Software Tools for

Technology Transfer (STTT)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

31 May 2024 Version 1.41

Title word cross-reference

∗ [62, 42]. ^{+} [828, 803]. ^{⋆} [751]. α [719]. ∫ [695]. k [802, 637]. LTL(F) [501]. μ [406, 407]. ω [471]. ϕ [231]. π [149]. ∥ [636].

-automata [471]. -calculus [406, 407, 149]. -induction [802, 637].

1394 [135, 18]. 1394a [79]. 17th [578]. 1998 [29]. 19th [819].

2 [305]. 2.0 [213, 793]. 2000 [48]. 2003 [253]. 2008 [377]. 2011 [470]. 2012 [583, 584, 530]. 2014 [715, 586]. 2015 [673]. 2017 [725, 882]. 2018 [763, 762, 754, 748].

2019 [781, 848, 810, 850, 854]. 2020 [886]. 2021 [911, 907]. 2nd [940].

3 [764, 299, 762, 765, 766, 307, 769, 306, 513, 512, 768, 302, 298, 309, 303, 308, 514, 515].

6 [846]. 6.1 [842]. 6.9 [879]. 653 [830].

= [7].

AADL [897, 761]. AbC [776]. abduction [661]. abort [422]. abort-aware [422]. ABS [455]. Abstract [435, 620, 856, 132, 419, 481, 425, 131, 122, 551]. Abstraction [597, 182, 489, 621, 323, 382, 123, 429, 359, 205, 326, 234, 743, 407, 385, 930, 325, 510, 40, 867, 426]. abstraction-based [326, 325].

Abstraction-guided [489]. abstraction-refinement [407]. abstraction/refinement [359]. abstractions [406, 217, 618, 664, 751, 321, 951, 947]. ABZ [763, 762, 642]. ABZ-2018 [763]. acceleration [310, 436, 888, 500]. accelerators [199]. access [778, 273, 906, 317]. accompanying [327]. accuracy [655]. accurate [880]. ACL2 [48]. action [77, 279, 705, 528]. action-based [705]. active [531, 361, 670, 938, 534]. activity [522, 472]. actors [704, 785]. actuator [704]. ad [721, 582]. ad-hoc [582]. adaptable [455]. adaptation [948]. Adapters [254]. adaptive [772, 771, 943, 775, 901, 606, 951, 940]. address [63]. ADTLang [800]. advanced [481]. Advancements [152]. Advances [684, 659, 605, 598, 667, 325, 680, 445, 183, 298]. adversaries [834, 835]. advisories [671]. aerospace [357]. affine [925]. against [608, 770]. AGATHA [129]. agent [293, 163, 633, 945]. agent-based [163]. agents [872, 946]. AGG [374]. aggregation [920, 29]. AHB [497]. AI [502]. aided [142]. air [860]. airborne [671, 604]. Aircraft [640]. airplane [544]. AJAX [315]. ALDÉBARAN [12]. algebra [336, 859, 246, 912]. Algebraic [920, 901, 937]. Algebraically [919]. Algorithm [599, 917, 952, 759, 141, 176, 202, 142, 73, 490]. Algorithmic [488, 297]. algorithmics [165]. algorithms [667, 835, 834, 886, 322, 588, 33, 330, 154, 901, 194, 214, 911, 882, 599, 108, 855, 449, 854, 150]. alias [466]. alias-based [466]. alive [181]. ALL-TIMES [463]. allocation [73]. along [249]. Alternating [446]. alternation [204]. alternation-free [204]. AMBA [497]. among [252]. amplification [873]. AMT [793]. AMulet2 [912]. analog [479]. analyser [347]. analyses [225]. Analysing [575, 26, 23, 24, 262, 73, 455]. Analysis [244, 894, 33, 924, 281, 598, 141, 15, 667, 380, 738, 889, 172, 789, 886, 442, 507, 91, 408, 241, 126, 16, 164, 31, 963, 862, 579, 114, 537, 799, 721, 470, 74, 427, 461, 287, 733, 810, 14, 658, 958, 930, 325, 840, 522, 214, 911, 484, 902, 481, 477, 897, 882, 694, 291, 393, 70, 410, 467, 900, 768, 680, 468, 464, 261, 504, 793, 910, 459, 344, 389, 231, 624, 560, 475, 528, 133, 220, 627, 632, 564, 854, 523, 741]. analysis-friendly [738]. Analytic [918, 474]. analyzer [17]. Analyzing [531, 928, 658, 22, 904, 704, 913]. Android [674, 557]. animate [156]. annotated [686]. anomalies [873]. antecedent [250]. anti [301]. anti-product [301]. ants [944]. AOP [672]. AOP-specific [672]. API [776]. APIs [602]. applicability [956]. Application [336, 63, 416, 61, 957, 557, 565, 430, 396, 552, 448, 820, 454, 95, 954, 665, 514, 52]. Application-controlled [63]. applications [759, 355, 700, 188, 316, 356, 774, 786, 411, 674, 484, 315, 351, 41, 57, 312, 801, 308, 398, 724, 576]. applied [189, 197, 315, 449]. Applying [746, 530, 142, 467, 240]. approach [772, 382, 406, 355, 189, 161, 356, 163, 476, 437, 730, 539, 769, 872, 288, 800, 454, 496, 155, 550, 129, 351, 128, 761, 945, 692, 302, 731, 354, 814, 426, 711, 869, 294, 946]. Approaches [537, 238, 824, 699, 735, 563, 910, 240, 409, 640, 269]. Approximate [883, 779, 572]. approximated [426]. Approximating [343]. apps [755]. APSET [557]. arc [703]. Architecture [465, 873, 774, 68, 162, 22]. architectures [26, 648]. ARINC [830]. ARINC-653 [830]. arithmetic [203, 56, 326, 816, 50, 491, 613, 817]. arms [334]. array [895]. art [450]. ASAP [736]. ASM [645]. Aspect [428, 254, 832, 340]. Aspect-oriented [254]. aspects [119]. aSPIN [132]. assembly [148]. assertions [228]. Assessing [699, 545]. assessment [461, 372, 798, 378, 540, 719, 801, 869]. assisted [141]. Assume [887]. Assumption [175]. Assumption-based [175]. assurance [775, 200, 748]. asynchronous [935, 459, 328]. asynchronously [657]. ATL [709]. atomicity [392]. attack [800, 527]. attitudes [99]. attribute [877]. attributed [712]. Authorized [501]. Authorship [736]. auto [670]. auto-active [670]. Automata [803, 723, 717, 746, 406, 531, 172, 217, 729, 808, 779, 884, 745, 684, 477, 905, 124, 342, 668, 423, 85, 471, 151, 593]. Automata-based [803]. Automated [611, 521, 330, 810, 542, 317, 900, 399, 124, 242, 216, 712, 686, 710, 598, 332, 650, 506, 842, 291, 397, 680, 373, 722, 502, 516, 898, 121, 693]. Automatic [404, 833, 747, 135, 251, 613, 27, 304, 565, 91, 316, 655, 313, 129, 896, 672, 169]. automatically [500]. Automating [396, 300, 791]. Automation [647, 49, 300, 242, 298]. automaton [34]. automotive [956]. autonomically [778]. autonomous [941, 872]. autopilot [695]. AutoProof [670, 589]. availability [200]. Average [423]. Average-price-per-reward [423]. avionic [963]. avionics [864]. avoidance [519, 671, 871, 604]. Avoiding [813]. aware [422, 623].

B [387, 763, 238, 336, 643, 156, 766, 721, 767, 291, 641, 358, 360, 634, 903, 640, 636]. backjumping [505]. backtracking [732]. bad [896]. Bandera [84]. based [915, 238, 802, 517, 239, 602, 893, 217, 773, 949, 529, 812, 618, 811, 734, 205, 273, 175, 326, 179, 335, 234, 416, 16, 164, 505, 163, 65, 246, 664, 751, 307, 396, 963, 650, 608, 536, 533, 580, 813, 230, 251, 750, 749, 461, 411, 524, 877, 830, 554, 648, 646, 739, 678, 607, 698, 147, 325, 417, 452, 656, 200, 803, 107, 496, 279, 597, 617, 474, 477, 11, 905, 550, 75, 599, 145, 768, 315, 705, 761, 376, 353, 534, 731, 755, 722, 445, 183, 130, 727, 240, 748, 557, 354, 814, 466, 548, 801]. based [627, 493, 947, 710, 785, 328, 638, 724, 609, 257, 346, 660]. basin [701]. Bayesian [404, 824, 341]. BDD [529, 60, 151]. BDD-based [529]. BDD-like [151]. BDDs [61, 203, 181, 57]. be [42]. Behavior [15, 147, 531, 928]. Behavior-based [147]. behavioral [416, 348, 551]. behaviour [487, 262]. behavioural [878]. Benchmark [307, 851, 526]. Benchmarking [38, 714, 885]. benchmarks [715, 532]. better [287]. between [478]. beyond [926]. biased [430]. bidirectional [353]. Binary [55, 863, 858, 56]. binding [399]. bioinformatics [626]. biological [566, 576]. birds [944]. Bisimulation [66, 176, 683]. bit [326, 509]. bit-vector [326]. bitcode [876]. Blast [276]. blind [274]. block [870]. blocks [628, 481]. blueprint [592]. BMC [179, 505, 178, 379]. BMC-based [179]. BMC’03 [177]. Boolean [123, 204, 61, 890, 120]. boosting [404]. bottom [944]. bound [616]. Bounded [324, 507, 494, 230, 905, 565, 203, 408, 184, 660, 637, 842, 530, 855, 328]. bounds [217]. brainiac [857]. branch [770]. branches [843]. bridging [679]. Bringing [861]. broadcast [620]. brute [527]. brute-force [527]. Büchi [884, 684, 668, 85, 471]. bugs [420]. building [359, 628, 80, 481, 242, 171]. bus [18]. business [602, 356, 862, 284, 472]. bytecode [630, 931]. bytecode-level [931].

C [727, 802, 123, 190, 332, 414, 637, 739, 550, 333]. CADP [470]. CAESAR SOLVE [204]. CakeML [914]. Calculus [790, 406, 407, 888, 149]. call [334]. caller [562]. caller-side [562]. can [444, 822]. Capability [263]. capturing [457]. CARA [141, 145, 140, 144, 143]. card [759, 130]. care [359]. Cartesian [123]. Case [236, 763, 565, 738, 639, 762, 338, 118, 963, 702, 861, 114, 536, 842, 377, 497, 69, 840, 142, 597, 477, 315, 623, 925, 370, 130, 610, 329, 347, 523]. case-centric [623]. cases [846]. causal [943]. CBTC [539]. CC [68]. CC-NUMA [68]. CCDL [255]. centered [355]. centric [904, 623]. Certification [963, 148]. CESRBDDs [858]. Chains [231, 92, 468, 610]. challenge [851, 329, 530]. challenges [473, 639, 184, 584, 567, 445, 589, 528, 527]. change [225, 798]. changes [356, 560]. channel [73, 825]. channels [822]. characterization [832]. charts [197, 125]. checkable [84]. checker [602, 276, 794, 46, 600, 881, 8, 633]. Checking [612, 266, 230, 222, 914, 346, 487, 432, 438, 666, 32, 565, 324, 123, 575, 367, 429, 738, 174, 117, 729, 363, 391, 700, 435, 434, 171, 175, 170, 760, 735, 691, 158, 577, 756, 690, 184, 456, 30, 501, 556, 566, 783, 65, 437, 664, 751, 118, 321, 31, 165, 137, 167, 88, 574, 611, 660, 637, 842, 132, 701, 68, 740, 424, 166, 287, 928, 247, 411, 385, 290, 571, 43, 82, 92, 744, 39, 875, 245, 168, 417, 616, 419, 656, 704, 684, 728, 852, 481, 96, 155, 757, 433, 567, 182, 613]. checking [472, 705, 320, 782, 41, 530, 485, 555, 910, 131, 389, 36, 358, 544, 624, 727, 568, 570, 366, 333, 394, 133, 409, 614, 508, 855, 86, 319, 42, 558, 711, 347, 601, 120, 149, 636, 218, 328, 523, 576]. checking-based [656]. checks [591, 668]. chemical [160, 347]. CHEOPS [160]. Chinese [523]. chip [335]. Chisel [713]. CINCO [692]. circuits [56, 50, 912, 479, 720]. circular [661]. Citrus [864]. CkTailv2 [878]. Clara [442]. class [362, 631, 290, 607, 69, 449]. classification [563]. classifying [809]. clinical [347]. clock [7, 905]. Clojure [891]. Cloned [582]. Closed [510]. Closed-loop [510]. cloud [954, 627]. cloud-based [627]. cloud-native [954]. CLP [699]. CLPS [156]. CLPS-B [156]. co [486, 68, 901]. co-simulation [486, 68, 901]. coarse [618, 128]. coarse-grained [618, 128].

COBOL [352]. CoCo [850]. Code

[634, 736, 517, 645, 356, 332, 454, 656, 467, 504, 722, 727, 148, 333, 352, 27]. CoDec [304]. coevolution [775]. CoLiS [889]. collective [772, 771, 943, 945, 951, 940]. collision [519, 671, 871, 604]. colonies [944]. color [274]. color-blind [274]. Colored [264, 608, 284, 280, 282, 25, 22]. Coloured [260, 278, 26, 23, 259, 21, 75, 24, 688]. combinational [61]. combined [537]. Combining [915, 385, 648, 739, 528, 372, 798, 426]. comeback [738]. commercial [235, 539]. Common [428]. communicating [657, 208, 392, 878]. communication [965, 15, 620, 137, 891, 744, 454, 346]. COMP [923, 849]. compact [138]. comparative [402]. Comparing [823, 735, 951, 770]. Comparison [97, 466, 824, 185, 315]. Competition [838, 839, 923, 848, 585, 651, 850, 849, 715]. competitions [847]. compilation [27]. compiled [148]. complement [136]. Complementary [734]. complemented [858]. Complete [607]. completeness [318]. complex [330, 195]. compliance [416]. Component [474, 254, 874, 777, 650, 270, 867, 498, 708, 256, 466, 947, 296, 257]. Component-based [474, 650, 947, 257]. Components [255, 734]. composable [778, 255]. composed [408]. Composition [254, 773, 708, 296]. Compositional [235, 462, 246, 818, 125, 475, 914, 635, 661, 381, 937]. compositionality [76]. compositions [399]. compression [716]. Computational [184, 180]. computer [336, 141, 142, 912]. computer-aided [142]. computer-assisted [141]. computing [73]. concatenation [459]. concept [765]. concepts [6, 3]. conceptual [355]. concurrency [965, 78, 420, 732, 441]. concurrency-related [420]. concurrency-specific [441]. Concurrent [599, 932, 662, 836, 883, 72, 362, 168, 260, 484, 169, 275, 502, 614, 532, 609]. condition [528]. conditioned [250]. conditions [414, 542]. Conference [819, 339, 374, 578]. Configuration [744, 457, 451, 453]. configuring [284]. conflict [108]. confluence [571, 850]. conformance [608, 515, 44]. Congestion [281]. connection [262, 281]. conquer [691]. conscious [946]. consensus [835, 834]. considerations [299]. consistency [666]. Consistent [353]. Constraint [505, 65, 156, 452, 787, 421, 797]. Constraint-based [505, 65, 452]. constraints [203, 735, 616, 742]. construction [667, 789, 886, 657, 470, 749, 147, 214, 911, 882, 905, 549, 854, 750]. constructions [290]. Constructive [461]. consumption [478]. Contention [360, 135, 79]. contest [369]. context [778, 535]. context-dependent [778]. contextual [296]. Continuous [6, 788, 243, 574, 703, 937, 534, 549, 309, 866]. continuous-time [937]. contract [631, 708]. contracts [235, 960, 258]. contributions [325]. Control [141, 40, 281, 778, 630, 746, 83, 906, 519, 538, 225, 511, 828, 317, 864, 540, 534, 719, 309, 860]. controllable [851, 443]. controlled [823, 63]. controller [70]. controllers [919]. convenient [375]. conversion [48]. converting [305]. Cooperative [812]. Coordinating [949, 776]. coordination [773, 941, 128]. Coping [295]. copy [728]. Coqoon [681]. CORBA [270]. core [750, 749, 601, 669, 683]. cornerstones [40]. Correct [831, 941, 749, 630, 657, 813, 750]. Correct-by-construction [749, 657, 750]. correcting [656]. Correction [836, 835, 885, 817, 750, 837, 952]. correctness [102, 441]. correlation [270]. cost [378, 304]. cost-optimal [378]. Counter [131, 122]. Counter-example [131]. counter-examples [122]. counterexample [271]. counterexamples [580]. counterstrategies [496]. Counting [758, 622, 318]. course [279]. Coverage [229, 430, 402, 404, 341, 288, 770, 927]. Coverage-biased [430]. Covering [138]. CoVeriTest [812, 840]. CPA [841]. CPA/Tiger [841]. CPA/Tiger-MGP [841]. CPN [260, 265]. CPS [833, 957]. CPSDebug [833]. Crawlability [397]. Creating [25, 953]. criteria [734, 288, 441]. Critical [959, 653, 645, 186, 187, 697, 652, 191, 658, 899, 694, 226, 693, 868]. critical-path [658]. cross [404]. cross-product [404]. CRV [715]. CSP [487, 600, 740, 358, 636]. CTL [249, 175, 751, 88, 616, 42]. CTL* [290]. CTL-property [249]. cube [215]. current [245]. custom [965]. CVT [27]. cyber [753, 678, 897, 936]. cyber-physical [753, 678, 897, 936].

Dafny [960]. Data [382, 380, 662, 836, 675, 138, 817, 816, 543, 448, 875, 484, 40, 623, 731, 932, 627, 564, 784, 215, 860, 151, 121]. Data-abstraction [382]. data-aware [623]. data-intensive [627]. dataflow [164]. Datagram [281]. DBM [905]. DBM-based [905]. Debian [889]. Debugger [737]. Debugging [496, 403, 738, 362, 388, 105, 737]. decentralized [861]. decidable [723]. deciding [501]. Decision [926, 328, 780, 858, 380, 326, 573, 55, 325, 919, 392, 572, 182, 925, 311, 58, 669]. Decision-diagram-based [328]. Decision-making [926]. decisions [457]. declarative [880]. decomposition [859]. decompositions [964]. Deductive [960, 413, 709, 65, 791]. deep [928, 913]. defect [398]. defect-prone [398]. defence [283]. defense [800]. defined [503]. definitely [201]. delta [403, 747, 581, 814]. delta-oriented [747]. demonstration [767]. dense [558]. dense-time [558]. dependence [698]. dependencies [414].

Dependency [786, 856, 878]. dependent [778]. depth [289, 599]. depth-first [289, 599]. described [653]. Description [255, 252, 101, 545]. descriptions [77, 280]. Design [249, 60, 238, 517, 357, 473, 827, 675, 941, 293, 335, 106, 579, 68, 750, 749, 465, 678, 674, 113, 95, 798, 474, 11, 832, 70, 36, 3, 295, 860]. design-space [579]. designs [109, 250]. desktop [398]. Detecting [53, 797, 392, 821, 557]. detection [517, 779, 853, 484, 902, 96, 896, 399, 108, 732, 394, 379, 508]. deterministic [884, 684, 871]. determinization [178]. developer [299, 19]. Development [134, 251, 645, 299, 273, 187, 559, 682, 258, 454, 113, 277, 381, 279, 349, 545, 360, 516, 240, 303, 301, 226]. Developments [506]. devices [299, 142, 510]. devoted [754]. DFT [869]. Diagnosing [808]. diagnosis [496]. diagram [328]. diagrams [796, 780, 858, 380, 56, 460, 55, 608, 522, 311, 58, 669, 120]. Diamont [935]. digital [509]. Directed [321, 137, 167, 924]. Discourje [891]. DisCoveR [880]. Discovering [282, 799]. discovery [880]. Discrete [703, 7, 575]. discrete-state [575]. distance [321, 219, 658]. distance-preserving [321]. Distributed [172, 192, 753, 198, 365, 776, 174, 117, 773, 176, 171, 170, 234, 935, 230, 470, 330, 647, 477, 11, 320, 936, 73, 475, 86, 150, 455]. distributing [109]. distribution [175, 264, 195]. distributions [549]. Diversity [752, 689, 450]. divide [691]. divide-and-conquer [691]. divider [104]. DivSIM [876]. document [864, 145]. documents [539, 961]. Domain [436, 649, 381, 410, 692, 731, 560, 860, 961]. Domain-specific [436, 692, 860, 961]. domains [268, 574, 425]. Don’t [359]. doors [519]. Downward [593]. DReAM [774]. Driven [540, 486, 906, 608, 258, 726, 692, 591, 526, 644]. driving [941]. DSSs [701]. during [465]. Dynamic [535, 342, 777, 158, 84, 774, 829, 935, 168, 399, 485, 731, 73, 365]. dynamically [622]. dynamics [945].

E-LOTOS [18]. early [465, 732]. Easy [78]. Ecdar [462]. eclipse [958, 354]. Eclipse-based [354]. Eddy [320]. edelta [865]. edge [858]. edge-specified [858]. edges [858]. Editor [20, 64, 90, 71, 152, 100, 395, 331]. Editorial [1]. editors [111, 146, 241]. EdSketch [726]. education [29, 503]. educational [72]. effect [596]. effective [502]. effectiveness [288, 285]. effects [545]. efficiency [288, 852]. Efficient [203, 93, 105, 664, 792, 230, 157, 616, 133, 58, 931, 286, 151, 880, 359, 794, 813, 674, 17, 841, 220, 720]. eighth [850]. electronic [299, 280, 3]. Electrum [765]. elimination [590]. embeddable [364]. Embedded [207, 478, 473, 241, 234, 332, 114, 750, 749, 513, 592, 475, 333, 466]. EMF [374]. empirical [930, 711]. emptiness [668]. emulated [225]. Enabling [788]. Encoding [336, 149]. end [196, 475]. end-of-production [196]. end-to-end [475]. Energy [832, 746]. enforce [444]. enforcement [823, 650]. engine [844]. engineering [772, 771, 539, 769, 457, 104, 140, 761, 302, 679, 242, 304, 81, 295, 269, 940]. engines [953]. enhancing [354]. enriched [212]. ensembles [778, 777]. Envelopes [871]. Environment [644, 77, 63, 486, 105, 727, 354, 428]. Environment-driven [644]. environments [196, 626]. equation [204, 890]. equations [32, 758]. Equivalence [631, 607, 613]. ERLANG [112, 134, 411]. Error [219, 153, 902]. errors [732, 394]. ERTMS [764, 762, 765, 766, 870, 769, 768]. ERTMS/ETCS [764, 762, 765, 870, 769, 768]. ESBMC [842]. Estimating [237, 356]. Estimation [468, 738, 465, 825]. ETCS [764, 762, 765, 870, 769, 767, 768]. ETI [4, 5]. evaluate [370]. Evaluating [402, 596, 288]. Evaluation [376, 383, 611, 920, 60, 910, 544]. event [238, 904, 894, 608, 805, 955, 360, 282, 528, 387, 763, 336, 643, 766, 721, 641, 634, 903, 640]. event-B [360, 387, 763, 643, 766, 721, 641, 634, 903, 640]. event-condition-action [528]. event-driven [608]. events [794, 341, 932, 558]. evergreens [363]. evolution [865, 559, 457, 512, 298, 514, 350, 395]. evolutionary [711]. evolving [561]. Exact [676]. examination [525]. example [349, 131]. examples [122]. exceptions [630]. Exchange [263]. Executable [74, 25]. executing [455]. Execution [737, 265, 323, 478, 738, 811, 844, 846, 114, 658, 958, 726, 344, 431, 329, 150, 22]. execution-driven [726]. execution-time [114]. exemplified [560]. experience [870, 113, 689, 144]. Experiences [240, 649, 467]. experiment [332, 18]. Experimental [372, 238]. experimentally [288]. experimentation [54]. experiments [639, 60, 384, 606]. ExPLAIn [918]. explainability [920]. explainable [919]. explanation [833, 219, 921, 918]. explicit [118, 137, 668, 409]. explicit-state [118, 137]. Exploiting [495, 287, 169, 379]. exploits [718]. exploration [478, 271, 202, 430, 465, 267, 240, 866]. Exploring [870, 154, 322]. Expressing [84]. Expression [120]. expressions [821, 286]. Expressive [520, 375, 931]. expressiveness [243]. Extended [856, 524, 892, 908, 938, 705, 793, 791, 898, 42]. Extending [857]. extensible [222]. extension [631, 820]. extensions [152]. Extracting [782]. Extrapolating [433, 348].

Facilitating [457]. factories [702]. facts [343]. failure [833]. Fair [706, 558]. FairFuzz [843]. FairFuzz-TC [843]. fairness [735]. false [860]. Falsification [483]. family [664, 751, 425]. family-based [664, 751]. Farkas [859]. FASE [810, 253, 748]. FASE’17 [707]. Fast [732, 375, 902, 310]. Fate [153]. fault [894, 709, 313, 191, 855]. fault-tolerance [313]. fault-tolerant [855]. FDR3 [600]. feasible [122]. feature [247]. featured [746, 745]. featuring [814]. Feyerabend [13]. field [319]. Fighting [118]. filters [509]. Fin [729]. Fin-less [729]. final [715]. Finding [420, 122]. Finite [231, 38, 779, 875, 606, 169, 133, 898, 676]. finite-state [38, 169, 676]. FireWire [18]. firm [477]. First [715, 839, 923, 289, 820, 651, 599, 806, 490]. first-order [820, 806]. FISh [37]. Flexibility [334]. Flexible [660, 359]. floating [915, 136, 104]. floating-point [915, 104]. flocks [944]. Florida [29]. flow [630, 817, 816, 467]. fluidic [295]. Flush [349]. fly [117, 266, 178, 204, 705, 601, 44]. FO [803]. focus [665]. force [527]. forces [756, 919]. Forest [921]. forests [920]. form [747]. Formal [238, 141, 232, 873, 874, 460, 519, 538, 225, 418, 721, 959, 191, 899, 70, 233, 140, 261, 108, 722, 815, 248, 693, 687, 868, 965, 357, 186, 713, 51, 870, 847, 187, 234, 229, 105, 788, 963, 650, 223, 769, 68, 697, 497, 652, 767, 542, 142, 29, 104, 40, 318, 107, 496, 103, 897, 900, 768, 689, 761, 61, 596, 183, 240, 431, 961, 2, 257]. Formalising [766]. formalism [636]. Formalization [45]. Formalizing [338]. Formally [411, 964, 143, 671]. formula [85]. formulas [448, 570]. formulation [179]. fragment [695]. Framework [308, 102, 487, 666, 442, 777, 234, 271, 718, 774, 361, 313, 109, 660, 887, 252, 452, 270, 67, 908, 381, 629, 937, 440, 311, 722, 348, 516, 222, 898, 785, 669]. frameworks [696, 689, 296]. free [747, 153, 204, 857].

FreeRTOS [521]. frequency [817, 816]. frequentist [824]. friendly [738]. FSAP [244]. FSAP/NuSMV [244]. FSAP/NuSMV-SA [244]. FTSyn [313]. Fujaba [161, 377]. Full [895, 870, 692]. Full-program [895]. Fully [650]. fUML [487]. Functional [414, 491, 185, 550, 402, 196, 262, 670, 647, 318, 198]. functions [414, 597, 787, 613, 797]. fundamental [269]. fuzzer [843].

game [382, 246]. games [180, 730, 703, 685, 890, 423, 685]. gap [679]. gas [560]. gear [643, 639, 642, 70, 641, 640]. general [666, 322, 391, 45]. generalized [729, 424, 668]. generated [500]. Generating [77, 948, 580, 115, 877, 502, 610, 961, 846]. generation [796, 565, 91, 158, 386, 271, 338, 611, 842, 543, 68, 448, 524, 542, 840, 671, 597, 845, 129, 692, 124, 731, 131, 672, 340, 722, 591, 242, 634, 841, 304, 216, 526, 532, 199, 364, 605, 604]. generator [17]. generators [752]. generic [713, 550, 204, 516]. genetic [154, 656]. GenUTest [340]. GIOP [45]. GNATprove [586]. GNU [118]. Go [908]. Go2Pins [908]. goal [896, 841]. good [81]. GPSO [759]. GPU [783]. GPUs [601]. GraBaTs [377]. grade [386]. grained [618, 128]. grammar [877]. grammars [813, 524]. Graph [369, 801, 753, 561, 813, 258, 59, 375, 376, 370, 712, 815, 814, 752]. Graph-based [801, 376]. graphical [682, 692]. graphics [391]. Graphillion [594]. graphs [630, 786, 856, 594, 878, 528]. Greedy [716, 759]. Greybox [530]. GrGen.NET [375]. grid [523]. GRL [896]. GROOVE [427]. group [691]. GSDetector [896]. guarantee [887]. guarantees [939]. Guard [617]. Guard-based [617]. Guest [152, 395, 241, 331]. guide [21]. Guided [618, 286, 489]. GUMP [921].

Haifa [339]. Handel [190]. Handel-C [190]. Handling [637]. hard [341]. hard-to-reach [341]. hardware [197, 179, 95, 67, 101, 36, 250]. Haskell [185]. HASL [575]. hazards [853]. HCI [238]. heads [318]. Healing [316]. Heerhugowaard [88]. held [29]. Herschel [556]. heterogeneous [429, 486, 107, 121]. Heuristics [166, 60, 431]. Hierarchical [103, 861]. High [696, 49, 779, 674, 200, 73, 358, 720]. High-automation [49]. high-availability [200]. High-level [696, 674, 73, 358, 720]. high-speed [779]. Higher [101, 550, 857]. Higher-level [101]. higher-order [550, 857]. Highlights [177]. highly [662, 836]. Hip [521]. Hip/Sleek [521]. HiPE [113]. History [511]. hit [341]. hoc [721, 582]. HOL [415]. home [861]. homes [861]. Hoorn [88]. Hoorn-Kersenboogerd [88]. horizon [231]. horizontal [871]. human [336]. human-computer [336]. Hybrid [717, 406, 442, 618, 408, 574, 769, 297, 513, 733, 767, 8, 417, 671, 103, 155, 6, 768, 708, 483, 595, 837, 423, 309, 956, 409, 764, 643, 762, 765, 766]. HybridUML [213]. hyperproperties [792]. hypervisor [418]. hypothesis [568]. HYTECH [8, 297].

i-protocol [118]. IC3 [757]. identification [676]. identify [394]. identities [867]. IEEE [135, 360, 18, 79]. IEEE-1394 [18]. ignorant [562]. ignoring [368]. II [847]. imaging [760]. impact [225]. imperfect [735]. Implementation [588, 307, 717, 628, 330, 674, 280, 311, 722]. Implementation-level [588]. implementations [474]. Implementing [76, 362]. improve [91]. Improved [388, 250, 475]. improvement [167]. improvements [700]. Improving [655, 290, 285, 912, 656]. IMS [307]. incomplete [782]. inconsistencies [453]. increase [468]. Incremental [565, 811, 249, 660, 814]. incrementalization [95]. independent [943]. induction [802, 895, 637, 50]. Inductive [228]. Indus [275]. Industrial [959, 186, 51, 196, 535, 386, 611, 652, 648, 899, 467, 770, 640, 52, 868]. industrialization [223]. industry [536, 596]. inequalities [425]. inference [802, 425]. influence [25]. informal [187]. information [735, 28, 813, 258, 215]. infrastructure [80, 869]. Infusion [141]. inheritance [785]. inhouse [130]. initialization [164]. injection [860]. inline [562]. Innovation [514]. input [795]. instead [324]. instrumentation [931]. integer [663, 613]. integrate [626]. Integrated [697, 487, 662, 836, 727, 514]. Integrating [4, 28, 701]. Integration [187, 161, 163, 788, 247, 162, 665, 160, 866, 449, 3]. intelligent [647]. intensive [743, 592, 627]. intent [557]. intent-based [557]. Inter [45, 744]. Inter-ORB [45]. inter-process [744]. Interacting [5, 950]. interactions [336]. Interactive [206, 737, 680, 876, 961]. Interface [864]. interfaces [474, 75, 864]. interleaving [840]. interlocking [542]. interlockings [541]. International [819, 578, 715, 839, 923, 395]. interoperability [338, 516]. interpolants [859]. interpretation [419, 481]. interval [417]. interval-based [417]. interworking [306]. introducing [545]. Introduction [374, 781, 299, 762, 929, 159, 725, 934, 819, 754, 907, 253, 292, 298, 439, 269, 488, 241, 331, 39, 152, 463, 395]. Introductory [177, 139, 119, 193, 227, 127, 186, 170, 165]. intrusion [779, 394]. Intuition [918]. invariant [802, 148]. invariants [883, 895]. Invasive [254]. investigate [945]. invisible [221]. involving [574]. IOA [330]. IOCO [446]. IoT [799, 832, 801]. ISDN [306]. isolation [476, 420]. Issue [929, 934, 754, 612, 886, 781, 819, 652, 911, 907, 882, 854]. Issues [171, 41, 110, 485]. ISUP/ISDN [306]. Iterative [403]. iUML [766]. iUML-B [766]. IVE [386].

jABC [629]. JAVA

[43, 915, 630, 645, 386, 126, 166, 726, 69, 351, 275, 931, 143, 86, 686, 710]. Jensen [20]. JML [188, 386, 258, 222, 686]. JML-annotated [686]. join [919]. Joint [756]. journey [906].

Kaveri [275]. KDM2PSM [953]. Kermeta [376]. kernel [728, 80]. Kersenboogerd [88]. KeY [915, 588, 892]. KeYmaera [595, 837]. kinetic [945]. KIV [585]. KLAIM [949]. KLEE [844]. Knowledge [954, 724]. Knowledge-based [724]. KRONOS [135, 9].

labeled [594]. Ladder [53, 900]. lambda [857]. lambda-free [857]. landing [643, 639, 642, 641, 544, 640]. Language [950, 307, 777, 164, 330, 805, 800, 677, 674, 826, 512, 62, 860, 84, 255]. language-to-language [826]. languages [16, 579, 581, 381, 185, 334]. large [429, 430, 154, 206, 594, 104, 572, 295, 453]. large-scale [453]. latencies [475]. Lava [106]. layer [18]. layered [332]. leaks [825]. learned [343, 377]. Learning [773, 795, 878, 711, 531, 963, 818, 342, 947, 784]. Learning-based [773]. LearnLib [348]. less [729]. lessons [377]. Let [587]. Level [764, 762, 765, 766, 769, 768, 196, 588, 161, 696, 327, 862, 767, 592, 59, 674, 467, 128, 101, 73, 358, 931, 720]. Leveraging [909, 918, 233]. libraries [454, 550, 498, 303]. library [362, 69, 594, 204]. life [767]. lifecycle [559]. light [67]. light-weight [67]. Lightweight [128, 573, 437, 933, 596]. like [151]. likely [784]. limit [884]. limit-deterministic [884]. Line [578, 266, 457, 454, 216, 281]. linear [405, 859, 408, 547, 884, 730, 491, 425, 787, 797]. linear-constraint [787, 797]. linearizability [615]. lines [456, 580, 274, 742, 582]. Link [18]. LinkedList [892]. linking [636]. Linux [728]. lists [614]. literature [537]. Live [197]. livelock [118]. Liveness [221, 412, 212]. LLVM [876]. load [863, 198]. load-time [863]. Local [31, 407]. locality [169]. localization [709]. locking [134]. locks [392]. Log [879]. logic [16, 884, 115, 678, 875, 61, 793, 815, 814, 849, 421, 857, 724, 53]. logic-based [16, 814]. logical [149]. logics [760]. logistics [649, 283]. logs [904, 282]. loop [895, 888, 500, 510]. loops [547, 637, 901, 787, 797]. loss [813]. LOTOS [18]. low [327]. low-level [327]. Lower [217]. LTL [438, 367, 495, 424, 820, 173, 803, 908, 358, 483, 366, 409, 379, 85]. LTL-FO [803]. Lurette [236]. Lustre [349].

Machine [947, 643, 963, 676, 784, 364]. Machine-learning [947]. machines [795, 919, 606, 898]. magic [590]. Main [110]. maintainer [889]. maintenance [399]. makes [81]. making [926]. malware [508]. managed [582]. management [63, 262, 281]. Managing [98]. manipulation [59, 58]. manipulator [121]. Manual [373]. Many [601]. Many-core [601]. Mapping [579]. March [29]. markings [758, 808]. Markov [573, 390, 92, 572, 231]. Markovian [894]. mastering [559]. Masterminding [798]. matching [372, 61, 256]. mathematical [597]. MCC [852]. MCMAS [633]. mCRL2 [874]. MDA [251]. MDA-based [251]. MDP [571]. MDPs [893]. me [215]. meaning [2]. measure [216]. Measurement [464]. Mechanical [79]. mechanisms [264]. Mechanized [52]. mechatronic [293, 24, 292, 294]. mecum [19]. medical [760, 142, 510]. meet [51, 499]. meets [589]. Melbourne [29]. Memory [237, 666, 367, 63, 756, 327, 791]. message [507, 125, 785]. message-passing [507]. meta [161]. meta-model [161]. metabolic [74]. metamodel [865]. metaSMT [665]. method [336, 266, 877, 291, 360, 216, 281]. method-based [877]. methodological [545]. Methods [959, 233, 357, 186, 51, 870, 847, 827, 234, 788, 538, 223, 418, 652, 142, 29, 899, 140, 689, 464, 73, 2, 693, 687, 868]. metric [815]. metrics [780, 229, 219, 397, 927]. MGP [841]. microgrid [577]. microprocessor [102]. MIDAS [649]. middleware [200]. migrating [352]. migration [351]. minimisation [683]. minimization [192, 549]. minimized [34]. Mining [678, 628]. missile [23]. mixed [479]. mixed-signal [479]. ML [185]. Mobile [237, 73, 22, 149]. mock [340]. modal [32, 751, 946]. Model [612, 729, 363, 700, 273, 456, 230, 448, 247, 524, 830, 43, 258, 646, 698, 481, 279, 472, 540, 371, 755, 445, 591, 36, 130, 544, 748, 333, 614, 86, 558, 636, 638, 609, 432, 438, 666, 238, 32, 565, 324, 602, 123, 575, 367, 429, 738, 174, 117, 773, 276, 486, 811, 391, 734, 435, 434, 171, 175, 170, 794, 760, 161, 735, 691, 158, 577, 327, 709, 690, 46, 184, 30, 501, 556, 566, 783, 906, 65, 437, 664, 751, 118, 321, 31, 817, 816, 165, 137, 167, 88, 574, 611, 422, 657, 580, 813]. model [637, 842, 132, 701, 68, 740, 750, 749, 424, 461, 166, 287, 928, 252, 411, 385, 290, 810, 767, 571, 82, 881, 8, 648, 92, 744, 39, 678, 607, 147, 245, 168, 417, 616, 510, 419, 270, 656, 704, 684, 728, 496, 852, 96, 155, 757, 433, 567, 182, 633, 705, 320, 373, 782, 41, 761, 530, 485, 555, 910, 131, 389, 358, 242, 595, 837, 624, 568, 222, 570, 394, 133, 409, 508, 855, 319, 42, 711, 347, 601, 149, 218, 523, 576]. Model-based [524, 830, 646, 698, 279, 755, 445, 130, 748, 638, 609, 734, 580, 750, 749, 461, 648, 678, 607, 496, 761]. Model-checking [247, 472, 86, 558, 429, 735, 68, 92, 616, 131]. Model-Driven [540, 258, 591, 486, 906].

Modeling [769, 480, 704, 318, 274, 897, 641, 374, 873, 383, 460, 579, 721, 581, 592, 674, 452, 6, 864, 692, 640, 334, 869, 720]. Modelling [262, 577, 944, 283, 427, 73, 264, 688, 387, 774, 541, 260, 160, 455, 523, 741]. Models [265, 487, 666, 965, 645, 880, 833, 355, 904, 753, 561, 416, 620, 430, 49, 74, 828, 390, 924, 75, 129, 896, 210, 348, 878, 903, 752, 784, 961, 25, 711]. Modular [267, 888]. modules [191]. modulo [603, 417]. moment [56]. monitor [822, 248]. Monitoring [939, 603, 479, 962, 824, 412, 753, 861, 935, 792, 553, 157, 443, 803, 938, 832, 937, 936, 933, 806, 562]. Monitors [957, 955]. monotony [963]. Montages [77]. MontiCore [381]. MoonLight [933]. Moore [795]. MOP [440]. MoTMoT [371]. movement [690]. moving [870]. MSC [459]. MSO [438]. MTL [243, 695]. MTL- [695]. Multi [862, 449, 683, 643, 293, 750, 749, 739, 685, 633, 945, 841, 86, 669]. multi-agent [293, 633, 945]. Multi-core [683, 750, 749, 669]. multi-goal [841]. Multi-level [862]. multi-machine [643]. Multi-objective [449]. multi-player [685]. multi-threaded [739, 86]. multiobjective [294]. multiphase [797]. multiple [949, 691, 536, 872, 685]. multiplier [912]. multiprocessor [68]. multithreaded [220]. multivehicle [871]. Mur [231]. mutation [126, 770, 752]. mutually [818].

NASA [248]. native [954]. navigation [354]. needed [356]. needs [51]. negative [289]. Net [80, 909, 758, 289, 74, 631, 703, 75, 73]. Nets [260, 278, 264, 174, 76, 699, 26, 608, 23, 284, 290, 259, 280, 21, 267, 24, 282, 25, 22, 688]. network [779, 928, 755, 195]. networks [404, 923, 948, 341, 283, 704, 924, 938, 242, 922, 913, 825, 927]. Neural [824, 923, 948, 928, 924, 938, 922, 913, 825, 927]. next [386, 671, 604]. next-generation [671]. nightly [866]. non [894, 887, 888, 871]. non-deterministic [871]. non-Markovian [894]. non-termination [888]. nondeterminism [794, 574]. nondeterministic [606]. nonlinear [417, 962]. notation [77, 511]. notations [136, 252]. novel [539]. NUMA [68]. numeric [385]. Numerical [218, 655, 425]. NUSMV [46]. NuSMV-SA [244]. nutshell [10].

object [904, 187, 109, 670, 553, 504, 455]. object-centric [904]. object-oriented [187, 109, 670, 553, 455]. objective [449]. objectives [685]. objects [208, 351]. obtaining [475]. OC [904]. OC-PM [904]. OCL [416, 212, 638]. ODE [417]. oil [560, 874]. omega [433]. omega- [433]. OMT [11]. On-the-fly [705, 44, 117, 266, 178, 204, 601]. one [358]. Online [220, 28, 115, 937, 806, 294]. only [182]. Ontology [251, 256]. Open [503, 387, 633, 471]. open-source [633]. OpenJDK [892]. operating [965]. operational [190, 272, 261, 228, 869]. operators [268, 705]. opinions [29]. optimal [378]. optimization [759, 83, 830, 373, 801, 449, 294]. optimizing [293, 292, 947]. options [451]. oracles [115, 415]. ORB [45]. Order [739, 780, 322, 35, 93, 167, 368, 571, 820, 617, 550, 890, 806, 449, 857, 365]. ordered [730]. ordering [60]. orders [480]. oriented [412, 408, 187, 747, 109, 670, 553, 648, 254, 455]. Oris [383]. oscillatory [575]. output [795]. overhead [443]. overview [188, 585, 463, 440, 309, 927]. OWL [251].

package [59]. packages [889]. PacoSuite [254]. PAG [17]. Paper [227, 186, 177, 170, 165, 892, 139, 119, 193, 127]. Papers [725, 911]. paradigm [775]. Paradigms [918]. Parallel [391, 733, 173, 62, 320, 170, 600, 722, 668, 431]. parallel-code [722]. parallelisations [831]. parameterised [890]. Parameterized [619, 621, 867, 624, 138, 437, 620, 480]. parametric [723, 893, 853, 390, 678, 742, 807]. parity [884, 730, 890]. Partial [32, 322, 167, 890, 35, 93, 368, 571, 480, 617, 365, 739]. Partial-order [322, 167, 890, 93, 617]. particle [759]. partition [830]. partitioning [841]. passing [507, 785]. past [297]. Path [408, 658, 787, 797]. Path-oriented [408]. PathFinder [43]. pathways [74]. patient [280]. Pattern [205, 943, 372, 62, 722, 593]. Pattern-based [205, 722]. patterns [517, 15, 821]. Paul [13]. pebbling [716]. peer [162]. peer-to-peer [162]. Pentium4 [104]. perform [356]. Performance [307, 388, 592, 373, 951, 242, 475]. periodic [106]. perspective [552, 36, 719, 956]. perspectives [473, 567, 450]. Petri [688, 909, 174, 76, 758, 26, 608, 74, 23, 284, 290, 260, 259, 278, 703, 280, 80, 21, 267, 75, 24, 264, 282, 25, 22]. phase [507]. phases [465]. PHAVer [297]. Phone [237, 22]. physical [753, 678, 897, 936]. physics [632]. piggyback [553]. pilot [352]. pipeline [853]. Plain [845]. planning [499, 535, 702, 572, 261, 502]. plans [500]. Platform [244, 889, 4, 386, 519, 737, 160, 3]. Platitudes [99]. player [685]. PLC [83]. Pleak [862]. PM [904]. point [915, 136, 299, 47, 104]. pointwise [243]. policies [273, 828, 317]. Policy [562, 624]. PolyGraph [816, 817]. polyhedral [909]. polynomial [730]. porting [400]. POSIX [416]. power [478, 925]. powerset [268]. Practical [649, 200, 467, 42, 40, 496]. practice [310, 55, 648]. practices [679]. practitioner [21]. Pragmatics [30]. pre [414]. pre-conditions [414]. precise [738, 509, 426]. Predicate [936, 327, 385, 840]. predictive [824, 220]. Preface [612, 401, 64, 90, 207, 652, 20, 71, 419, 111, 804, 100, 345, 146]. prefix [792]. presented [810, 707, 659, 748]. preservation [247, 814]. preserving [321, 922]. prialt [190]. price [423]. priced [746]. principles [767, 194]. printed [28]. printfs [287]. PRISM [135, 684, 155, 685]. PRISM-games [685]. privacy [862]. ProB [642, 291]. Probabilistic [390, 155, 604, 391, 881, 744, 684, 705, 711, 218]. probabilities [926]. probability [549]. problem [179, 368, 295, 449]. problems [909, 746, 696, 376, 347]. procedural [818]. procedure [326, 663, 392, 182]. procedures [262, 325, 281]. Process [249, 336, 953, 645, 880, 904, 628, 299, 246, 559, 744, 261, 545, 27, 516, 160]. processes [602, 575, 573, 862, 622, 470, 284, 392, 572, 623, 149]. processor [26]. processors [738, 391]. Product [578, 404, 299, 456, 580, 457, 454, 274, 742, 582, 301]. production [196, 11]. productivity [468]. products [457]. profile [213, 209, 211]. Program [89, 14, 82, 930, 518, 492, 517, 134, 713, 531, 831, 488, 895, 16, 661, 848, 520, 287, 481, 17, 370, 493, 710, 736]. Programming [829, 955, 37, 949, 412, 422, 800, 656, 185, 62, 724]. programs [915, 53, 802, 776, 630, 123, 507, 362, 895, 126, 655, 935, 670, 637, 663, 166, 553, 43, 739, 484, 908, 613, 900, 275, 502, 220, 931, 526, 86, 791]. progress [216]. project [72, 891, 463, 352]. ProM [879]. PROMELA [45, 41, 380]. prone [398]. Proof [104, 428, 335, 716]. proof-based [335]. proofs [405, 547, 49, 661, 888, 482, 379]. Propagation [914]. Properties [136, 312, 942, 691, 84, 49, 745, 447, 887, 266, 247, 157, 678, 900, 479, 962, 933, 555, 591, 483, 360, 712, 409, 898, 676, 784, 601]. Property [526, 249, 924, 785]. property-based [785]. property-directed [924]. Property-driven [526]. propositions [558]. PROSPER [94]. Protocol [12, 45, 135, 118, 31, 631, 110, 263, 41, 97, 18, 79, 721, 360, 281]. protocol-extension [631]. protocols [602, 205, 93, 137, 891, 472]. Protos2CPN [284]. prototypes [550]. prototyping [145, 503]. Provably [630]. prove [595, 837]. Proved [360]. prover [50, 48, 857, 415]. Proving [224, 47, 318]. PRTest [845]. public [690]. publish [162]. publish/subscribe [162]. Publisher [952]. Pump [141]. purpose [391, 338]. Pushdown [508]. Putting [232].

QoS [555]. Qualitative [694, 522, 793, 719]. quality [775, 707, 200, 534, 302, 748]. quantified [448]. Quantifying [512]. quantiles [684]. Quantitative [745, 522, 825, 523, 741, 615, 241, 33, 694, 937, 793, 564]. quantized [509]. quasi [730]. quasi-linear [730]. quasi-polynomial [730]. queries [753]. query [875].

Rabin [729]. race [484]. races [53]. radio [503]. radiotherapy [199]. railway [88, 538, 542, 541, 540, 688, 869]. railways [689]. RAMBUTANS [672]. random [430, 920, 420, 845]. randomisation [215]. randomization [732]. Randomized [902, 835, 834]. ranking [221, 787, 797]. rapid [920, 503]. rare [794, 894, 843]. rating [565]. Rby [719]. RCU [728]. RE [355]. RE-UWA [355]. reach [341]. reachability [909, 172, 408, 327, 733, 390, 325, 902, 133, 644, 286, 328]. reachable [34]. Reactive [790, 211, 750, 749, 258, 208, 525, 651, 610, 490, 121]. read [201, 728]. read-copy [728]. Real [7, 546, 207, 965, 759, 653, 383, 241, 33, 743, 476, 635, 462, 114, 805, 209, 767, 704, 474, 6, 477, 742, 910, 755, 130, 360, 9, 257]. real-life [767]. Real-time [7, 546, 965, 653, 383, 241, 33, 743, 476, 635, 462, 114, 805, 209, 704, 474, 6, 477, 742, 910, 360, 9, 257]. real-world [759]. realistic [526]. realizability [501]. reasoning [915, 387, 731, 712]. reasoning-based [731].

Rebeca [785]. recognition [466]. reconfigurable [950, 774, 829, 191]. record [280]. recover [355]. Recovery [412, 650]. recurrent [924]. recursive [106, 818]. reduce [304]. Reducing [405]. reduction [758, 176, 322, 779, 35, 167, 368, 740, 571, 617, 890, 731, 133, 365, 739]. reductions [909, 858, 93, 168]. Redundancy [914]. Refactoring [305, 747, 371, 370]. reference [562]. refined [579]. Refinement [107, 615, 382, 359, 335, 407, 600, 510, 597, 182, 768, 426, 593, 790]. Refinement-based [107, 768]. refinements [349]. refutation [802]. refuting [898]. registries [626]. regression [396, 563, 638]. Regular [432, 438, 436, 435, 434, 437, 887, 433]. reinterpretation [393]. related [420]. Relating [458]. relations [158]. relaxation [615]. Relay [53]. reliability [478, 346]. Reliable [714, 773, 885]. remove [747]. remove-free [747]. reordering [684]. repair [887, 814]. repairable [894]. repairing [453]. repeatable [718]. replace [747]. replace-free [747]. replaying [457]. report [113, 689, 850]. repository [675, 471]. representation [34, 954]. representations [698]. Require [654]. Requirements [145, 265, 714, 885, 675, 197, 49, 539, 769]. Requirements-document-based [145]. RERS [530, 851, 528, 527]. research [386, 81, 503, 319]. resets [423]. resolution [405, 204, 379, 149]. resource [134]. Restoring [561]. result [289, 52]. results [715, 852, 782, 389, 866]. resuscitation [141, 142]. Reusable [965, 343]. Revealing [453]. review [537, 548]. revisited [756, 556, 424, 554]. reward [423]. rewrite [375]. rewriting [863]. rich [932, 319, 636]. rigid [835, 834]. Rigorous [645, 473, 771, 525, 940, 827, 922]. Risk [534, 537, 536, 533, 798, 560, 869]. Risk-based [534, 536, 533]. river [701]. road [916]. Robotics [957]. robots [949]. robust [361]. Robustness [957]. Rodin [387]. role [675]. Root [360, 135, 79]. ROS [949]. ROS-based [949]. round [835, 834]. round-rigid [835, 834]. Routing [721]. RTAMT [957]. RTLola [916]. Rubik [215]. Rule [517, 239, 554, 915, 811, 948, 807]. rule- [915]. Rule-based [517, 239, 554, 811]. rule-specific [948]. rules [715, 820, 258, 48]. run [891]. run-time [891]. run.time [753]. Runtime [715, 929, 552, 934, 819, 695, 441, 957, 412, 753, 861, 444, 650, 809, 805, 553, 554, 820, 440, 591, 932, 439]. runway [248].

SA [244]. SaBRe [863]. safe [865, 671, 871, 782]. safely [343]. Safety [244, 490, 645, 412, 187, 756, 266, 697, 461, 191, 157, 542, 694, 483, 595, 956, 220, 248, 855, 226, 601, 837]. safety-critical [645, 187, 697, 191, 694, 226]. Sampling [742, 573]. sans [895]. SAT [324, 602, 660, 230, 317, 417, 912, 962, 183, 710]. SAT-based [660, 602, 230, 183]. satellite [454]. Satisfiability [337, 120, 182, 366]. SATMC [602]. saturation [202]. SBIP [555]. SCADE [349]. SCADE/Lustre [349]. scalability [119]. Scalable [367, 738, 117, 806, 709, 872, 851, 425, 128, 627]. scale [453]. Scenario [893, 416, 775]. Scenario-based [893, 416]. scenarios [628, 755, 676]. Schedulability [556, 477, 910]. schedule [83]. scheduler [521]. schedules [378, 782]. Scheduling [374, 499, 179, 696, 830, 647, 279, 347]. schemas [501]. sciences [564]. Scientific [628, 625, 629]. screen [519]. scripting [677]. scripts [889]. SDL [11]. search [618, 173, 599]. secrecy [205]. Section [233, 186, 170, 762, 30, 64, 771, 165, 201, 578, 90, 325, 71, 259, 278, 419, 111, 314, 100, 253, 350, 292, 298, 439, 269, 940, 395, 146, 20]. secure [277, 798]. SEcurity [557, 116, 560, 602, 734, 189, 561, 93, 559, 563, 418, 798, 314, 624, 724]. SED [737]. Selected [485, 725]. Selection [431, 734, 917, 952, 609]. selective [863]. self [773, 293, 292, 947]. self-composition [773]. self-optimizing [293, 292, 947]. semantic [382, 393]. Semantics [16, 713, 406, 190, 243, 246, 272, 208, 228, 922]. Semantics-based [16]. semantics-preserving [922]. sensor [704]. separation [849]. sequence [796, 197, 158, 608, 125]. Sequential [174, 181]. sequentialization [739]. sequentialization-based [739]. serial [18]. service [877, 648, 646, 399, 549]. service-oriented [648]. services [773, 351, 472, 626, 198, 352]. session [5]. set [343, 841]. sets [594, 491, 706]. setting [788, 770]. Seven [358, 689]. shalls [224]. shape [821]. shared [238, 367]. sharing [691, 138]. Should [42, 201]. side [825, 562]. side-channel [825]. signal [479, 793]. signalling [688, 263]. signals [937]. signatures [550, 394]. similarities [396]. simple [547]. simplicity [692]. simplicity-driven [692]. simulated [150]. simulating [688]. simulation [486, 894, 519, 68, 901, 11, 75, 41, 210, 160, 446]. simulator [23, 876]. Simulink [679]. single [63, 787, 797]. single-path [787, 797]. SIP [306]. SIP-ISUP [306]. SIP-ISUP/ISDN [306]. site [353]. sites [239, 354]. size [405]. Skeleton [879]. sketching [726, 492]. SL-COMP [849]. SLA [399]. Sleek [521]. Slicing [713, 709, 275, 846, 89, 41, 807, 250]. small [215]. Smart [573, 759, 777, 960, 702, 861, 523]. SMC [569]. smells [896]. SMT [915, 324, 359, 699, 388, 597, 903, 917, 952, 415]. SMT-based [915, 597]. SNIP [456]. social [564, 946]. socket [346]. Software [612, 578, 443, 707, 237, 679, 450, 866, 319, 226, 469, 324, 889, 276, 529, 839, 838, 363, 187, 561, 456, 400, 337, 246, 963, 165, 225, 580, 413, 697, 750, 749, 288, 385, 592, 454, 594, 168, 384, 757, 274, 458, 145, 140, 761, 485, 545, 256, 344, 389, 445, 222, 582, 748, 466, 81, 13, 503, 632, 453, 269, 22, 605, 257, 346]. software-defined [503]. software-intensive [592]. solution [195]. solutions [714, 885, 539]. solver [156, 317, 417, 665]. solvers [324, 359, 388, 962, 849, 415]. Solving [368, 584, 347, 180, 909, 337, 730, 912, 376, 903, 295, 421, 710]. Some [598, 667, 245, 363, 589]. sorters [106]. SOTA [772]. Sound [571, 650]. Soundness [623]. Source [736, 467, 633, 333, 503]. source-level [467]. space [63, 192, 618, 202, 35, 579, 730, 716, 267, 261, 732, 240, 364]. spaces [176, 154, 206, 312, 295]. SPARK [586]. Spatial [760, 942, 627]. Spatio [690, 962, 933]. Spatio-temporal [690, 962, 933]. Special [186, 170, 929, 165, 934, 754, 578, 325, 259, 278, 314, 233, 350, 612, 886, 781, 762, 30, 771, 819, 201, 652, 911, 419, 907, 882, 253, 292, 298, 439, 854, 269, 940, 395]. specific [436, 948, 381, 692, 672, 441, 860, 961]. Specification [72, 84, 207, 662, 836, 778, 235, 870, 156, 197, 696, 766, 68, 497, 261, 722]. specifications [141, 196, 874, 388, 635, 546, 861, 520, 306, 115, 964, 496, 302, 459, 358, 222, 710, 551, 257]. specified [858, 330]. Specifying [828, 821, 143]. speed [779, 500, 379]. SPIN [725, 727, 781, 754, 39, 45, 907, 97, 42, 44, 764, 87]. Spin-based [727]. split [691]. SPLs [747]. spontaneous [773]. stage [490]. standard [769, 768, 185]. Standardized [306]. standards [200]. standards-based [200]. State [35, 38, 575, 176, 192, 322, 202, 118, 137, 154, 206, 290, 954, 606, 267, 474, 905, 261, 732, 312, 169, 450, 528, 133, 548, 898, 676, 364, 636]. state-based [474, 548]. state-rich [636]. state-space [202]. Stateflow [460, 272]. Stateless [728]. statements [102]. states [34]. Static [796, 799, 484, 296, 91, 287, 528]. stations [88]. Statistical [429, 566, 574, 567, 555, 570, 576, 575, 794, 577, 556, 928, 571, 568, 218]. status [511]. steps [901]. stepwise [351]. still [181]. stochastic [575, 76, 574, 964, 685, 378, 523]. Store [471]. Storm [881]. strategies [735, 372, 698, 703, 742]. strategy [942, 505, 872, 685]. Stream [805]. streaming [750, 749]. streams [805, 875]. Striver [805]. stroke [358]. strong [176, 423, 2]. structural [927]. structure [138, 495, 354, 526, 784]. structured [520]. structures [662, 883, 925, 215, 151, 836]. STTT [30]. stubborn [706]. studies [377, 236]. Study [852, 763, 639, 762, 535, 118, 702, 861, 536, 497, 69, 142, 477, 315, 925, 370, 130, 784, 711, 347, 523]. study-based [315]. sub [691]. sub-task [691]. subject [901]. SubPolyhedra [425]. subscribe [162]. subway [519]. suitability [370]. suite [305, 841, 303, 455, 515]. summaries [663]. super [361]. superscalar [102, 873, 26]. supervisory [746, 719]. Support [237, 950, 293, 411, 919]. Supporting [865]. suppressed [57]. suppression [823]. surrogate [924]. Survey [543, 32, 344, 183]. survivability [801]. Swarm [783, 759, 240]. swarm-based [240]. sweep [266, 216, 281]. sweep-line [266, 216, 281]. switching [421]. SWOT [632]. Sylvan [669]. Symbiotic [846]. Symbolic [180, 323, 393, 504, 409, 117, 844, 33, 846, 202, 46, 635, 88, 524, 737, 739, 155, 182, 530, 131, 344, 286, 121, 683, 737]. Symmetric [87]. Symmetry [740, 168]. Symposium [395]. synchronization [943, 813, 489]. synchronizing [622]. Synchronous [211, 164, 897, 855]. syntactic [698]. SYNTCOMP [651]. Synthesis [661, 497, 498, 378, 823, 746, 942, 488, 313, 495, 494, 872, 581, 901, 851, 651, 491, 685, 101, 490, 493, 489, 604]. Synthesizing [656, 421, 526]. SyReNN [913]. SysML [522]. System [196, 335, 68, 823, 7, 780, 645, 643, 773, 72, 827, 189, 906, 539, 592, 285, 647, 325, 375, 671, 113, 280, 798, 642, 474, 641, 128, 140, 24, 708, 591, 560, 627, 640, 453, 688, 195, 604, 141, 374]. System-level [196, 592, 128]. System-on-chip [335]. Systematic [581, 537, 563, 95, 516, 548]. Systems [790, 959, 350, 395, 772, 965, 238, 950, 653, 357, 883, 186, 478, 667, 429, 380, 789, 886, 618, 486, 941, 408, 383, 241, 753, 777, 293, 179, 33, 234, 853, 690, 84, 743, 476, 635, 462, 566, 771, 437, 751, 361, 829, 574, 114, 650, 519, 538, 657, 943, 297, 818, 775, 749, 207, 513, 652, 733, 258, 592, 8, 480, 525, 678, 417, 214, 260, 911, 277, 902, 899, 6, 274, 477, 11, 897, 882, 832, 694, 633, 864, 867, 314, 540, 204, 373, 936, 945, 951, 890, 555, 695, 169, 445, 483, 595, 837, 807, 240]. systems [475, 878, 292, 309, 333, 815, 528, 610, 466, 956, 409, 490, 644, 947, 860, 854, 558, 940, 294, 455, 121, 605, 9, 328, 609, 693, 687, 868, 750].

TaaWS [877]. tabled [149]. tables [542]. TACAS [886, 911, 882, 854]. TACAS’13 [659]. tactic [682]. Tactical [708]. tagging [454]. Tailored [532]. targeting [843]. Tarjan [599]. task [691, 480, 280, 910]. taught [215]. taxonomy [809, 533]. TC [843]. TCP [262]. technique [662, 836, 500, 95, 596, 801]. Techniques [541, 180, 434, 35, 506, 810, 66, 707, 152, 852, 481, 11, 315, 61, 101, 502, 475, 748, 426, 328]. technologies [518]. technology [906, 47, 458, 13]. Template [493, 677, 212]. Template-based [493]. Temporal [476, 875, 690, 884, 247, 115, 678, 96, 900, 962, 933, 793, 821, 815, 676]. Temporal-logic [875]. Termination [547, 888]. Test [158, 303, 199, 195, 654, 565, 734, 91, 535, 846, 305, 338, 611, 842, 543, 68, 448, 511, 524, 115, 877, 840, 597, 11, 845, 129, 302, 124, 731, 672, 340, 300, 841, 304, 298, 610, 548, 866, 449, 605, 515, 609]. test-case [840, 597]. test-goal [841]. test-suite [841]. testable [447]. Testing [635, 513, 309, 85, 551, 759, 196, 812, 839, 838, 734, 362, 416, 332, 430, 396, 537, 608, 536, 533, 563, 306, 284, 511, 288, 648, 646, 647, 110, 677, 607, 698, 674, 315, 314, 397, 518, 534, 755, 344, 445, 130, 342, 568, 516, 815, 752, 308, 514, 866, 927, 706, 195, 301, 605, 638, 724, 609, 44, 557]. TestREx [718]. tests [482, 502, 198, 860]. textual [579]. TGV [194]. Thai [688]. their [698, 57, 312]. Theorem [47, 48, 415]. theories [603, 384]. theory [310, 774, 55, 194, 719, 215]. Thoth [162]. Thoughtful [527]. thread [782]. threaded [739, 86]. three [406, 923]. three-valued [406]. tier [361]. Tiger [374]. Tiger-MGP [841]. Time [207, 965, 7, 653, 863, 478, 405, 738, 383, 241, 33, 356, 743, 476, 635, 462, 546, 114, 730, 943, 805, 209, 290, 891, 704, 212, 474, 6, 477, 742, 937, 910, 360, 329, 558, 9, 257]. time-independent [943]. Timed [477, 723, 172, 217, 808, 171, 703, 902, 267, 905, 124, 210, 815, 644, 151, 593]. timed-arc [703]. Timed-automata [477]. TIMES [463]. Timing [459, 873, 465, 208, 107, 468, 464]. TiMo [653]. Tinker [682]. TLA [828]. together [232]. token [883]. tolerance [313, 191]. tolerant [855]. too [509]. Tool [161, 293, 237, 27, 3, 713, 383, 126, 163, 112, 132, 682, 411, 162, 958, 92, 911, 103, 279, 410, 468, 896, 921, 933, 672, 340, 369, 557, 160, 913, 19, 329, 455, 9]. tool-integration [160]. toolbox [470]. toolchain [942]. ToolDAy [410]. toolkit [94]. Tools [469, 789, 886, 214, 260, 911, 277, 882, 265, 854, 667, 715, 827, 4, 188, 400, 809, 543, 810, 14, 707, 80, 384, 899, 6, 11, 185, 626, 692, 370, 748, 956, 548, 503, 632, 334, 226, 515, 868]. toolset [387, 12, 862, 608, 291, 790]. TOOLympics [838, 847]. topics [245]. trace [654, 958, 793, 807, 898]. TRACE4CPS [958]. traces [795, 287, 658, 958, 153]. tracing [464]. Tradeoff [478]. traffic [860]. trail [167]. trajectories [544]. trajectory [98]. transactional [422]. transfer [906, 47]. transformation [953, 811, 561, 655, 258, 958, 826, 373, 353, 370, 369, 815]. transformational [550]. Transformations [249, 709, 922]. transition [158, 751, 274, 169, 528]. translating [252]. translation [148, 85, 686]. transport [690, 687]. traversal [732]. tree [435, 434]. trees [894, 138, 792, 800, 919]. trend [82]. trends [575, 511, 458, 344]. TriCore [873]. triple [813]. trustworthy [720]. TSTL [677]. TTCN [511, 299, 305, 307, 306, 513, 512, 302, 298, 309, 303, 308, 514, 515]. TTCN-2 [305]. TTCN-3 [511, 299, 307, 306, 513, 512, 302, 298, 309, 303, 308, 514, 515]. turns [871]. tutorial [569, 595, 837]. two [136, 490]. two-stage [490]. Type [164, 466]. Type-based [164, 466]. typed [550, 925]. types [121]. typestate [442]. typing [116].

UML [207, 796, 196, 213, 416, 608, 209, 277, 129, 265, 210, 211, 257]. UML-based [257]. UML/OCL [416]. unbounded [742, 570, 614]. uncertain [893]. uncertainty [926, 935]. Underapproximation [663]. understanding [41, 922]. unfolding [609]. unfolding-based [609]. unfoldings [289]. unified [620]. uniform [406, 311]. uninterpreted [613]. Unit [144, 340]. unknown [938]. unnecessary [813]. unreliable [822]. UNSAT [914]. until [570]. untimed [347]. update [728]. updating [353]. upgrade [660]. upon [247]. UPPAAL [902, 569, 10, 905]. Uppaal2k [79]. upper [217, 616]. upper-bound [616]. usability [388]. Usage [237, 649, 828, 648]. usage-based [648]. use [357, 537]. user [355, 675, 47, 5]. Using [859, 91, 106, 88, 341, 50, 626, 394, 150, 487, 402, 336, 565, 883, 764, 324, 404, 808, 56, 874, 561, 33, 158, 35, 93, 16, 462, 556, 751, 396, 861, 611, 792, 769, 842, 68, 427, 154, 284, 524, 767, 571, 43, 892, 66, 930, 480, 317, 45, 656, 48, 912, 704, 420, 496, 852, 103, 642, 274, 182, 867, 371, 73, 695, 264, 275, 222, 815, 79, 308, 421, 286, 784, 281, 688, 601, 120, 149, 44]. Utilizing [853]. UWA [355].

V&V [540]. V2 [236]. Vacuity [96, 379]. vade [19]. Validating [765, 210, 15]. Validation [764, 207, 767, 642, 129, 265, 27, 238, 336, 499, 72, 460, 137, 260, 45, 458, 119, 903, 961]. value [840]. valued [406]. VAMP [232]. VANETs [785]. variability [743, 664, 751, 579, 452, 451, 741]. variability-intensive [743]. Variable [780, 684]. variables [7]. variants [582]. Variations [668]. various [666]. VBS [476]. VDM [686]. vector [326, 69, 919]. vehicular [690]. VeriFast [584]. Verifiable [872]. Verification [802, 653, 499, 834, 54, 83, 56, 743, 929, 934, 901, 48, 263, 18, 469, 619, 621, 662, 836, 239, 357, 780, 883, 893, 717, 232, 529, 873, 12, 699, 923, 189, 588, 383, 197, 874, 181, 33, 960, 696, 709, 229, 93, 105, 573, 462, 135, 337, 138, 620, 246, 848, 861, 786, 552, 809, 519, 521, 413, 506, 112, 297, 670, 697, 520, 500, 805, 631, 891, 191, 554, 820, 737, 739, 285, 66, 317, 69, 325, 510, 419, 104, 152, 67, 40, 924, 318, 908, 107, 103, 685, 572, 185, 633, 867, 518, 440, 61]. verification [708, 108, 921, 210, 695, 97, 169, 659, 27, 183, 441, 144, 98, 431, 528, 956, 932, 248, 79, 439, 493, 19, 791, 52, 296, 589, 250, 151, 121, 339, 785, 9, 693, 604, 715, 835, 509, 819]. Verified [384, 914, 134, 939, 671]. verifier [812, 231]. verifier-based [812]. verifiers [38]. verify [587, 106, 88, 444]. Verifying [615, 892, 855, 915, 950, 205, 691, 234, 895, 828, 964, 541, 50, 912, 898, 150]. VerifyThis [848, 585, 583, 673, 584]. VeriTech [252]. VeriVANca [785]. version [511, 908, 938, 791]. very [154, 594]. VHDL [89]. via [615, 414, 661, 664, 637, 520, 280, 178, 924, 393, 534, 342, 502, 922]. viable [475]. VIATRA2 [372]. view [621, 299, 620, 47, 867]. violation [399]. violations [173, 392]. virtual [341, 364]. virtually [897]. VIS [97]. vision [632]. Visualization [958, 451, 400, 206, 866]. Visualizer [879]. VNN [923]. VNN-COMP [923]. vs [218]. VSE [51]. vulnerabilities [799, 557].

wands [590]. waves [363, 871]. WCET [467]. weak [684, 791, 2]. weak-memory [791]. weakest [414]. Web [395, 400, 877, 514, 724, 239, 355, 316, 396, 75, 315, 314, 397, 472, 353, 350, 354, 198, 352, 308, 398]. Web-based [396, 75]. weight [67]. weighted [616, 311]. Why3 [587]. Widening [268, 434]. wider [357]. will [153]. wireless [704]. within [181, 560]. Witnessing [590]. word [59]. word-level [59]. work [264]. workarounds [316, 710]. workbench [626]. workflow [699, 501, 280]. Workflows [564, 628, 625, 629]. workloads [402]. workshop [177, 29]. world [759]. Worst [114, 738, 329]. Worst-case [114, 738, 329].

X [949]. X-KLAIM [949]. Xenon [418]. XSB [16]. XSpeed [733].

Year [48]. years [923].

Z [358]. Zero [57]. Zero-suppressed [57]. Zeus [171]. zone [217, 721]. zone-based [217].

References

Cleaveland:1997:E

[1] W. Rance Cleaveland, Tiziana Margaria, and Bernhard Steffen. Editorial. International Journal on Software Tools for Technology Transfer (STTT), 1(1– 2):1–5, December 1997. CODEN ????

ISSN 1433-2779 (print), 1433-2787 (electronic).

Wolper:1997:MFW

[2] Pierre Wolper. The meaning of “formal”: from weak to strong formal methods. International Journal on Software Tools for Technology Transfer (STTT), 1(1–2):6–8, December 1997. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Steffen:1997:ETI

[3] Bernhard Steffen, Tiziana Margaria, and Volker Braun. The Electronic Tool Integration platform: concepts and design. International Journal on Software Tools for Technology Transfer (STTT), 1(1– 2):9–30, December 1997. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Braun:1997:ITE

[4] Volker Braun, Tiziana Margaria, and Carsten Weise. Integrating tools in the ETI platform. International Journal on Software Tools for Technology Transfer (STTT), 1(1–2):31–48, December 1997. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Margaria:1997:IEU

[5] Tiziana Margaria, Volker Braun, and Jürgen Kreileder. Interacting with ETI: a user session. International Journal on Software Tools for Technology Transfer (STTT), 1(1–2):49–63, December 1997. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Larsen:1997:CMR

[6] Kim G. Larsen, B. Steffen, and C. Weise. Continuous modeling of real-time and hybrid systems: from concepts to tools. International Journal on Software Tools for Technology Transfer (STTT), 1(1– 2):64–85, December 1997. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Alur:1997:RTS

[7] Rajeev Alur and Thomas A. Real-time system = discrete system + clock variables. International Journal on Software Tools for Technology Transfer (STTT), 1(1–2):86–109, December 1997. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Henzinger:1997:HMC

[8] Thomas A. Henzinger, Pei-Hsin Ho, and Howard Wong-Toi. HYTECH: a model checker for hybrid systems. International Journal on Software Tools for Technology Transfer (STTT), 1(1– 2):110–122, December 1997. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Yovine:1997:KVT

[9] Sergio Yovine. KRONOS: a verification tool for real-time systems. International Journal on Software Tools for Technology Transfer (STTT), 1(1– 2):123–133, December 1997. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Larsen:1997:UN

[10] Kim G. Larsen, Paul Pettersson, and Wang Yi. Uppaal in a nutshell. International Journal on Software Tools for Technology Transfer (STTT), 1(1– 2):134–152, December 1997. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Leblanc:1997:OSB

[11] Philippe Leblanc. OMT and SDL based techniques and tools for design, simulation and test production of distributed systems. International Journal on Software Tools for Technology Transfer (STTT), 1(1–2):153–165, December 1997. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Bozga:1997:PVA

[12] Marius Bozga, Jean-Claude Fernandez, Alain Kerbrat, and Laurent Mounier. Protocol verification with the ALDÉBARAN toolset. International Journal on Software Tools for Technology Transfer (STTT), 1(1–2):166– 183, December 1997. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Snelting:1998:PFS

[13] Gregor Snelting. Paul Feyerabend and software technology. International Journal on Software Tools for Technology Transfer (STTT), 2(1):1–5, November 1998. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Hankin:1998:PAT

[14] Chris Hankin. Program analysis tools. International Journal on Software Tools for Technology Transfer (STTT), 2(1): 6–12, November 1998. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Amtoft:1998:BAV

[15] Torben Amtoft, Hanne Riis Nielson, and Flemming Nielson. Behavior analysis for validating communication patterns. International Journal on Software Tools for Technology Transfer (STTT), 2(1): 13–28, November 1998. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Codish:1998:SBP

[16] Michael Codish, Bart Demoen, and Konstantinos Sagonas. Semantics-based program analysis for logic-based languages using XSB. International Journal on Software Tools for Technology Transfer (STTT), 2(1):29–45, November 1998. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Martin:1998:PEP

[17] Florian Martin. PAG — an efficient program analyzer generator. International Journal on Software Tools for Technology Transfer (STTT), 2(1):46– 67, November 1998. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Sighireanu:1998:VLL

[18] Mihaela Sighireanu and Radu Mateescu. Verification of the Link layer protocol of the IEEE-1394 serial bus (FireWire): an experiment with E-LOTOS. International Journal on Software Tools for Technology Transfer (STTT), 2(1): 68–88, November 1998. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Stevens:1998:VTD

[19] Perdita Stevens. A verification tool developer’s vade mecum. International Journal on Software Tools for Technology Transfer (STTT), 2(2):89–94, December 1998. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Jensen:1998:PSE

[20] Kurt Jensen. Preface by the Section Editor: Kurt Jensen. International Journal on Software Tools for Technology Transfer (STTT), 2(2):95–97, December 1998. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Kristensen:1998:PGC

[21] Lars M. Kristensen, Soren Christensen, and Kurt Jensen. The practitioner’s guide to coloured Petri nets. International Journal on Software Tools for Technology Transfer (STTT), 2(2):98– 132, December 1998. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Xu:1998:AEA

[22] Jianli Xu and Juha Kuusela. Analyzing the execution architecture of mobile phone software with colored Petri nets. International Journal on Software Tools for Technology Transfer (STTT), 2(2): 133–143, December 1998. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Gordon:1998:AMS

[23] Steven Gordon and Jonathan Billington. Analysing a missile simulator with coloured Petri nets. International Journal on Software Tools for Technology Transfer (STTT), 2(2):144–159, December 1998. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Moncelet:1998:AMS

[24] Gilles Moncelet, Søren Christensen, Hamid Demmou, Mario Paludetto, and José Porras. Analysing a mechatronic system with coloured Petri nets. International Journal on Software Tools for Technology Transfer (STTT), 2(2): 160–167, December 1998. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Wagenhals:1998:CEM

[25] Lee W. Wagenhals, Insub Shin, and Alexander H. Levis. Creating executable models of influence nets with colored Petri nets. International Journal on Software Tools for Technology Transfer (STTT), 2(2):168–181, December 1998. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Burns:1998:ASP

[26] F. P. Burns, A. M. Koelmans, and A. V. Yakovlev. Analysing superscalar processor architectures with coloured Petri nets. International Journal on Software Tools for Technology Transfer (STTT), 2(2):182–191, December 1998. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Pnueli:1998:CVT

[27] A. Pnueli, O. Shtrichman, and M. Siegel. The Code Validation Tool (CVT): Automatic verification of a compilation process. International Journal on Software Tools for Technology Transfer (STTT), 2(2):192–201, December 1998. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Friese:1998:IPO

[28] Thomas Friese, Tiziana Margaria, and Alfred Hofmann. Integrating printed and online information. International Journal on Software Tools for Technology Transfer (STTT), 2(2):202, December 1998. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Johnson:1999:WFM

[29] Steven D. Johnson. A workshop on formal methods education: held at Melbourne Florida in March 1998 [5]: an aggregation of opinions. International Journal on Software Tools for Technology Transfer (STTT), 2(3):203– 207, November 1999. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Cleaveland:1999:PMC

[30] Rance Cleaveland. Pragmatics of model checking: an STTT special section. International Journal on Software Tools for Technology Transfer (STTT), 2(3): 208–218, November 1999. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Du:1999:LMC

[31] Xiaoqun Du, Scott A. Smolka, and Rance Cleaveland. Local model checking and protocol analysis. International Journal on Software Tools for Technology Transfer (STTT), 2(3):219– 241, November 1999. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Andersen:1999:PMC

[32] Henrik Reif Andersen and Jorn Lind-Nielsen. Partial model checking of modal equations: a survey. International Journal on Software Tools for Technology Transfer (STTT), 2(3):242–259, November 1999. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Campos:1999:AVR

[33] Sérgio Vale Aguiar Campos and Edmund Clarke. Analysis and verification of real-time systems using quantitative symbolic algorithms. International Journal on Software Tools for Technology Transfer (STTT), 2(3):260– 269, November 1999. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Holzmann:1999:MAR

[34] Gerard J. Holzmann and Anuj Puri. A minimized automaton representation of reachable states. International Journal on Software Tools for Technology Transfer (STTT), 2(3):270–278, November 1999. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Clarke:1999:SSR

[35] E. M. Clarke, O. Grumberg, M. Minea, and D. Peled. State space reduction using partial order techniques. International Journal on Software Tools for Technology Transfer (STTT), 2(3):279– 287, November 1999. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Pixley:1999:MCH

[36] Carl Pixley and Vigyan Singhal. Model checking: a hardware design perspective. International Journal on Software Tools for Technology Transfer (STTT), 2(3):288–306, November 1999. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Jay:1999:PF

[37] C. Barry Jay. Programming in FISh. International Journal on Software Tools for Technology Transfer (STTT), 2(3): 307–315, November 1999. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Avrunin:2000:BFS

[38] George S. Avrunin, James C. Corbett, and Matthew B. Dwyer. Benchmarking finite-state verifiers. International Journal on Software Tools for Technology Transfer (STTT), 2(4):317–320, March 2000. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Holzmann:2000:SMC

[39] Gerard Holzmann, Eli Najm, and Ahmed Serhrouchni. SPIN model checking: an introduction. International Journal on Software Tools for Technology Transfer (STTT), 2(4):321–327, March 2000. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Kesten:2000:CDA

[40] Yonit Kesten and Amir Pnueli. Control and data abstraction: the cornerstones of practical formal verification. International Journal on Software Tools for Technology Transfer (STTT), 2(4): 328–342, March 2000. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Millett:2000:ISP

[41] Lynette I. Millett and Tim Teitelbaum. Issues in slicing PROMELA and its applications to model checking, protocol understanding, and simulation. International Journal on Software Tools for Technology Transfer (STTT), 2(4): 343–349, March 2000. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Visser:2000:PCM

[42] Willem Visser and Howard Barringer. Practical CTL^{∗} model checking: Should SPIN be extended? International Journal on Software Tools for Technology Transfer (STTT), 2(4):350–365, March 2000. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Havelund:2000:MCJ

[43] Klaus Havelund and Thomas Pressburger. Model checking JAVA programs using JAVA PathFinder. International Journal on Software Tools for Technology Transfer (STTT), 2(4):366–381, March 2000. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

deVries:2000:FCT

[44] René G. de Vries and Jan Tretmans. Onthe-fly conformance testing using SPIN. International Journal on Software Tools for Technology Transfer (STTT), 2(4): 382–393, March 2000. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Kamel:2000:FVG

[45] Moataz Kamel and Stefan Leue. Formalization and validation of the General Inter-ORB Protocol (GIOP) using PROMELA and SPIN. International Journal on Software Tools for Technology Transfer (STTT), 2(4):394–409, March 2000. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Cimatti:2000:NNS

[46] Alessandro Cimatti, Edmund Clarke, Fausto Giunchiglia, and Marco Roveri. NUSMV: a new symbolic model checker. International Journal on Software Tools for Technology Transfer (STTT), 2(4): 410–425, March 2000. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Giunchiglia:2000:TPT

[47] Fausto Giunchiglia and Paolo Traverso. Theorem proving in technology transfer: the user’s point of view. International Journal on Software Tools for Technology Transfer (STTT), 3(1): 1–12, September 2000. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Kaufmann:2000:VYC

[48] Matt Kaufmann. Verification of Year 2000 conversion rules using the ACL2 theorem prover. International Journal on Software Tools for Technology Transfer (STTT), 3(1):13–19, September 2000. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

DiVito:2000:HAP

[49] Ben L. Di Vito. High-automation proofs for properties of requirements models. International Journal on Software Tools for Technology Transfer (STTT), 3(1): 20–31, September 2000. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Kapur:2000:UIP

[50] Deepak Kapur and Mahadevan Subramaniam. Using an induction prover for verifying arithmetic circuits. International Journal on Software Tools for Technology Transfer (STTT), 3(1): 32–65, September 2000. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Autexier:2000:VFM

[51] Serge Autexier, Dieter Hutter, Bruno Langenstein, Heiko Mantel, Georg Rock, Axel Schairer, Werner Stephan, Roland Vogt, and Andreas Wolpers. VSE: formal methods meet industrial needs. International Journal on Software Tools for Technology Transfer (STTT), 3(1): 66–77, September 2000. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Traverso:2000:MRV

[52] Paolo Traverso and Piergiorgio Bertoli. Mechanized result verification: an industrial application. International Journal on Software Tools for Technology Transfer (STTT), 3(1):78–92, September 2000. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Aiken:2000:DRR

[53] Alexander Aiken, Manuel Fähndrich, and Zhendong Su. Detecting races in Relay Ladder Logic programs. International Journal on Software Tools for Technology Transfer (STTT), 3(1):93– 105, September 2000. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Brinksma:2001:VE

[54] Ed Brinksma. Verification is experimentation! International Journal on Software Tools for Technology Transfer (STTT), 3(2):107–111, May 2001. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Drechsler:2001:BDD

[55] Rolf Drechsler and Detlef Sieling. Binary decision diagrams in theory and practice. International Journal on Software Tools for Technology Transfer (STTT), 3(2):112–136, May 2001. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Bryant:2001:VAC

[56] Randal E. Bryant and Yirng-An Chen. Verification of arithmetic circuits using binary moment diagrams. International Journal on Software Tools for Technology Transfer (STTT), 3(2):137–155, May 2001. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Minato:2001:ZSB

[57] Shin ichi Minato. Zero-suppressed BDDs and their applications. International Journal on Software Tools for Technology Transfer (STTT), 3(2):156–170, May 2001. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Somenzi:2001:EMD

[58] Fabio Somenzi. Efficient manipulation of decision diagrams. International Journal on Software Tools for Technology Transfer (STTT), 3(2):171–181, May 2001. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Horeth:2001:WLG

[59] Stefan Höreth. A word-level graph manipulation package. International Journal on Software Tools for Technology Transfer (STTT), 3(2):182–192, May 2001. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Harlow:2001:DEE

[60] Justin E. Harlow III and Franc Brglez. Design of experiments and evaluation of BDD ordering heuristics. International Journal on Software Tools for Technology Transfer (STTT), 3(2):193–206, May 2001. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Mohnke:2001:ABB

[61] Janett Mohnke, Paul Molitor, and Sharad Malik. Application of BDDs in Boolean matching techniques for formal logic combinational verification. International Journal on Software Tools for Technology Transfer (STTT), 3(2): 207–216, May 2001. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Massingill:2001:PPP

[62] Berna L. Massingill, Timothy G. Mattson, and Beverly A. Sanders. Parallel programming with a pattern language ∗. International Journal on Software Tools for Technology Transfer (STTT), 3(2):217–234, May 2001. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Bartoli:2001:ACM

[63] Alberto Bartoli, Gianluca Dini, and Lanfranco Lopriore. Applicationcontrolled memory management in a single address space environment. International Journal on Software Tools for Technology Transfer (STTT), 3(2): 235–245, May 2001. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Cleaveland:2001:PSE

[64] Rance Cleaveland. Preface by the section editor. International Journal on Software Tools for Technology Transfer (STTT), 3(3):247–249, August 2001.

CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Delzanno:2001:CBD

[65] Giorgio Delzanno and Andreas Podelski. Constraint-based deductive model checking. International Journal on Software Tools for Technology Transfer (STTT), 3(3):250–270, August 2001. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Hirschkoff:2001:BVU

[66] Daniel Hirschkoff. Bisimulation verification using the up to techniques. International Journal on Software Tools for Technology Transfer (STTT), 3(3): 271–285, August 2001. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Kern:2001:LWF

[67] Christoph Kern, Tarik Ono-Tesfaye, and Mark R. Greenstreet. A light-weight framework for hardware verification. International Journal on Software Tools for Technology Transfer (STTT), 3(3): 286–313, August 2001. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Garavel:2001:SDC

[68] Hubert Garavel, César Viho, and Massimo Zendri. System design of a CC-NUMA multiprocessor architecture using formal specification, model-checking, co-simulation, and test generation. International Journal on Software Tools for Technology Transfer (STTT), 3(3): 314–331, August 2001. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Huisman:2001:CSC

[69] Marieke Huisman, Bart Jacobs, and Joachim van den Berg. A case study in class library verification: Java’s vector class. International Journal on Software Tools for Technology Transfer (STTT), 3(3):332–352, August 2001. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Lindahl:2001:FDA

[70] Magnus Lindahl, Paul Pettersson, and Wang Yi. Formal design and analysis of a gear controller. International Journal on Software Tools for Technology Transfer (STTT), 3(3):353–368, August 2001. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Jensen:2001:PSE

[71] Kurt Jensen. Preface by the section editor. International Journal on Software Tools for Technology Transfer (STTT), 3(4):369–371, September 2001. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Berthelot:2001:SVC

[72] Gérard Berthelot and Laure Petrucci. Specification and validation of a concurrent system: an educational project. International Journal on Software Tools for Technology Transfer (STTT), 3(4): 372–381, September 2001. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Ojala:2001:MAD

[73] Leo Ojala, Nisse Husberg, and Teemu Tynjälä. Modelling and analysing a distributed dynamic channel allocation algorithm for mobile computing using high-level net methods. International Journal on Software Tools for Technology Transfer (STTT), 3(4):382– 393, September 2001. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Genrich:2001:EPN

[74] Hartmann Genrich, Robert Küffner, and Klaus Voss. Executable Petri net models for the analysis of metabolic pathways. International Journal on Software Tools for Technology Transfer (STTT), 3(4): 394–404, September 2001. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Lindstrom:2001:WBI

[75] Bo Lindstrom. Web-based interfaces for simulation of coloured Petri net models. International Journal on Software Tools for Technology Transfer (STTT), 3(4): 405–416, September 2001. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Bernardi:2001:ICS

[76] S. Bernardi, S. Donatelli, and A. Horváth. Implementing compositionality for stochastic Petri nets. International Journal on Software Tools for Technology Transfer (STTT), 3(4):417–430, September 2001. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Anlauff:2001:GAN

[77] Matthias Anlauff, Samarjit Chakraborty, Philipp W. Kutter, Alfonso Pierantonio, and Lothar Thiele. Generating an action notation environment from Montages descriptions. International Journal on Software Tools for Technology Transfer (STTT), 3(4):431–455, September 2001. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Aizman:2001:EC

[78] Alex Aizman. Easy concurrency. International Journal on Software Tools for Technology Transfer (STTT), 3(4): 456–468, September 2001. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Simons:2001:MVI

[79] David P. L. Simons and Mariëlle I. A. Stoelinga. Mechanical verification of the IEEE 1394a root contention protocol using Uppaal2k. International Journal on Software Tools for Technology Transfer (STTT), 3(4):469–485, September 2001. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Kindler:2001:PNK

[80] Ekkart Kindler and Michael Weber. The Petri Net Kernel: an infrastructure for building Petri net tools. International Journal on Software Tools for Technology Transfer (STTT), 3(4):486– 497, September 2001. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Shaw:2002:WMG

[81] Mary Shaw. What makes good research in software engineering? International Journal on Software Tools for Technology Transfer (STTT), 4(1):1–7, October 2002. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Havelund:2002:PMC

[82] Klaus Havelund and Willem Visser. Program model checking as a new trend. International Journal on Software Tools for Technology Transfer (STTT), 4(1): 8–20, October 2002. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Brinksma:2002:VOP

[83] Ed Brinksma, Angelika Mader, and Ansgar Fehnker. Verification and optimization of a PLC control schedule. International Journal on Software Tools for Technology Transfer (STTT), 4(1): 21–33, October 2002. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Corbett:2002:ECP

[84] James C. Corbett, Matthew B. Dwyer, John Hatcliff, and Robby. Expressing checkable properties of dynamic systems: the Bandera Specification Language. International Journal on Software Tools for Technology Transfer (STTT), 4(1):34–56, October 2002. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Tauriainen:2002:TLF

[85] Heikki Tauriainen and Keijo Heljanko. Testing LTL formula translation into Büchi automata. International Journal on Software Tools for Technology Transfer (STTT), 4(1):57–70, October 2002. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Stoller:2002:MCM

[86] Scott D. Stoller. Model-checking multithreaded distributed Java programs. International Journal on Software Tools for Technology Transfer (STTT), 4(1): 71–91, October 2002. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Bosnacki:2002:SS

[87] Dragan Bošnački, Dennis Dams, and Leszek Holenderski. Symmetric Spin. International Journal on Software Tools for Technology Transfer (STTT), 4(1): 92–106, October 2002. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Eisner:2002:USC

[88] Cindy Eisner. Using symbolic CTL model checking to verify the railway stations of Hoorn-Kersenboogerd and Heerhugowaard. International Journal on Software Tools for Technology Transfer (STTT), 4(1):107–124, October 2002. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Clarke:2002:PSV

[89] E. M. Clarke, M. Fujita, S. P. Rajan, T. Reps, S. Shankar, and T. Teitelbaum. Program slicing for VHDL. International Journal on Software Tools for Technology Transfer (STTT), 4(1): 125–137, October 2002. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Graf:2003:PSE

[90] Susanne Graf. Preface by the section editor. International Journal on Software Tools for Technology Transfer (STTT), 4(2):139–141, February 2003. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Bozga:2003:USA

[91] Marius Bozga, Jean-Claude Fernandez, and Lucian Ghirvu. Using static analysis to improve automatic test generation. International Journal on Software Tools for Technology Transfer (STTT), 4(2):142–152, February 2003. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Hermanns:2003:TMC

[92] Holger Hermanns, Joost-Pieter Katoen, Joachim Meyer-Kayser, and Markus Siegle. A tool for model-checking Markov chains. International Journal on Software Tools for Technology Transfer (STTT), 4(2):153–172, February 2003. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Clarke:2003:EVS

[93] Edmund Clarke, Somesh Jha, and Will Marrero. Efficient verification of security protocols using partial-order reductions. International Journal on Software Tools for Technology Transfer (STTT), 4(2):173–188, February 2003. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Dennis:2003:PT

[94] Louise A. Dennis, Graham Collins, Michael Norrish, Richard J. Boulton, Konrad Slind, and Thomas F. Melham. The PROSPER toolkit. International Journal on Software Tools for Technology Transfer (STTT), 4(2): 189–210, February 2003. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Johnson:2003:SIT

[95] Steven D. Johnson, Yanhong A. Liu, and Yuchen Zhang. A systematic incrementalization technique and its application to hardware design. International Journal on Software Tools for Technology Transfer (STTT), 4(2):211–223, February 2003. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Kupferman:2003:VDT

[96] Orna Kupferman and Moshe Y. Vardi. Vacuity detection in temporal model checking. International Journal on Software Tools for Technology Transfer (STTT), 4(2):224–233, February 2003. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Peng:2003:CSV

[97] Hong Peng, Sofiène Tahar, and Ferhat Khendek. Comparison of SPIN and VIS for protocol verification. International Journal on Software Tools for Technology Transfer (STTT), 4(2): 234–245, February 2003. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Ruys:2003:MVT

[98] Theo C. Ruys and Ed Brinksma. Managing the verification trajectory. International Journal on Software Tools for Technology Transfer (STTT), 4(2): 246–259, February 2003. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Wing:2003:PA

[99] Jeannette M. Wing. Platitudes and attitudes. International Journal on Software Tools for Technology Transfer (STTT), 4(3):261–265, May 2003. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Margaria:2003:PSE

[100] Tiziana Margaria. Preface by the section editor. International Journal on Software Tools for Technology Transfer (STTT), 4(3):266–270, May 2003. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Mycroft:2003:HLT

[101] Alan Mycroft and Richard Sharp. Higher-level techniques for hardware description and synthesis. International Journal on Software Tools for Technology Transfer (STTT), 4(3):271–297, May 2003. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Aagaard:2003:FSM

[102] Mark D. Aagaard, Byron Cook, Nancy A. Day, and Robert B. Jones. A framework for superscalar microprocessor correctness statements. International Journal on Software Tools for Technology Transfer (STTT), 4(3): 298–312, May 2003. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Kort:2003:HFV

[103] Skander Kort, Sofiène Tahar, and Paul Curzon. Hierarchical formal verification using a hybrid tool. International Journal on Software Tools for Technology Transfer (STTT), 4(3):313–322, May 2003. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Kaivola:2003:PEL

[104] Roope Kaivola and Katherine Kohatsu.

Proof engineering in the large: formal verification of Pentium(R)4 floatingpoint divider. International Journal on Software Tools for Technology Transfer (STTT), 4(3):323–334, May 2003. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Copty:2003:EDF

[105] Fady Copty, Amitai Irron, Osnat Weissberg, Nathan Kropp, and Gila Kamhi. Efficient debugging in a formal verification environment. International Journal on Software Tools for Technology Transfer (STTT), 4(3):335–348, May 2003. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Claessen:2003:ULD

[106] Koen Claessen, Mary Sheeran, and Satnam Singh. Using Lava to design and verify recursive and periodic sorters. International Journal on Software Tools for Technology Transfer (STTT), 4(3): 349–358, May 2003. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Kong:2003:RBF

[107] Xiaohua Kong, Radu Negulescu, and Larry Weidong Ying. Refinementbased formal verification with heterogeneous timing. International Journal on Software Tools for Technology Transfer (STTT), 4(3):359–370, May 2003. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Munoz:2003:FVC

[108] César Muñoz, Víctor Carreño, Gilles Dowek, and Ricky Butler. Formal verification of conflict detection algorithms. International Journal on Software Tools for Technology Transfer (STTT), 4(3): 371–380, May 2003. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

ElGuemhioui:2003:FDO

[109] Karim El Guemhioui. A framework for distributing object-oriented designs. International Journal on Software Tools for Technology Transfer (STTT), 4(3): 381–396, May 2003. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Hogrefe:2003:MIP

[110] Dieter Hogrefe. Main issues in protocol testing. International Journal on Software Tools for Technology Transfer (STTT), 4(4):397–400, August 2003. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Jonsson:2003:PSE

[111] Bengt Jonsson and Konstantinos Sagonas. Preface by the section editors. International Journal on Software Tools for Technology Transfer (STTT), 4(4): 401–404, August 2003. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Fredlund:2003:VTE

[112] Lars-Åke Fredlund, Dilian Gurov, Thomas Noll, Mads Dam, Thomas Arts, and Gennady Chugunov. A verification tool for ERLANG. International Journal on Software Tools for Technology Transfer (STTT), 4(4):405–420, August 2003. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Johansson:2003:DHS

[113] Erik Johansson, Mikael Pettersson, Konstantinos Sagonas, and Thomas Lindgren. The development of the HiPE system: design and experience report. International Journal on Software Tools for Technology Transfer (STTT), 4(4): 421–436, August 2003. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Engblom:2003:WCE

[114] Jakob Engblom, Andreas Ermedahl, Mikael Sjödin, Jan Gustafsson, and Hans Hansson. Worst-case executiontime analysis for embedded real-time systems. International Journal on Software Tools for Technology Transfer (STTT), 4(4):437–455, August 2003. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Haakansson:2003:GOT

[115] John Håkansson, Bengt Jonsson, and Ola Lundqvist. Generating online test oracles from temporal logic specifications. International Journal on Software Tools for Technology Transfer (STTT), 4(4):456–471, August 2003. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Debbabi:2003:ST

[116] Mourad Debbabi, Nancy Durgin, Mohamed Mejri, and John C. Mitchell. Security by typing. International Journal on Software Tools for Technology Transfer (STTT), 4(4):472–495, August 2003. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Ben-David:2003:SDF

[117] Shoham Ben-David, Orna Grumberg, Tamir Heyman, and Assaf Schuster. Scalable distributed on-the-fly symbolic model checking. International Journal on Software Tools for Technology Transfer (STTT), 4(4):496–504, August 2003. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Dong:2003:FLG

[118] Yifei Dong, Xiaoqun Du, Gerard J. Holzmann, and Scott A. Smolka. Fighting livelock in the GNU i-protocol: a case study in explicit-state model checking. International Journal on Software Tools for Technology Transfer (STTT), 4(4):505–528, August 2003. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Margaria:2003:IPS

[119] Tiziana Margaria and Wang Yi. Introductory paper: scalability aspects of validation. International Journal on Software Tools for Technology Transfer (STTT), 5(1):1–3, November 2003. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Williams:2003:SCU

[120] Poul F. Williams, Henrik R. Andersen, and Henrik Hulgaard. Satisfiability checking using Boolean Expression Diagrams. International Journal on Software Tools for Technology Transfer (STTT), 5(1):4–14, November 2003. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Yavuz-Kahveci:2003:SMA

[121] Tuba Yavuz-Kahveci and Tevfik Bultan. A symbolic manipulator for automated verification of reactive systems with heterogeneous data types. International Journal on Software Tools for Technology Transfer (STTT), 5(1): 15–33, November 2003. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Pasareanu:2003:FFA

[122] Corina S. Păsăreanu, Matthew B.

Dwyer, and Willem Visser. Finding feasible abstract counter-examples. International Journal on Software Tools for Technology Transfer (STTT), 5(1): 34–48, November 2003. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Ball:2003:BCA

[123] Thomas Ball, Andreas Podelski, and Sriram K. Rajamani. Boolean and Cartesian abstraction for model checking C programs. International Journal on Software Tools for Technology Transfer (STTT), 5(1):49–58, November 2003. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Nielsen:2003:ATG

[124] Brian Nielsen and Arne Skou. Automated test generation from timed automata. International Journal on Software Tools for Technology Transfer (STTT), 5(1):59–77, November 2003. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Gunter:2003:CMS

[125] Elsa L. Gunter, Anca Muscholl, and Doron Peled. Compositional message sequence charts. International Journal on Software Tools for Technology Transfer (STTT), 5(1):78–89, November 2003.

CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Chevalley:2003:MAT

[126] P. Chevalley and P. Thévenod-Fosse. A mutation analysis tool for Java programs. International Journal on Software Tools for Technology Transfer (STTT), 5(1):90–103, November 2003. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

vandePol:2004:IP

[127] Jaco van de Pol. Introductory paper. International Journal on Software Tools for Technology Transfer (STTT), 5(2– 3):105–106, March 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Margaria:2004:LCG

[128] Tiziana Margaria and Bernhard Steffen. Lightweight coarse-grained coordination: a scalable system-level approach. International Journal on Software Tools for Technology Transfer (STTT), 5(2–3):107–123, March 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Lugato:2004:VAT

[129] David Lugato, Céline Bigot, Yannick Valot, Jean-Pierre Gallois, Sébastien Gérard, and François Terrier. Validation and automatic test generation on UML models: the AGATHA approach. International Journal on Software Tools for Technology Transfer (STTT), 5(2– 3):124–139, March 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Pretschner:2004:MBT

[130] A. Pretschner, O. Slotosch, E. Aiglstorfer, and S. Kriebel. Model-based testing for real: The inhouse card case study. International Journal on Software Tools for Technology Transfer (STTT), 5(2– 3):140–157, March 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Pace:2004:CEG

[131] Gordon Pace, Nicolas Halbwachs, and Pascal Raymond. Counter-example generation in symbolic abstract modelchecking. International Journal on Software Tools for Technology Transfer (STTT), 5(2–3):158–164, March 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Gallardo:2004:ATA

[132] María del Mar Gallardo, Jesús Martínez, Pedro Merino, and Ernesto Pimentel. aSPIN: a tool for abstract model checking. International Journal on Software Tools for Technology Transfer (STTT), 5(2–3):165–184, March 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Schuppan:2004:ERF

[133] Viktor Schuppan and Armin Biere. Efficient reduction of finite state model checking to reachability analysis. International Journal on Software Tools for Technology Transfer (STTT), 5(2– 3):185–204, March 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Arts:2004:DVE

[134] Thomas Arts, Clara Benac Earle, and John Derrick. Development of a verified Erlang program for resource locking. International Journal on Software Tools for Technology Transfer (STTT), 5(2–3):205–220, March 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Daws:2004:AVI

[135] Conrado Daws, Marta Kwiatkowska, and Gethin Norman. Automatic verification of the IEEE 1394 root contention protocol with KRONOS and PRISM. International Journal on Software Tools for Technology Transfer (STTT), 5(2– 3):221–236, March 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Boldo:2004:PTC

[136] Sylvie Boldo and Marc Daumas. Properties of two’s complement floating point notations. International Journal on Software Tools for Technology Transfer (STTT), 5(2–3):237–246, March 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//perso.ens-lyon.fr/marc.daumas/
SoftArith/BolDau04a.pdf.

Edelkamp:2004:DES

[137] Stefan Edelkamp, Stefan Leue, and Alberto Lluch-Lafuente. Directed explicitstate model checking in the validation of communication protocols. International Journal on Software Tools for Technology Transfer (STTT), 5(2–3):247–267, March 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Delzanno:2004:CST

[138] Giorgio Delzanno, Jean-François Raskin, and Laurent Van Begin. Covering sharing trees: a compact data structure for parameterized verification. International Journal on Software Tools for Technology Transfer (STTT), 5(2–3): 268–297, March 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Iyer:2004:IP

[139] S. Purushothaman Iyer, David Hislop, Paul L. Jones, Jaime Lee, Frederick Pearce, and Stephen Van Albert. Introductory paper. International Journal on Software Tools for Technology Transfer (STTT), 5(4):299–300, May 2004. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Martin:2004:FMS

[140] John C. Martin. Formal methods software engineering for the CARA system. International Journal on Software Tools for Technology Transfer (STTT), 5(4):301–307, May 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Alur:2004:FSA

[141] Rajeev Alur, David Arney, Elsa L.

Gunter, Insup Lee, Jaime Lee, Wonhong Nam, Frederick Pearce, Steve Van Albert, and Jiaxiang Zhou. Formal specifications and analysis of the computer-assisted resuscitation algorithm (CARA) Infusion Pump Control System. International Journal on Software Tools for Technology Transfer (STTT), 5(4):308–319, May 2004. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Jetley:2004:CSA

[142] Raoul Praful Jetley, Cohan Carlos, and S. Purushothaman Iyer. A case study on applying formal methods to medical devices: computer-aided resuscitation algorithm. International Journal on Software Tools for Technology Transfer (STTT), 5(4):320–330, May 2004. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Stark:2004:FSC

[143] Eugene W. Stark. Formally specifying CARA in Java. International Journal on Software Tools for Technology Transfer (STTT), 5(4):331–350, May 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Ray:2004:UVC

[144] Arnab Ray and Rance Cleaveland. Unit verification: the CARA experience. International Journal on Software Tools for Technology Transfer (STTT), 5(4): 351–369, May 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Luqi:2004:RDB

[145] Luqi, Z. Guan, V. Berzins, L. Zhang, D. Floodeen, V. Coskun, J. Puett, and M. Brown. Requirements-document-based prototyping of CARA software. International Journal on Software Tools for Technology Transfer (STTT), 5(4): 370–390, May 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Zuck:2004:PSE

[146] Lenore Zuck, Paul Attie, and Agostino Cortesi. Preface by the section editors. International Journal on Software Tools for Technology Transfer (STTT), 6(1):1–3, July 2004. CODEN ????

ISSN 1433-2779 (print), 1433-2787 (electronic).

Hungar:2004:BBM

[147] Hardi Hungar and Bernhard Steffen.

Behavior-based model construction. International Journal on Software Tools for Technology Transfer (STTT), 6 (1):4–14, July 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Rival:2004:CCA

[148] Xavier Rival. Certification of compiled assembly code by invariant translation. International Journal on Software Tools for Technology Transfer (STTT), 6(1):15–37, July 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Yang:2004:LEC

[149] Ping Yang, C. R. Ramakrishnan, and Scott A. Smolka. A logical encoding of the π-calculus: model checking mobile processes using tabled resolution. International Journal on Software Tools for Technology Transfer (STTT), 6(1):38– 66, July 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Win:2004:USE

[150] Toh Ne Win, Michael D. Ernst, Stephen J. Garland, Dilsun Kimathrlimath, and Nancy A. Lynch. Using simulated execution in verifying distributed algorithms. International Journal on Software Tools for Technology Transfer (STTT), 6(1):67–76, July 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Wang:2004:EVT

[151] Farn Wang. Efficient verification of timed automata with BDD-like data structures. International Journal on Software Tools for Technology Transfer (STTT), 6(1):77–97, July 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Katoen:2004:GEI

[152] Joost-Pieter Katoen and Perdita Stevens. Guest editors’ introduction: Advancements and extensions of verification techniques. International Journal on Software Tools for Technology Transfer (STTT), 6(2):99–101, August 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Jin:2004:FFW

[153] HoonSang Jin, Kavita Ravi, and Fabio Somenzi. Fate and free will in error traces. International Journal on Software Tools for Technology Transfer (STTT), 6(2):102–116, August 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Godefroid:2004:EVL

[154] Patrice Godefroid and Sarfraz Khurshid. Exploring very large state spaces using genetic algorithms. International Journal on Software Tools for Technology Transfer (STTT), 6(2):117–127, August 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Kwiatkowska:2004:PSM

[155] Marta Kwiatkowska, Gethin Norman, and David Parker. Probabilistic symbolic model checking with PRISM: a hybrid approach. International Journal on Software Tools for Technology Transfer (STTT), 6(2):128–142, August 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Bouquet:2004:CBC

[156] Fabrice Bouquet, Bruno Legeard, and Fabien Peureux. CLPS-B — a constraint solver to animate a B specification. International Journal on Software Tools for Technology Transfer (STTT), 6(2): 143–157, August 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Havelund:2004:EMS

[157] Klaus Havelund and Grigore Roşu. Efficient monitoring of safety properties. International Journal on Software Tools for Technology Transfer (STTT), 6(2): 158–173, August 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Campos:2004:TSG

[158] Sérgio Campos, Orna Grumberg, Karen Yorav, and Copty Fady. Test sequence generation and model checking using dynamic transition relations. International Journal on Software Tools for Technology Transfer (STTT), 6(2):174–182, August 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Dorr:2004:I

[159] Heiko Dörr and Andy Schürr. Introduction. International Journal on Software Tools for Technology Transfer (STTT), 6(3):183–185, August 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Schopfer:2004:CTI

[160] G. Schopfer, A. Yang, L. von Wedel, and W. Marquardt. CHEOPS: a toolintegration platform for chemical process modelling and simulation. International Journal on Software Tools for Technology Transfer (STTT), 6(3): 186–202, August 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Burmester:2004:TIM

[161] Sven Burmester, Holger Giese, Jörg Niere, Matthias Tichy, Jörg P. Wadsack, Robert Wagner, Lothar Wendehals, and Albert Zündorf. Tool integration at the meta-model level: the Fujaba approach. International Journal on Software Tools for Technology Transfer (STTT), 6(3): 203–218, August 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Hansen:2004:TPS

[162] Klaus Marius Hansen. Thoth: a publish/subscribe architecture for peer-to-peer tool integration. International Journal on Software Tools for Technology Transfer (STTT), 6(3):219–230, August 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Corradini:2004:ABA

[163] Flavio Corradini, Leonardo Mariani, and Emanuela Merelli. An agentbased approach to tool integration. International Journal on Software Tools for Technology Transfer (STTT), 6(3): 231–244, August 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Colaco:2004:TBI

[164] Jean-Louis Colaço and Marc Pouzet.

Type-based initialization analysis of a synchronous dataflow language. International Journal on Software Tools for Technology Transfer (STTT), 6(3): 245–255, August 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Dwyer:2004:SSA

[165] Matthew Dwyer and Stefan Leue. Special section on the algorithmics of software model checking: Introductory paper. International Journal on Software Tools for Technology Transfer (STTT), 6(4):257–259, August 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=6&issue=4&spage=257.

Groce:2004:HMC

[166] Alex Groce and Willem Visser. Heuristics for model checking Java programs. International Journal on Software Tools for Technology Transfer (STTT), 6(4):260–276, August 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=6&issue=4&spage=260.

Edelkamp:2004:POR

[167] Stefan Edelkamp, Stefan Leue, and Alberto Lluch-Lafuente. Partial-order reduction and trail improvement in directed model checking. International Journal on Software Tools for Technology Transfer (STTT), 6(4):277–301, August 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1433-2779&volume=6&issue=4&spage=
277.

Iosif:2004:SRM

[168] Radu Iosif. Symmetry reductions for model checking of concurrent dynamic software. International Journal on Software Tools for Technology Transfer (STTT), 6(4):302–319, August 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=6&issue=4&spage=302.

Penna:2004:ETL

[169] Giuseppe Della Penna, Benedetto Intrigila, Igor Melatti, Enrico Tronci, and Marisa Venturini Zilli. Exploiting transition locality in automatic verification of finite-state concurrent systems. International Journal on Software Tools for Technology Transfer (STTT), 6(4):320–341, August 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=6&issue=4&spage=320.

Brim:2005:SSP

[170] Luboš Brim and Orna Grumberg. Special section on parallel and distributed model checking: Introductory paper. International Journal on Software Tools for Technology Transfer (STTT), 7(1):1–3, February 2005. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=7&issue=1&spage=1.

Braberman:2005:IDT

[171] Víctor Braberman, Alfredo Olivero, and Fernando Schapachnik. Issues in distributed timed model checking: Building Zeus. International Journal on Software Tools for Technology Transfer (STTT), 7(1):4–18, February 2005. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=7&issue=1&spage=4.

Behrmann:2005:DRA

[172] Gerd Behrmann. Distributed reachability analysis in timed automata. International Journal on Software Tools for Technology Transfer (STTT), 7 (1):19–30, February 2005. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=7&issue=1&spage=19.

Jones:2005:PSL

[173] Michael D. Jones and Jacob Sorber. Parallel search for LTL violations. International Journal on Software Tools for Technology Transfer (STTT), 7(1):31–42, February 2005. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=7&issue=1&spage=31.

Bell:2005:SDM

[174] Alexander Bell and Boudewijn R. Haverkort. Sequential and distributed model checking of Petri nets. International Journal on Software Tools for Technology Transfer (STTT), 7 (1):43–60, February 2005. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=7&issue=1&spage=43.

Brim:2005:ABD

[175] Luboš Brim, Karen Yorav, and Jitka Žídková. Assumption-based distribution of CTL model checking. International Journal on Software Tools for Technology Transfer (STTT), 7 (1):61–73, February 2005. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=7&issue=1&spage=61.

Blom:2005:DAS

[176] Stefan Blom and Simona Orzan. A distributed algorithm for strong bisimulation reduction of state spaces. International Journal on Software Tools for Technology Transfer (STTT), 7 (1):74–86, February 2005. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=7&issue=1&spage=74.

Biere:2005:IPH

[177] Armin Biere and Ofer Strichman. Introductory paper: Highlights from the BMC’03 workshop and more. International Journal on Software Tools for Technology Transfer (STTT), 7 (2):87–88, April 2005. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=7&issue=2&spage=87.

Jussila:2005:BFD

[178] Toni Jussila, Keijo Heljanko, and Ilkka Niemelä. BMC via on-the-fly determinization. International Journal on Software Tools for Technology Transfer (STTT), 7(2):89–101, April 2005. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=7&issue=2&spage=89.

Cabodi:2005:BBF

[179] Gianpiero Cabodi, Alex Kondratyev, Luciano Lavagno, Sergio Nocco, Stefano Quer, and Yosinori Watanabe. A BMC-based formulation for the scheduling problem of hardware systems. International Journal on Software Tools for Technology Transfer (STTT), 7(2):102–117, April 2005. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=7&issue=2&spage=102.

Alur:2005:SCT

[180] Rajeev Alur, P. Madhusudan, and Wonhong Nam. Symbolic computational techniques for solving games. International Journal on Software Tools for Technology Transfer (STTT), 7 (2):118–128, April 2005. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=7&issue=2&spage=118.

Cabodi:2005:BSA

[181] Gianpiero Cabodi, Sergio Nocco, and Stefano Quer. Are BDDs still alive within sequential verification? International Journal on Software Tools for Technology Transfer (STTT), 7 (2):129–142, April 2005. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=7&issue=2&spage=129.

Li:2005:ARS

[182] Bing Li, Chao Wang, and Fabio Somenzi. Abstraction refinement in symbolic model checking using satisfiability as the only decision procedure. International Journal on Software Tools for Technology Transfer (STTT), 7(2):143–155, April 2005. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=7&issue=2&spage=143.

Prasad:2005:SRA

[183] Mukul R. Prasad, Armin Biere, and Aarti Gupta. A survey of recent advances in SAT-based formal verification. International Journal on Software Tools for Technology Transfer (STTT), 7(2):156–173, April 2005. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=7&issue=2&spage=156.

Clarke:2005:CCB

[184] Edmund Clarke, Daniel Kroening, Joël Ouaknine, and Ofer Strichman. Computational challenges in bounded model checking. International Journal on Software Tools for Technology Transfer (STTT), 7(2):174–183, April 2005. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=7&issue=2&spage=174.

Leucker:2005:FPL

[185] Martin Leucker, Thomas Noll, Perdita Stevens, and Michael Weber. Functional programming languages for verification tools: a comparison of Standard ML and Haskell. International Journal on Software Tools for Technology Transfer (STTT), 7(2):184–194, April 2005. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=7&issue=2&spage=184.

Arts:2005:SSF

[186] Thomas Arts and Jaco van de Pol.

Special section on formal methods for industrial critical systems: Introductory paper. International Journal on Software Tools for Technology Transfer (STTT), 7(3):195–196, June 2005. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=7&issue=3&spage=195.

Bubel:2005:IIF

[187] Richard Bubel and Reiner Hähnle. Integration of informal and formal development of object-oriented safety-critical software. International Journal on Software Tools for Technology Transfer (STTT), 7(3):197–211, June 2005. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=7&issue=3&spage=197.

Burdy:2005:OJT

[188] Lilian Burdy, Yoonsik Cheon, David R. Cok, Michael D. Ernst, Joseph R. Kiniry, Gary T. Leavens, K. Rustan M. Leino, and Erik Poll. An overview of JML tools and applications. International Journal on Software Tools for Technology Transfer (STTT), 7(3):212–232, June 2005. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=7&issue=3&spage=212.

Brucker:2005:VAA

[189] Achim D. Brucker and Burkhart Wolff. A verification approach to applied system security. International Journal on Software Tools for Technology Transfer (STTT), 7(3):233–247, June 2005. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=7&issue=3&spage=233.

Butterfield:2005:PHC

[190] Andrew Butterfield and Jim Woodcock. prialt in Handel-C: an operational semantics. International Journal on Software Tools for Technology Transfer (STTT), 7(3):248–267, June 2005. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=7&issue=3&spage=248.

Hammarberg:2005:FVF

[191] Jerker Hammarberg and Simin Nadjm-Tehrani. Formal verification of fault tolerance in safety-critical reconfigurable modules. International Journal on Software Tools for Technology Transfer (STTT), 7(3):268–279, June 2005. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=7&issue=3&spage=268.

Blom:2005:DSS

[192] Stefan Blom and Simona Orzan. Distributed state space minimization. International Journal on Software Tools for Technology Transfer (STTT), 7 (3):280–291, June 2005. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=7&issue=3&spage=280.

Margaria:2005:IP

[193] Tiziana Margaria. Introductory paper. International Journal on Software Tools for Technology Transfer (STTT), 7(4):293–296, August 2005. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=1433-2779&
                                        volume=7&issue=4&spage=293.

Jard:2005:TTP

[194] Claude Jard and Thierry Jéron. TGV: theory, principles and algorithms. International Journal on Software Tools for Technology Transfer (STTT), 7 (4):297–315, August 2005. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=7&issue=4&spage=297.

Viho:2005:TDS

[195] César Viho. Test distribution: a solution for complex network system testing. International Journal on Software Tools for Technology Transfer (STTT), 7(4):316–325, August 2005. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=7&issue=4&spage=316.

Baldini:2005:SLF

[196] Andrea Baldini, Alfredo Benso, and Paolo Prinetto. System-level functional testing from UML specifications in end-of-production industrial environments. International Journal on Software Tools for Technology Transfer (STTT), 7(4):326–340, August 2005. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=7&issue=4&spage=326.

Bunker:2005:LSC

[197] Annette Bunker, Ganesh Gopalakrishnan, and Konrad Slind. Live sequence charts applied to hardware requirements specification and verification. International Journal on Software Tools for Technology Transfer (STTT), 7(4):341–350, August 2005. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=7&issue=4&spage=341.

Schieferdecker:2005:DFL

[198] Ina Schieferdecker, George Din, and Dimitrios Apostolidis. Distributed functional and load tests for Web services. International Journal on Software Tools for Technology Transfer (STTT), 7(4):351–360, August 2005. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=7&issue=4&spage=351.

Turner:2005:TGR

[199] Kenneth J. Turner. Test generation for radiotherapy accelerators. International Journal on Software Tools for Technology Transfer (STTT), 7(4):361–375, August 2005. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1433-2779&volume=7&issue=4&spage=
361.

Kellerer:2005:PQA

[200] Bartholomäus Kellerer and Manfred Reitenspiess. Practical quality assurance for standards-based, high-availability middleware. International Journal on Software Tools for Technology Transfer (STTT), 7(4):376–387, August 2005. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=7&issue=4&spage=376.

Garavel:2006:WYS

[201] Hubert Garavel and John Hatcliff. Why you should definitely read this special section. International Journal on Software Tools for Technology Transfer (STTT), 8(1):1–3, February 2006. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=8&issue=1&spage=1.

Ciardo:2006:SAS

[202] Gianfranco Ciardo, Robert Marmorstein, and Radu Siminiceanu. The saturation algorithm for symbolic statespace exploration. International Journal on Software Tools for Technology Transfer (STTT), 8(1):4–25, February 2006. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=8&issue=1&spage=4.

Bartzis:2006:EBB

[203] Constantinos Bartzis and Tevfik Bultan. Efficient BDDs for bounded arithmetic constraints. International Journal on Software Tools for Technology Transfer (STTT), 8(1):26–36, February 2006. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=1433-2779&
volume=8&issue=1&spage=26.

Mateescu:2006:CGL

[204] Radu Mateescu. CAESAR SOLVE: a generic library for on-the-fly resolution of alternation-free Boolean equation systems. International Journal on Software Tools for Technology Transfer (STTT), 8(1):37–56, February 2006. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=8&issue=1&spage=37.

Bozga:2006:PBA

[205] L. Bozga, Y. Lakhnech, and M. Périn.

Pattern-based abstraction for verifying secrecy in protocols. International Journal on Software Tools for Technology Transfer (STTT), 8(1):57–76, February 2006. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1433-2779&volume=8&issue=1&spage=
57.

Groote:2006:IVL

[206] Jan Friso Groote and Frank van Ham.

Interactive visualization of large state spaces. International Journal on Software Tools for Technology Transfer (STTT), 8(1):77–91, February 2006. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=8&issue=1&spage=77.

Graf:2006:PSV

[207] Susanne Graf, Øystein Haugen, Ileana Ober, and Bran Selic. Preface of “Specification and Validation of Real Time and Embedded systems in UML”. International Journal on Software Tools for Technology Transfer (STTT), 8(2):93–96, April 2006. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=8&issue=2&spage=93.

Hooman:2006:SCR

[208] Jozef Hooman and Mark B. van der Zwaag. A semantics of communicating reactive objects with timing. International Journal on Software Tools for Technology Transfer (STTT), 8(2):97–112, April 2006. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=8&issue=2&spage=97.

Graf:2006:RTP

[209] Susanne Graf, Ileana Ober, and Iulian Ober. A real-time profile for UML. International Journal on Software Tools for Technology Transfer (STTT), 8(2):113–127, April 2006. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=8&issue=2&spage=113.

Ober:2006:VTU

[210] Iulian Ober, Susanne Graf, and Ileana Ober. Validating timed UML models by simulation and verification. International Journal on Software Tools for Technology Transfer (STTT), 8 (2):128–145, April 2006. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=8&issue=2&spage=128.

deSimone:2006:TSR

[211] Robert de Simone and Charles André.

Towards a “synchronous reactive” UML profile? International Journal on Software Tools for Technology Transfer (STTT), 8(2):146–155, April 2006. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=8&issue=2&spage=146.

Kuster-Filipe:2006:TEO

[212] Juliana Küster-Filipe and Stuart Anderson. On a time enriched OCL liveness template. International Journal on Software Tools for Technology Transfer (STTT), 8(2):156–166, April 2006. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=8&issue=2&spage=156.

Berkenkotter:2006:HPU

[213] Kirsten Berkenkötter, Stefan Bisanz, Ulrich Hannemann, and Jan Peleska. The HybridUML profile for UML 2.0. International Journal on Software Tools for Technology Transfer (STTT), 8(2):167–176, April 2006. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=8&issue=2&spage=167.

Jensen:2006:TAC

[214] Kurt Jensen and Andreas Podelski. Tools and algorithms for the construction and analysis of systems. International Journal on Software Tools for Technology Transfer (STTT), 8 (3):177–179, June 2006. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=8&issue=3&spage=177.

Valmari:2006:WSR

[215] Antti Valmari. What the small Rubik’s cube taught me about data structures, information theory, and randomisation. International Journal on Software Tools for Technology Transfer (STTT), 8(3):180–194, June 2006. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=8&issue=3&spage=180.

Schmidt:2006:AGP

[216] Karsten Schmidt. Automated generation of a progress measure for the sweepline method. International Journal on Software Tools for Technology Transfer (STTT), 8(3):195–203, June 2006. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=8&issue=3&spage=195.

Behrmann:2006:LUB

[217] Gerd Behrmann, Patricia Bouyer, Kim G. Larsen, and Radek Pelánek. Lower and upper bounds in zone-based abstractions of timed automata. International Journal on Software Tools for Technology Transfer (STTT), 8 (3):204–215, June 2006. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=8&issue=3&spage=204.

Younes:2006:NVS

[218] Håkan L. S. Younes, Marta Kwiatkowska, Gethin Norman, and David Parker. Numerical vs. statistical probabilistic model checking. International Journal on Software Tools for Technology Transfer (STTT), 8(3):216–228, June 2006. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=8&issue=3&spage=216.

Groce:2006:EED

[219] Alex Groce, Sagar Chaki, Daniel Kroening, and Ofer Strichman. Error explanation with distance metrics. International Journal on Software Tools for Technology Transfer (STTT), 8(3):229–247, June 2006. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1433-2779&volume=8&issue=3&spage=
229.

Sen:2006:OEP

[220] Koushik Sen, Grigore Rosu, and Gul Agha. Online efficient predictive safety analysis of multithreaded programs. International Journal on Software Tools for Technology Transfer (STTT), 8(3):248–260, June 2006. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=8&issue=3&spage=248.

Fang:2006:LIR

[221] Yi Fang, Nir Piterman, Amir Pnueli, and Lenore Zuck. Liveness with invisible ranking. International Journal on Software Tools for Technology Transfer (STTT), 8(3):261–279, June 2006. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=8&issue=3&spage=261.

Robby:2006:CJS

[222] Robby, Edwin Rodríguez, Matthew B. Dwyer, and John Hatcliff. Checking JML specifications using an extensible software model checking framework. International Journal on Software Tools for Technology Transfer (STTT), 8(3):280–299, June 2006. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=8&issue=3&spage=280.

Fitzgerald:2006:IFM

[223] John S. Fitzgerald, Stefania Gnesi, and Dino Mandrioli. The industrialization of formal methods. International Journal on Software Tools for Technology Transfer (STTT), 8(4–5):301–302, August 2006. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1433-2779&volume=8&issue=4&spage=

Miller:2006:PS

[224] Steven P. Miller, Alan C. Tribble, Michael W. Whalen, and Mats P. E. Heimdahl. Proving the shalls. International Journal on Software Tools for Technology Transfer (STTT), 8 (4–5):303–319, August 2006. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=8&issue=4&spage=303.

Fidge:2006:FCI

[225] C. J. Fidge. Formal change impact analyses for emulated control software. International Journal on Software Tools for Technology Transfer (STTT), 8 (4–5):321–335, August 2006. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=8&issue=4&spage=321.

Wassyng:2006:STS

[226] Alan Wassyng and Mark Lawford. Software tools for safety-critical software development. International Journal on Software Tools for Technology Transfer (STTT), 8(4–5):337–354, August 2006. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=8&issue=4&spage=337.

Tronci:2006:IP

[227] Enrico Tronci. Introductory paper. International Journal on Software Tools for Technology Transfer (STTT), 8 (4–5):355–358, August 2006. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=8&issue=4&spage=355.

Moore:2006:IAO

[228] J. Strother Moore. Inductive assertions and operational semantics. International Journal on Software Tools for Technology Transfer (STTT), 8(4–5):359–371, August 2006. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1433-2779&volume=8&issue=4&spage=
359.

Chockler:2006:CMF

[229] Hana Chockler, Orna Kupferman, and Moshe Vardi. Coverage metrics for formal verification. International Journal on Software Tools for Technology Transfer (STTT), 8(4–5):373–386, August 2006. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1433-2779&volume=8&issue=4&spage=
373.

Ganai:2006:EDS

[230] Malay K. Ganai, Aarti Gupta, Zijiang Yang, and Pranav Ashar. Efficient distributed SAT and SAT-based distributed bounded model checking. International Journal on Software Tools for Technology Transfer (STTT), 8 (4–5):387–396, August 2006. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=8&issue=4&spage=387.

Penna:2006:FHA

[231] Giuseppe Della Penna, Benedetto Intrigila, Igor Melatti, Enrico Tronci, and Marisa Venturini Zilli. Finite horizon analysis of Markov chains with the Murϕ verifier. International Journal on Software Tools for Technology Transfer (STTT), 8(4–5):397–409, August 2006. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=8&issue=4&spage=397.

Beyer:2006:PIA

[232] Sven Beyer, Christian Jacobi, Daniel Kröning, Dirk Leinenbach, and Wolfgang J. Paul. Putting it all together — formal verification of the VAMP. International Journal on Software Tools for Technology Transfer (STTT), 8 (4–5):411–430, August 2006. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=8&issue=4&spage=411.

Margaria:2006:SSL

[233] Tiziana Margaria and Bernhard Steffen. Special section on “leveraging formal methods”. International Journal on Software Tools for Technology Transfer (STTT), 8(6):467–469, November 2006. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=8&issue=6&spage=467.

Carcenac:2006:FFV

[234] François Carcenac and Frederic Boniol.

A formal framework for verifying distributed embedded systems based on abstraction methods. International Journal on Software Tools for Technology Transfer (STTT), 8(6):471–484, November 2006. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1433-2779&volume=8&issue=6&spage=
471.

Andersen:2006:CSC

[235] Jesper Andersen, Ebbe Elsborg, Fritz Henglein, Jakob Grue Simonsen, and Christian Stefansen. Compositional specification of commercial contracts. International Journal on Software Tools for Technology Transfer (STTT), 8 (6):485–516, November 2006. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=8&issue=6&spage=485.

Jahier:2006:CSL

[236] Erwan Jahier, Pascal Raymond, and Philippe Baufreton. Case studies with Lurette V2. International Journal on Software Tools for Technology Transfer (STTT), 8(6):517–530, November 2006. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=8&issue=6&spage=517.

Jorgensen:2006:TSE

[237] Jens Bæk Jørgensen, Søren Christensen, Antti-Pekka Tuovinen, and Jianli Xu. Tool support for estimating the memory usage of mobile phone software. International Journal on Software Tools for Technology Transfer (STTT), 8 (6):531–545, November 2006. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=8&issue=6&spage=531.

Ait-Ameur:2006:FEV

[238] Yamine Ait-Ameur and Mickael Baron.

Formal and experimental validation approaches in HCI systems design based on a shared event B model. International Journal on Software Tools for Technology Transfer (STTT), 8 (6):547–563, November 2006. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=8&issue=6&spage=547.

Alpuente:2006:RBV

[239] M. Alpuente, D. Ballis, and M. Falaschi.

Rule-based verification of Web sites. International Journal on Software Tools for Technology Transfer (STTT), 8 (6):565–585, November 2006. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=8&issue=6&spage=565.

Rouff:2006:EAF

[240] Christopher A. Rouff, Michael G.

Hinchey, Walter F. Truszkowski, and James L. Rash. Experiences applying formal approaches in the development of swarm-based space exploration systems. International Journal on Software Tools for Technology Transfer (STTT), 8(6):587–603, November 2006. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=8&issue=6&spage=587.

Buchholz:2006:GEI

[241] Peter Buchholz, Joost-Pieter Katoen, and Marcel Verhoef. Guest Editors’ introduction: quantitative analysis of real-time embedded systems. International Journal on Software Tools for Technology Transfer (STTT), 8 (6):605–606, November 2006. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=8&issue=6&spage=605.

Ploennigs:2006:AMG

[242] Joern Ploennigs, Mario Neugebauer, and Klaus Kabitzsch. Automated model generation for performance engineering of building automation networks. International Journal on Software Tools for Technology Transfer (STTT), 8 (6):607–620, November 2006. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=8&issue=6&spage=607.

DSouza:2007:EMP

[243] Deepak D’Souza and Pavithra Prabhakar. On the expressiveness of MTL in the pointwise and continuous semantics. International Journal on Software Tools for Technology Transfer (STTT), 9(1):1–4, February 2007. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=9&issue=1&spage=1.

Bozzano:2007:FNS

[244] Marco Bozzano and Adolfo Villafiorita.

The FSAP/NuSMV-SA safety analysis platform. International Journal on Software Tools for Technology Transfer (STTT), 9(1):5–24, February 2007. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=9&issue=1&spage=5.

Huth:2007:SCT

[245] Michael Huth. Some current topics in model checking. International Journal on Software Tools for Technology Transfer (STTT), 9(1):25–36, February 2007. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=9&issue=1&spage=25.

Dimovski:2007:CSV

[246] Aleksandar Dimovski and Ranko Lazić.

Compositional software verification based on game semantics and process algebra. International Journal on Software Tools for Technology Transfer (STTT), 9(1):37–51, February 2007. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=9&issue=1&spage=37.

Guelev:2007:MCP

[247] Dimitar P. Guelev, Mark D. Ryan, and Pierre Yves Schobbens. Modelchecking the preservation of temporal properties upon feature integration. International Journal on Software Tools for Technology Transfer (STTT), 9(1):53–62, February 2007. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=9&issue=1&spage=53.

Siminiceanu:2007:FVN

[248] Radu I. Siminiceanu and Gianfranco Ciardo. Formal verification of the NASA runway safety monitor. International Journal on Software Tools for Technology Transfer (STTT), 9 (1):63–76, February 2007. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=9&issue=1&spage=63.

Braunstein:2007:CPT

[249] Cécile Braunstein and Emmanuelle Encrenaz. CTL-property transformations along an incremental design process. International Journal on Software Tools for Technology Transfer (STTT), 9(1):77–88, February 2007. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=9&issue=1&spage=77.

Vasudevan:2007:IVH

[250] Shobha Vasudevan, E. Allen Emerson, and Jacob A. Abraham. Improved verification of hardware designs through antecedent conditioned slicing. International Journal on Software Tools for Technology Transfer (STTT), 9(1):89–101, February 2007. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=9&issue=1&spage=89.

Gasevic:2007:MBA

[251] Dragan Gašević, Dragan Djurić, and Vladan Devedžić. MDA-based automatic OWL ontology development. International Journal on Software Tools for Technology Transfer (STTT), 9 (2):103–117, March 2007. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=9&issue=2&spage=103.

Grumberg:2007:VFT

[252] Orna Grumberg and Shmuel Katz.

VeriTech: a framework for translating among model description notations. International Journal on Software Tools for Technology Transfer (STTT), 9(2):119–132, March 2007. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=9&issue=2&spage=119.

Pezze:2007:ISS

[253] Mauro Pezzè. Introduction to the special section on FASE 2003. International Journal on Software Tools for Technology Transfer (STTT), 9(2):133–134, March 2007. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=

                                        1433-2779&volume=9&issue=2&spage=
                                        133.

Vanderperren:2007:AOC

[254] Wim Vanderperren, Davy Suvée, Bruno De Fraine, and Viviane Jonckers. Aspect-oriented component composition in PacoSuite through invasive composition adapters. International Journal on Software Tools for Technology Transfer (STTT), 9(2):135–154, March 2007. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=9&issue=2&spage=135.

Sora:2007:CCC

[255] Ioana Sora, Pierre Verbaeten, and Yolande Berbers. CCDL: the Composable Components Description Language. International Journal on Software Tools for Technology Transfer (STTT), 9(2):155–168, March 2007. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=9&issue=2&spage=155.

Pahl:2007:OSC

[256] Claus Pahl. An ontology for software component matching. International Journal on Software Tools for Technology Transfer (STTT), 9(2):169–178, March 2007. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1433-2779&volume=9&issue=2&spage=
                                        169.

delBianco:2007:TUB

[257] Vieri del Bianco, Luigi Lavazza, Marco Mauri, and Giuseppe Occorso. Towards UML-based formal specifications of component-based real-time software. International Journal on Software Tools for Technology Transfer (STTT), 9(2):179–192, March 2007. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=9&issue=2&spage=179.

Heckel:2007:MDD

[258] Reiko Heckel and Marc Lohmann.

Model-driven development of reactive information systems: from graph transformation rules to JML contracts. International Journal on Software Tools for Technology Transfer (STTT), 9 (2):193–207, March 2007. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=9&issue=2&spage=193.

Jensen:2007:SSC

[259] Kurt Jensen. Special section on coloured Petri nets. International Journal on Software Tools for Technology Transfer (STTT), 9(3–4):209–212, June 2007. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=9&issue=3&spage=209.

Jensen:2007:CPN

[260] Kurt Jensen, Lars Michael Kristensen, and Lisa Wells. Coloured Petri nets and CPN tools for modelling and validation of concurrent systems. International Journal on Software Tools for Technology Transfer (STTT), 9(3–4):213–254, June 2007. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1433-2779&volume=9&issue=3&spage=
213.

Mitchell:2007:FSS

[261] Brice Mitchell, Lars Michael Kristensen, and Lin Zhang. Formal specification and state space analysis of an operational planning process. International Journal on Software Tools for Technology Transfer (STTT), 9(3–4):255–267, June 2007. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=9&issue=3&spage=255.

Billington:2007:MAF

[262] Jonathan Billington and Bing Han. Modelling and analysing the functional behaviour of TCP’s connection management procedures. International Journal on Software Tools for Technology Transfer (STTT), 9(3–4):269–304, June 2007. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=9&issue=3&spage=269.

Liu:2007:VCE

[263] Lin Liu and Jonathan Billington. Verification of the Capability Exchange Signalling protocol. International Journal on Software Tools for Technology Transfer (STTT), 9(3–4):305–326, June 2007.

CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=9&issue=3&spage=305.

Pesic:2007:MWD

[264] Maja Pesic and Wil M. P. van der Aalst.

Modelling work distribution mechanisms using colored Petri nets. International Journal on Software Tools for Technology Transfer (STTT), 9(3–4):327–352, June 2007. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1433-2779&volume=9&issue=3&spage=
327.

Machado:2007:RVE

[265] Ricardo J. Machado, Kristian Bisgaard Lassen, Sérgio Oliveira, Marco Couto, and Patrícia Pinto. Requirements validation: Execution of UML models with CPN tools. International Journal on Software Tools for Technology Transfer (STTT), 9(3–4):353–369, June 2007. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=9&issue=3&spage=353.

Gallasch:2007:CSP

[266] Guy Edward Gallasch, Jonathan Billington, Somsak Vanit-Anunchai, and Lars Michael Kristensen. Checking safety properties on-the-fly with the sweep-line method. International Journal on Software Tools for Technology Transfer (STTT), 9(3–4):371–391, June 2007. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1433-2779&volume=9&issue=3&spage=
                                        371.

Lakos:2007:MSS

[267] C. Lakos and L. Petrucci. Modular state space exploration for timed Petri nets. International Journal on Software Tools for Technology Transfer (STTT), 9(3–4):393–411, June 2007. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=9&issue=3&spage=393.

Bagnara:2007:WOP

[268] Roberto Bagnara, Patricia M. Hill, and Enea Zaffanella. Widening operators for powerset domains. International Journal on Software Tools for Technology Transfer (STTT), 9(3–4):413–414, June 2007. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1433-2779&volume=9&issue=3&spage=
                                        413.

Wermelinger:2007:ISS

[269] Michel Wermelinger, Tiziana Margaria, and Maura Cerioli. Introduction to the special section on fundamental approaches to software engineering. International Journal on Software Tools for Technology Transfer (STTT), 9 (5–6):415–416, October 2007. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=9&issue=5&spage=415.

Jung:2007:CFC

[270] Georg Jung and John Hatcliff. A correlation framework for the CORBA component model. International Journal on Software Tools for Technology Transfer (STTT), 9(5–6):417–427, October 2007. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=9&issue=5&spage=417.

Chechik:2007:FCG

[271] Marsha Chechik and Arie Gurfinkel.

A framework for counterexample generation and exploration. International Journal on Software Tools for Technology Transfer (STTT), 9(5–6):429–445, October 2007. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1433-2779&volume=9&issue=5&spage=
429.

Hamon:2007:OSS

[272] Grégoire Hamon and John Rushby.

An operational semantics for Stateflow. International Journal on Software Tools for Technology Transfer (STTT), 9(5–6):447–456, October 2007. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=9&issue=5&spage=447.

Breu:2007:MBD

[273] Ruth Breu, Gerhard Popp, and Muhammad Alam. Model based development of access policies. International Journal on Software Tools for Technology Transfer (STTT), 9(5–6):457–470, October 2007.

CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=9&issue=5&spage=457.

Larsen:2007:MSP

[274] Kim G. Larsen, Ulrik Nyman, and Andrzej Wasowski. Modeling software product lines using color-blind transition systems. International Journal on Software Tools for Technology Transfer (STTT), 9(5–6):471–487, October 2007. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=9&issue=5&spage=471.

Ranganath:2007:SCJ

[275] Venkatesh Prasad Ranganath and John Hatcliff. Slicing concurrent Java programs using Indus and Kaveri. International Journal on Software Tools for Technology Transfer (STTT), 9 (5–6):489–504, October 2007. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=9&issue=5&spage=489.

Beyer:2007:SMC

[276] Dirk Beyer, Thomas A. Henzinger, Ranjit Jhala, and Rupak Majumdar. The software model checker Blast. International Journal on Software Tools for Technology Transfer (STTT), 9 (5–6):505–525, October 2007. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=1433-2779&
volume=9&issue=5&spage=505.

Jurjens:2007:TSS

[277] Jan Jürjens and Pasha Shabalin. Tools for secure systems development with UML. International Journal on Software Tools for Technology Transfer (STTT), 9(5–6):527–544, October 2007. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=9&issue=5&spage=527.

Jensen:2008:SSC

[278] Kurt Jensen. Special section on coloured Petri nets. International Journal on Software Tools for Technology Transfer (STTT), 10(1):1–3, January 2008. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=10&issue=1&spage=1.

Kristensen:2008:MBD

[279] Lars M. Kristensen, Peter Mechlenborg, Lin Zhang, Brice Mitchell, and Guy E. Gallasch. Model-based development of a course of action scheduling tool. International Journal on Software Tools for Technology Transfer (STTT), 10(1):5–14, January 2008. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=10&issue=1&spage=5.

Jorgensen:2008:TDC

[280] Jens Bæk Jørgensen, Kristian Bisgaard Lassen, and Wil M. P. van der Aalst.

From task descriptions via colored Petri nets towards an implementation of a new electronic patient record workflow system. International Journal on Software Tools for Technology Transfer (STTT), 10(1):15–28, January 2008. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=10&issue=1&spage=15.

Vanit-Anunchai:2008:ADC

[281] Somsak Vanit-Anunchai, Jonathan Billington, and Guy Edward Gallasch. Analysis of the Datagram Congestion Control Protocol’s connection management procedures using the sweepline method. International Journal on Software Tools for Technology Transfer (STTT), 10(1):29–56, January 2008. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=10&issue=1&spage=29.

Rozinat:2008:DCP

[282] A. Rozinat, R. S. Mans, M. Song, and W. M. P. van der Aalst. Discovering colored Petri nets from event logs. International Journal on Software Tools for Technology Transfer (STTT), 10(1):57–74, January 2008. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=10&issue=1&spage=57.

Gallasch:2008:MDL

[283] Guy Edward Gallasch, Nimrod Lilith, Jonathan Billington, Lin Zhang, Axel Bender, and Benjamin Francis. Modelling defence logistics networks. International Journal on Software Tools for Technology Transfer (STTT), 10 (1):75–93, January 2008. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=10&issue=1&spage=75.

Gottschalk:2008:PUC

[284] F. Gottschalk, W. M. P. van der Aalst, M. H. Jansen-Vullers, and H. M. W. Verbeek. Protos2CPN: using colored Petri nets for configuring and testing business processes. International Journal on Software Tools for Technology Transfer (STTT), 10(1):95–110, January 2008. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=10&issue=1&spage=95.

Hermanns:2008:IES

[285] Holger Hermanns and Jens Palsberg.

Improving the effectiveness of system verification. International Journal on Software Tools for Technology Transfer (STTT), 10(2):111–112, March 2008. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=10&issue=2&spage=111.

Thomas:2008:EGS

[286] Dina Thomas, Supratik Chakraborty, and Paritosh Pandya. Efficient guided symbolic reachability using reachability expressions. International Journal on Software Tools for Technology Transfer (STTT), 10(2):113–129, March 2008. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=10&issue=2&spage=113.

Groce:2008:ETS

[287] Alex Groce and Rajeev Joshi. Exploiting traces in static program analysis: better model checking through printfs. International Journal on Software Tools for Technology Transfer (STTT), 10(2):131–144, March 2008. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=10&issue=2&spage=131.

Gupta:2008:AEE

[288] Atul Gupta and Pankaj Jalote. An approach for experimentally evaluating effectiveness and efficiency of coverage criteria for software testing. International Journal on Software Tools for Technology Transfer (STTT), 10 (2):145–160, March 2008. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=10&issue=2&spage=145.

Esparza:2008:NRD

[289] Javier Esparza, Pradeep Kanade, and Stefan Schwoon. A negative result on depth-first net unfoldings. International Journal on Software Tools for Technology Transfer (STTT), 10 (2):161–166, March 2008. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=10&issue=2&spage=161.

Hadjidj:2008:ISC

[290] Rachid Hadjidj and Hanifa Boucheneb. Improving state class constructions for CTL* model checking of time Petri nets. International Journal on Software Tools for Technology Transfer (STTT), 10(2):167–184, March 2008. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=10&issue=2&spage=167.

Leuschel:2008:PAA

[291] Michael Leuschel and Michael Butler.

ProB: an automated analysis toolset for the B method. International Journal on Software Tools for Technology Transfer (STTT), 10(2):185–203, March 2008. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=10&issue=2&spage=185.

Schafer:2008:ISS

[292] Wilhelm Schäfer and Matthias Tichy.

Introduction to the special section on self-optimizing mechatronic systems. International Journal on Software Tools for Technology Transfer (STTT), 10(3):205–206, June 2008. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=10&issue=3&spage=205.

Burmester:2008:TSD

[293] Sven Burmester, Holger Giese, Eckehard Münch, Oliver Oberschelp, Florian Klein, and Peter Scheideler. Tool support for the design of selfoptimizing mechatronic multi-agent systems. International Journal on Software Tools for Technology Transfer (STTT), 10(3):207–222, June 2008. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=10&issue=3&spage=207.

Witting:2008:NAO

[294] Katrin Witting, Bernd Schulz, Michael Dellnitz, Joachim Böcker, and Norbert Fröhleke. A new approach for online multiobjective optimization of mechatronic systems. International Journal on Software Tools for Technology Transfer (STTT), 10(3):223–231, June 2008. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=10&issue=3&spage=223.

Stein:2008:CLD

[295] Benno Stein. Coping with large design spaces: design problem solving in fluidic engineering. International Journal on Software Tools for Technology Transfer (STTT), 10(3):233–245, June 2008. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=10&issue=3&spage=233.

Trofin:2008:SVC

[296] Mircea Trofin and John Murphy. Static verification of component composition in contextual composition frameworks. International Journal on Software Tools for Technology Transfer (STTT), 10(3):247–261, June 2008. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=10&issue=3&spage=247.

Frehse:2008:PAV

[297] Goran Frehse. PHAVer: algorithmic verification of hybrid systems past HyTech. International Journal on Software Tools for Technology Transfer (STTT), 10(3):263–279, June 2008. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=10&issue=3&spage=263.

Schieferdecker:2008:ISS

[298] Ina Schieferdecker and Jens Grabowski.

Introduction to the special section on advances in test automation: the evolution of TTCN-3. International Journal on Software Tools for Technology Transfer (STTT), 10(4): 281–283, August 2008. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=10&issue=4&spage=281.

Botteck:2008:ITP

[299] Martin Botteck and Thomas Deiß. Introduction of TTCN-3 into the product development process: considerations from an electronic devices developer point of view. International Journal on Software Tools for Technology Transfer (STTT), 10(4):285–289, August 2008. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=10&issue=4&spage=285.

Pietschker:2008:ATA

[300] Andrej Pietschker. Automating test automation. International Journal on Software Tools for Technology Transfer (STTT), 10(4):291–295, August 2008. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=10&issue=4&spage=291.

Warken:2008:TAP

[301] Markus Warken. From testing to anti-product development. International Journal on Software Tools for Technology Transfer (STTT), 10(4): 297–307, August 2008. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=10&issue=4&spage=297.

Neukirchen:2008:AQE

[302] Helmut Neukirchen, Benjamin Zeiss, and Jens Grabowski. An approach to quality engineering of TTCN-3 test specifications. International Journal on Software Tools for Technology Transfer (STTT), 10(4):309–326, August 2008. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=1433-2779&
volume=10&issue=4&spage=309.

Schulz:2008:TSD

[303] Stephan Schulz. Test suite development with TTCN-3 libraries. International Journal on Software Tools for Technology Transfer (STTT), 10 (4):327–336, August 2008. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=10&issue=4&spage=327.

Sabiguero:2008:ACG

[304] Ariel Sabiguero, Anthony Baire, and César Viho. Automatic CoDec generation to reduce test engineering cost. International Journal on Software Tools for Technology Transfer (STTT), 10(4):337–346, August 2008. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=10&issue=4&spage=337.

Deiss:2008:RCT

[305] Thomas Deiß. Refactoring and converting a TTCN-2 test suite. International Journal on Software Tools for Technology Transfer (STTT), 10 (4):347–352, August 2008. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=10&issue=4&spage=347.

Glaser:2008:STS

[306] Michael Gläser, Sebastian Müller, Axel Rennoch, and Peter Schmitting. Standardized TTCN-3 specifications for SIP-ISUP/ISDN interworking testing. International Journal on Software Tools for Technology Transfer (STTT), 10 (4):353–358, August 2008. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=10&issue=4&spage=353.

Din:2008:IPB

[307] George Din. An IMS performance benchmark implementation based on the TTCN-3 language. International Journal on Software Tools for Technology Transfer (STTT), 10(4): 359–370, August 2008. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=10&issue=4&spage=359.

Stepien:2008:FTW

[308] Bernard Stepien, Liam Peyton, and Pulei Xiong. Framework testing of Web applications using TTCN-3. International Journal on Software Tools for Technology Transfer (STTT), 10 (4):371–381, August 2008. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=10&issue=4&spage=371.

Schieferdecker:2008:THC

[309] Ina Schieferdecker and Juergen Grossmann. Testing hybrid control systems with TTCN-3: an overview on continuous TTCN-3. International Journal on Software Tools for Technology Transfer (STTT), 10(4):383–400, August 2008. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=10&issue=4&spage=383.

Bardin:2008:FAT

[310] Sébastien Bardin, Alain Finkel, Jérôme Leroux, and Laure Petrucci. FAST: acceleration from theory to practice. International Journal on Software Tools for Technology Transfer (STTT), 10 (5):401–424, October 2008. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=10&issue=5&spage=401.

Ossowski:2008:UFW

[311] Jörn Ossowski and Christel Baier. A uniform framework for weighted decision diagrams and its implementation. International Journal on Software Tools for Technology Transfer (STTT), 10(5):425–441, October 2008. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=10&issue=5&spage=425.

Pelanek:2008:PSS

[312] Radek Pelánek. Properties of state spaces and their applications. International Journal on Software Tools for Technology Transfer (STTT), 10 (5):443–454, October 2008. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=1433-2779&
                                        volume=10&issue=5&spage=443.

Ebnenasir:2008:FFA

[313] Ali Ebnenasir, Sandeep S. Kulkarni, and Anish Arora. FTSyn: a framework for automatic synthesis of faulttolerance. International Journal on Software Tools for Technology Transfer (STTT), 10(5):455–471, October 2008. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=10&issue=5&spage=455.

Marchetto:2008:SST

[314] Alessandro Marchetto. Special section on testing and security of Web systems. International Journal on Software Tools for Technology Transfer (STTT), 10(6):473–476, December 2008. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=10&issue=6&spage=473.

Marchetto:2008:CSB

[315] Alessandro Marchetto, Filippo Ricca, and Paolo Tonella. A case studybased comparison of Web testing techniques applied to AJAX Web applications. International Journal on Software Tools for Technology Transfer (STTT), 10(6):477–492, December 2008. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=10&issue=6&spage=477.

Carzaniga:2008:HWA

[316] Antonio Carzaniga, Alessandra Gorla, and Mauro Pezzè. Healing Web applications through automatic workarounds. International Journal on Software Tools for Technology Transfer (STTT), 10 (6):493–502, December 2008. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=10&issue=6&spage=493.

Hughes:2008:AVA

[317] Graham Hughes and Tevfik Bultan.

Automated verification of access control policies using a SAT solver. International Journal on Software Tools for Technology Transfer (STTT), 10 (6):503–520, December 2008. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=10&issue=6&spage=503.

Kinder:2008:MPF

[318] Sebastian Kinder and Rolf Drechsler.

Modeling and proving functional completeness in formal verification of counting heads. International Journal on Software Tools for Technology Transfer (STTT), 10(6):521–534, December 2008. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=10&issue=6&spage=521.

Valmari:2009:SMC

[319] Antti Valmari. Software model checking is a rich research field. International Journal on Software Tools for Technology Transfer (STTT), 11 (1):1–11, February 2009. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=11&issue=1&spage=1.

Melatti:2009:PDM

[320] I. Melatti, R. Palmer, G. Sawaya, Y. Yang, R. M. Kirby, et al. Parallel and distributed model checking in Eddy. International Journal on Software Tools for Technology Transfer (STTT), 11(1):13–25, February 2009. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=11&issue=1&spage=13.

Drager:2009:DMC

[321] Klaus Dräger, Bernd Finkbeiner, and Andreas Podelski. Directed model checking with distance-preserving abstractions. International Journal on Software Tools for Technology Transfer (STTT), 11(1):27–37, February 2009. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=11&issue=1&spage=27.

Bosnacki:2009:POR

[322] Dragan Bosnacki, Stefan Leue, and Alberto Lluch Lafuente. Partial-order reduction for general state exploring algorithms. International Journal on Software Tools for Technology Transfer (STTT), 11(1):39–51, February 2009. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=11&issue=1&spage=39.

Anand:2009:SEA

[323] Saswat Anand, Corina S. Pasareanu, and Willem Visser. Symbolic execution with abstraction. International Journal on Software Tools for Technology Transfer (STTT), 11 (1):53–67, February 2009. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=11&issue=1&spage=53.

Armando:2009:BMC

[324] Alessandro Armando, Jacopo Mantovani, and Lorenzo Platania. Bounded model checking of software using SMT solvers instead of SAT solvers. International Journal on Software Tools for Technology Transfer (STTT), 11 (1):69–83, February 2009. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=11&issue=1&spage=69.

Huth:2009:SSA

[325] Michael Huth and Orna Grumberg. Special section on advances in reachability analysis and decision procedures: contributions to abstraction-based system verification. International Journal on Software Tools for Technology Transfer (STTT), 11(2):85–94, April 2009. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=1433-2779&
                                        volume=11&issue=2&spage=85.

Bryant:2009:ABD

[326] Randal E. Bryant, Daniel Kroening, Joël Ouaknine, Sanjit A. Seshia, Ofer Strichman, and Bryan Brady. An abstractionbased decision procedure for bit-vector arithmetic. International Journal on Software Tools for Technology Transfer (STTT), 11(2):95–104, April 2009. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=11&issue=2&spage=95.

Chatterjee:2009:LLM

[327] Shaunak Chatterjee, Shuvendu K. Lahiri, Shaz Qadeer, and Zvonimir Rakamarić. A low-level memory model and an accompanying reachability predicate. International Journal on Software Tools for Technology Transfer (STTT), 11(2):105–116, April 2009. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=11&issue=2&spage=105.

Yu:2009:DDB

[328] Andy Jinqing Yu, Gianfranco Ciardo, and Gerald Lüttgen. Decision-diagram-based techniques for bounded reachability checking of asynchronous systems. International Journal on Software Tools for Technology Transfer (STTT), 11(2):117–131, April 2009. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.

asp?genre=article&issn=1433-2779&
volume=11&issue=2&spage=117.

Tan:2009:WCE

[329] Lili Tan. The worst-case execution time tool challenge 2006. International Journal on Software Tools for Technology Transfer (STTT), 11 (2):133–152, April 2009. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=11&issue=2&spage=133.

Georgiou:2009:AIC

[330] Chryssis Georgiou, Nancy Lynch, Panayiotis Mavrommatis, and Joshua A. Tauber. Automated implementation of complex distributed algorithms specified in the IOA language. International Journal on Software Tools for Technology Transfer (STTT), 11 (2):153–171, April 2009. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=11&issue=2&spage=153.

Hinchey:2009:GEI

[331] Mike Hinchey, Tiziana Margaria, and Bernhard Steffen. Guest Editor’s introduction. International Journal on Software Tools for Technology Transfer (STTT), 11(3):173–174, July 2009. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=11&issue=3&spage=173.

Chetali:2009:ATE

[332] Boutheina Chetali and Quang-Huy Nguyen. An automated testing experiment for layered embedded C code. International Journal on Software Tools for Technology Transfer (STTT), 11(3):175–185, July 2009. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=11&issue=3&spage=175.

Schlich:2009:MCC

[333] Bastian Schlich and Stefan Kowalewski. Model checking C source code for embedded systems. International Journal on Software Tools for Technology Transfer (STTT), 11(3):187–202, July 2009. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=11&issue=3&spage=187.

VanWyk:2009:FML

[334] Eric Van Wyk and Mats Per Erik Heimdahl. Flexibility in modeling languages and tools: a call to arms. International Journal on Software Tools for Technology Transfer (STTT), 11(3):203–215, July 2009. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=11&issue=3&spage=203.

Cansell:2009:SCD

[335] Dominique Cansell, Dominique Méry, and Cyril Proch. System-on-chip design by proof-based refinement. International Journal on Software Tools for Technology Transfer (STTT), 11 (3):217–238, July 2009. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=11&issue=3&spage=217.

Ait-Ameur:2009:EPA

[336] Yamine Ait-Ameur, Mickael Baron, Nadjet Kamel, and Jean-Marc Mota. Encoding a process algebra using the Event B method application to the validation of human-computer interactions. International Journal on Software Tools for Technology Transfer (STTT), 11(3):239–253, July 2009. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=11&issue=3&spage=239.

Deharbe:2009:SSS

[337] David Déharbe and Silvio Ranise. Satisfiability solving for software verification. International Journal on Software Tools for Technology Transfer (STTT), 11(3):255–260, July 2009. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=11&issue=3&spage=255.

Desmoulin:2009:FIT

[338] Alexandra Desmoulin and César Viho.

Formalizing interoperability for test case generation purpose. International Journal on Software Tools for Technology Transfer (STTT), 11 (3):261–267, July 2009. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=11&issue=3&spage=261.

Yorav:2009:HVC

[339] Karen Yorav. Haifa verification conference 2007. International Journal on Software Tools for Technology Transfer (STTT), 11(4):269–272, October 2009. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=11&issue=4&spage=269.

Pasternak:2009:GUT

[340] Benny Pasternak, Shmuel Tyszberowicz, and Amiram Yehudai. GenUTest: a unit test and mock aspect generation tool. International Journal on Software Tools for Technology Transfer (STTT), 11(4):273–290, October 2009. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=11&issue=4&spage=273.

Fine:2009:UBN

[341] Shai Fine, Laurent Fournier, and Avi Ziv. Using Bayesian networks and virtual coverage to hit hard-to-reach events. International Journal on Software Tools for Technology Transfer (STTT), 11(4):291–305, October 2009. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=11&issue=4&spage=291.

Raffelt:2009:DTA

[342] Harald Raffelt, Maik Merten, Bernhard Steffen, and Tiziana Margaria. Dynamic testing via automata learning. International Journal on Software Tools for Technology Transfer (STTT), 11(4):307–324, October 2009. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=11&issue=4&spage=307.

Babic:2009:ASR

[343] Domagoj Babić and Alan J. Hu. Approximating the safely reusable set of learned facts. International Journal on Software Tools for Technology Transfer (STTT), 11(4):325–338, October 2009. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=11&issue=4&spage=325.

Pasareanu:2009:SNT

[344] Corina S. Pasareanu and Willem Visser.

A survey of new trends in symbolic execution for software testing and analysis. International Journal on Software Tools for Technology Transfer (STTT), 11(4):339–353, October 2009. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=11&issue=4&spage=339.

Margaria:2009:P

[345] Tiziana Margaria and Mieke Massink.

Preface. International Journal on Software Tools for Technology Transfer (STTT), 11(5):355–357, November 2009. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1433-2779&volume=11&issue=5&spage=
355.

delaCamara:2009:CRS

[346] Pedro de la Cámara, María del Mar Gallardo, Pedro Merino, and David Sanán. Checking the reliability of socket based communication software. International Journal on Software Tools for Technology Transfer (STTT), 11 (5):359–374, November 2009. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=11&issue=5&spage=359.

Wijs:2009:SSP

[347] Anton J. Wijs, Jaco C. van de Pol, and Elena M. Bortnik. Solving scheduling problems by untimed model checking: The clinical chemical analyser case study. International Journal on Software Tools for Technology Transfer (STTT), 11 (5):375–392, November 2009. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=11&issue=5&spage=375.

Raffelt:2009:LFE

[348] Harald Raffelt, Bernhard Steffen, Therese Berg, and Tiziana Margaria. LearnLib: a framework for extrapolating behavioral models. International Journal on Software Tools for Technology Transfer (STTT), 11 (5):393–407, November 2009. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=11&issue=5&spage=393.

Mikac:2009:FED

[349] Jan Mikác and Paul Caspi. Flush:

an example of development by refinements in SCADE/Lustre. International Journal on Software Tools for Technology Transfer (STTT), 11 (5):409–418, November 2009. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=11&issue=5&spage=409.

Ricca:2009:SSW

[350] Filippo Ricca and Liu Chao. Special section on Web Systems Evolution. International Journal on Software Tools for Technology Transfer (STTT), 11(6):419–425, December 2009. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=11&issue=6&spage=419.

Marchetto:2009:OST

[351] Alessandro Marchetto and Filippo Ricca. From objects to services: toward a stepwise migration approach for Java applications. International Journal on Software Tools for Technology Transfer (STTT), 11(6):427–440, December 2009. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=1433-2779&
                                        volume=11&issue=6&spage=427.

Sneed:2009:PPM

[352] Harry M. Sneed. A pilot project for migrating COBOL code to Web services. International Journal on Software Tools for Technology Transfer (STTT), 11(6):441–451, December 2009. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=11&issue=6&spage=441.

Nakano:2009:CWS

[353] Keisuke Nakano, Zhenjiang Hu, and Masato Takeichi. Consistent Web site updating based on bidirectional transformation. International Journal on Software Tools for Technology Transfer (STTT), 11(6):453–468, December 2009. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=11&issue=6&spage=453.

Scanniello:2009:AEB

[354] Giuseppe Scanniello, Damiano Distante, and Michele Risi. An approach and an Eclipse-based environment for enhancing the navigation structure of Web sites. International Journal on Software Tools for Technology Transfer (STTT), 11(6):469–484, December 2009. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=11&issue=6&spage=469.

Bernardi:2009:RUA

[355] Mario Luca Bernardi, Giuseppe Antonio Di Lucca, and Damiano Distante. The RE-UWA approach to recover user centered conceptual models from Web applications. International Journal on Software Tools for Technology Transfer (STTT), 11(6):485–501, December 2009. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=11&issue=6&spage=485.

Chan:2009:AET

[356] Brian Chan, King Chun Foo, Lionel Marks, and Ying Zou. An approach for estimating the time needed to perform code changes in business applications. International Journal on Software Tools for Technology Transfer (STTT), 11(6):503–515, December 2009. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=11&issue=6&spage=503.

Ameur:2010:TWU

[357] Yamine Ait Ameur, Frédéric Boniol, and Virginie Wiels. Toward a wider use of formal methods for aerospace systems design and verification. International Journal on Software Tools for Technology Transfer (STTT), 12(1):1–7, February 2010. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=12&issue=1&spage=1.

Plagge:2010:SOS

[358] Daniel Plagge and Michael Leuschel. Seven at one stroke: LTL model checking for high-level specifications in B, Z, CSP, and more. International Journal on Software Tools for Technology Transfer (STTT), 12(1):9–21, February 2010. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=12&issue=1&spage=9.

Bauer:2010:DCS

[359] Andreas Bauer, Martin Leucker, Christian Schallhart, and Michael Tautschnig. Don’t care in SMT: building flexible yet efficient abstraction/refinement solvers. International Journal on Software Tools for Technology Transfer (STTT), 12(1):23–37, February 2010. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=12&issue=1&spage=23.

Rehm:2010:PDR

[360] Joris Rehm. Proved development of the real-time properties of the IEEE 1394 Root Contention Protocol with the event-B method. International Journal on Software Tools for Technology Transfer (STTT), 12 (1):39–51, February 2010. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=12&issue=1&spage=39.

Dolev:2010:FRA

[361] Shlomi Dolev and Ori Gersten. A framework for robust active super tier systems. International Journal on Software Tools for Technology Transfer (STTT), 12(1):53–67, February 2010. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=12&issue=1&spage=53.

Carver:2010:CLI

[362] Richard H. Carver and Yu Lei. A class library for implementing, testing, and debugging concurrent programs. International Journal on Software Tools for Technology Transfer (STTT), 12(1):69–88, February 2010. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=12&issue=1&spage=69.

Bosnacki:2010:MCS

[363] Dragan Bosnacki and Stefan Edelkamp.

Model checking software: on some new waves and some evergreens. International Journal on Software Tools for Technology Transfer (STTT), 12 (2):89–95, May 2010. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=12&issue=2&spage=89.

Weber:2010:EVM

[364] Michael Weber. An embeddable virtual machine for state space generation. International Journal on Software Tools for Technology Transfer (STTT), 12(2):97–111, May 2010. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=12&issue=2&spage=97.

Yang:2010:DDP

[365] Yu Yang, Xiaofang Chen, Ganesh Gopalakrishnan, and Robert M. Kirby. Distributed dynamic partial order reduction. International Journal on Software Tools for Technology Transfer (STTT), 12(2):113–122, May 2010. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=12&issue=2&spage=113.

Rozier:2010:LSC

[366] Kristin Y. Rozier and Moshe Y. Vardi. LTL satisfiability checking. International Journal on Software Tools for Technology Transfer (STTT), 12 (2):123–137, May 2010. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=12&issue=2&spage=123.

Barnat:2010:SSM

[367] J. Barnat, L. Brim, and P. Rockai. Scalable shared memory LTL model checking. International Journal on Software Tools for Technology Transfer (STTT), 12(2):139–153, May 2010. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=12&issue=2&spage=139.

Evangelista:2010:SIP

[368] Sami Evangelista and Christophe Pajault. Solving the ignoring problem for partial order reduction. International Journal on Software Tools for Technology Transfer (STTT), 12 (2):155–170, May 2010. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=12&issue=2&spage=155.

Rensink:2010:GTT

[369] Arend Rensink and Pieter Van Gorp.

Graph transformation tool contest 2008. International Journal on Software Tools for Technology Transfer (STTT), 12(3–4):171–181, July 2010. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=12&issue=3&spage=171.

Perez:2010:CSE

[370] Javier Pérez, Yania Crespo, Berthold Hoffmann, and Tom Mens. A case study to evaluate the suitability of graph transformation tools for program refactoring. International Journal on Software Tools for Technology Transfer (STTT), 12(3–4):183–199, July 2010. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=12&issue=3&spage=183.

Muliawan:2010:MRU

[371] Olaf Muliawan and Dirk Janssens.

Model refactoring using MoTMoT. International Journal on Software Tools for Technology Transfer (STTT), 12 (3–4):201–209, July 2010. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=12&issue=3&spage=201.

Horvath:2010:EAC

[372] Ákos Horváth, Gábor Bergmann, István Ráth, and Dániel Varró. Experimental assessment of combining pattern matching strategies with VIA-TRA2. International Journal on Software Tools for Technology Transfer (STTT), 12(3–4):211–230, July 2010. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=12&issue=3&spage=211.

Meszaros:2010:MAP

[373] Tamás Mészáros, Gergely Mezei, Tihamér Levendovszky, and Márk Asztalos. Manual and automated performance optimization of model transformation systems. International Journal on Software Tools for Technology Transfer (STTT), 12(3–4):231–243, July 2010. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=12&issue=3&spage=231.

Biermann:2010:IAE

[374] Enrico Biermann, Claudia Ermel, Leen Lambers, Ulrike Prange, Olga Runge, et al. Introduction to AGG and EMF Tiger by modeling a Conference Scheduling System. International Journal on Software Tools for Technology Transfer (STTT), 12(3– 4):245–261, July 2010. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=12&issue=3&spage=245.

Jakumeit:2010:GNE

[375] Edgar Jakumeit, Sebastian Buchwald, and Moritz Kroll. GrGen.NET: The expressive, convenient and fast graph rewrite system. International Journal on Software Tools for Technology Transfer (STTT), 12(3–4):263–271, July 2010. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=12&issue=3&spage=263.

Moha:2010:EKS

[376] Naouel Moha, Sagar Sen, Cyril Faucher, Olivier Barais, and Jean-Marc Jézéquel. Evaluation of kermeta for solving graphbased problems. International Journal on Software Tools for Technology Transfer (STTT), 12(3–4):273–285, July 2010. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=12&issue=3&spage=273.

Geiger:2010:FCS

[377] Leif Geiger and Albert Zündorf. Fujaba case studies for GraBaTs 2008: lessons learned. International Journal on Software Tools for Technology Transfer (STTT), 12(3–4):287–304, July 2010. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=1433-2779&
                                        volume=12&issue=3&spage=287.

Mader:2010:SSA

[378] Angelika Mader, Henrik Bohnenkamp, Yaroslav S. Usenko, David N. Jansen, Johann Hurink, and Holger Hermanns. Synthesis and stochastic assessment of cost-optimal schedules. International Journal on Software Tools for Technology Transfer (STTT), 12 (5):305–318, September 2010. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=12&issue=5&spage=305.

Simmonds:2010:ERP

[379] Jocelyn Simmonds, Jessica Davies, Arie Gurfinkel, and Marsha Chechik. Exploiting resolution proofs to speed up LTL vacuity detection for BMC. International Journal on Software Tools for Technology Transfer (STTT), 12 (5):319–335, September 2010. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=12&issue=5&spage=319.

Beaudenon:2010:DDD

[380] Vincent Beaudenon, Emmanuelle Encrenaz, and Sami Taktak. Data decision diagrams for Promela systems analysis. International Journal on Software Tools for Technology Transfer (STTT), 12(5):337–352, September 2010. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.

asp?genre=article&issn=1433-2779&
volume=12&issue=5&spage=337.

Krahn:2010:MFC

[381] Holger Krahn, Bernhard Rumpe, and Steven Völkel. MontiCore: a framework for compositional development of domain specific languages. International Journal on Software Tools for Technology Transfer (STTT), 12 (5):353–372, September 2010. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=12&issue=5&spage=353.

Bakewell:2010:DAR

[382] Adam Bakewell, Aleksandar Dimovski, Dan R. Ghica, and Ranko Lazić. Data-abstraction refinement: a game semantic approach. International Journal on Software Tools for Technology Transfer (STTT), 12(5):373– 389, September 2010. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=12&issue=5&spage=373.

Bucci:2010:OTM

[383] Giacomo Bucci, Laura Carnevali, Lorenzo Ridi, and Enrico Vicario. Oris: a tool for modeling, verification and evaluation of real-time systems. International Journal on Software Tools for Technology Transfer (STTT), 12 (5):391–403, September 2010. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=1433-2779&
volume=12&issue=5&spage=391.

Kroening:2010:VST

[384] Daniel Kroening and Tiziana Margaria. Verified software: theories, tools and experiments. International Journal on Software Tools for Technology Transfer (STTT), 12 (6):405–408, November 2010. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=12&issue=6&spage=405.

Gurfinkel:2010:CPN

[385] Arie Gurfinkel and Sagar Chaki. Combining predicate and numeric abstraction for software model checking. International Journal on Software Tools for Technology Transfer (STTT), 12 (6):409–427, November 2010. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=12&issue=6&spage=409.

Chalin:2010:TIG

[386] Patrice Chalin, Robby, Perry R. James, Jooyong Lee, and George Karabotsos. Towards an industrial grade IVE for Java and next generation research platform for JML. International Journal on Software Tools for Technology Transfer (STTT), 12 (6):429–446, November 2010. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=12&issue=6&spage=429.

Abrial:2010:ROT

[387] Jean-Raymond Abrial, Michael Butler, Stefan Hallerstede, Thai Son Hoang, Farhad Mehta, et al. Rodin: an open toolset for modelling and reasoning in Event-B. International Journal on Software Tools for Technology Transfer (STTT), 12(6):447– 466, November 2010. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=12&issue=6&spage=447.

Cok:2010:IUP

[388] David R. Cok. Improved usability and performance of SMT solvers for debugging specifications. International Journal on Software Tools for Technology Transfer (STTT), 12 (6):467–481, November 2010. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=12&issue=6&spage=467.

Pasareanu:2011:NRS

[389] Corina S. Pasareanu. New results in software model checking and analysis. International Journal on Software Tools for Technology Transfer (STTT), 13(1):1–2, January 2011. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=13&issue=1&spage=1.

Hahn:2011:PRP

[390] Ernst Moritz Hahn, Holger Hermanns, and Lijun Zhang. Probabilistic reachability for parametric Markov models. International Journal on Software Tools for Technology Transfer (STTT), 13(1):3–19, January 2011. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=13&issue=1&spage=3.

Bosnacki:2011:PPM

[391] Dragan Bosnacki, Stefan Edelkamp, Damian Sulewski, and Anton Wijs. Parallel probabilistic model checking on general purpose graphics processors. International Journal on Software Tools for Technology Transfer (STTT), 13(1):21–35, January 2011. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=13&issue=1&spage=21.

Kidd:2011:DPD

[392] Nicholas Kidd, Peter Lammich, Tayssir Touili, and Thomas Reps. A decision procedure for detecting atomicity violations for communicating processes with locks. International Journal on Software Tools for Technology Transfer (STTT), 13(1):37–60, January 2011. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=13&issue=1&spage=37.

Lim:2011:SAS

[393] Junghee Lim, Akash Lal, and Thomas Reps. Symbolic analysis via semantic reinterpretation. International Journal on Software Tools for Technology Transfer (STTT), 13(1):61–87, January 2011. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=13&issue=1&spage=61.

Schmerl:2011:UMC

[394] Sebastian Schmerl, Michael Vogel, and Hartmut König. Using model checking to identify errors in intrusion detection signatures. International Journal on Software Tools for Technology Transfer (STTT), 13(1):89–106, January 2011. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=13&issue=1&spage=89.

Zou:2011:GEI

[395] Ying Zou, Ji Wu, and Kenny Wong.

Guest editors’ introduction to the special section from the International Symposium on Web Systems Evolution. International Journal on Software Tools for Technology Transfer (STTT), 13(2):107–109, April 2011. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=13&issue=2&spage=107.

Dobolyi:2011:ART

[396] Kinga Dobolyi, Elizabeth Soechting, and Westley Weimer. Automating regression testing using Webbased application similarities. International Journal on Software Tools for Technology Transfer (STTT), 13 (2):111–129, April 2011. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=13&issue=2&spage=111.

Marchetto:2011:CMA

[397] Alessandro Marchetto, Roberto Tiella, Paolo Tonella, Nadia Alshahwan, and Mark Harman. Crawlability metrics for automated Web testing. International Journal on Software Tools for Technology Transfer (STTT), 13 (2):131–149, April 2011. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=13&issue=2&spage=131.

Torchiano:2011:WAM

[398] Marco Torchiano, Filippo Ricca, and Alessandro Marchetto. Are Web applications more defect-prone than desktop applications? International Journal on Software Tools for Technology Transfer (STTT), 13(2):151–166, April 2011. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=13&issue=2&spage=151.

Mosincat:2011:AMS

[399] Adina Mosincat and Walter Binder. Automated maintenance of service compositions with SLA violation detection and dynamic binding. International Journal on Software Tools for Technology Transfer (STTT), 13(2):167–179, April 2011. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.

asp?genre=article&issn=1433-2779&
volume=13&issue=2&spage=167.

DAmbros:2011:PSV

[400] Marco D’Ambros, Michele Lanza, Mircea Lungu, and Romain Robbes. On porting software visualization tools to the web. International Journal on Software Tools for Technology Transfer (STTT), 13(2):181–200, April 2011. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=13&issue=2&spage=181.

Chockler:2011:P

[401] Hana Chockler and Alan J. Hu. Preface. International Journal on Software Tools for Technology Transfer (STTT), 13(3):201–205, June 2011. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=13&issue=3&spage=201.

Adler:2011:EWU

[402] Yoram Adler, Dale Blue, Thomas Conti, Richard Prewitt, and Shmuel Ur. Evaluating workloads using comparative functional coverage. International Journal on Software Tools for Technology Transfer (STTT), 13(3):207–221, June 2011. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=13&issue=3&spage=207.

Artho:2011:IDD

[403] Cyrille Artho. Iterative delta debugging. International Journal on Software Tools for Technology Transfer (STTT), 13(3):223–246, June 2011. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=13&issue=3&spage=223.

Baras:2011:ABC

[404] Dorit Baras, Shai Fine, Laurent Fournier, Dan Geiger, and Avi Ziv. Automatic boosting of cross-product coverage using Bayesian networks. International Journal on Software Tools for Technology Transfer (STTT), 13 (3):247–261, June 2011. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=13&issue=3&spage=247.

Bar-Ilan:2011:RSR

[405] Omer Bar-Ilan, Oded Fuhrmann, Shlomo Hoory, Ohad Shacham, and Ofer Strichman. Reducing the size of resolution proofs in linear time. International Journal on Software Tools for Technology Transfer (STTT), 13 (3):263–272, June 2011. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=13&issue=3&spage=263.

Bauer:2011:UAT

[406] K. Bauer, R. Gentilini, and K. Schneider. A uniform approach to threevalued semantics for μ-calculus on abstractions of hybrid automata. International Journal on Software Tools for Technology Transfer (STTT), 13 (3):273–287, June 2011. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=13&issue=3&spage=273.

Fecher:2011:LAR

[407] Harald Fecher and Sharon Shoham. Local abstraction-refinement for the μcalculus. International Journal on Software Tools for Technology Transfer (STTT), 13(4):289–306, August 2011. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=13&issue=4&spage=289.

Bu:2011:POB

[408] Lei Bu and Xuandong Li. Pathoriented bounded reachability analysis of composed linear hybrid systems. International Journal on Software Tools for Technology Transfer (STTT), 13(4):307–317, August 2011. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=13&issue=4&spage=307.

Sebastiani:2011:SSE

[409] Roberto Sebastiani, Stefano Tonetta, and Moshe Y. Vardi. Symbolic systems, explicit properties: on hybrid approaches for LTL symbolic model checking. International Journal on Software Tools for Technology Transfer (STTT), 13(4):319–335, August 2011. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=1433-2779&
                                        volume=13&issue=4&spage=319.

Lisboa:2011:TTD

[410] Liana Barachisio Lisboa, Vinicius Cardoso Garcia, Eduardo Santana de Almeida, and Silvio Romero de Lemos Meira. ToolDAy: a tool for domain analysis. International Journal on Software Tools for Technology Transfer (STTT), 13(4):337–353, August 2011. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=13&issue=4&spage=337.

Guo:2011:FBT

[411] Qiang Guo and John Derrick. Formally based tool support for model checking Erlang applications. International Journal on Software Tools for Technology Transfer (STTT), 13 (4):355–376, August 2011. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=13&issue=4&spage=355.

Brukman:2011:ROP

[412] Olga Brukman and Shlomi Dolev. Recovery oriented programming: runtime monitoring of safety and liveness. International Journal on Software Tools for Technology Transfer (STTT), 13(4):377–395, August 2011. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=13&issue=4&spage=377.

Filliatre:2011:DSV

[413] Jean-Christophe Filliâtre. Deductive software verification. International Journal on Software Tools for Technology Transfer (STTT), 13(5): 397–403, October 2011. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=13&issue=5&spage=397.

Cuoq:2011:FDC

[414] Pascal Cuoq, Benjamin Monate, Anne Pacalet, and Virgile Prevosto. Functional dependencies of C functions via weakest pre-conditions. International Journal on Software Tools for Technology Transfer (STTT), 13 (5):405–417, October 2011. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=13&issue=5&spage=405.

Weber:2011:SSN

[415] Tjark Weber. SMT solvers: new oracles for the HOL theorem prover. International Journal on Software Tools for Technology Transfer (STTT), 13 (5):419–429, October 2011. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=13&issue=5&spage=419.

Castillos:2011:SBT

[416] Kalou Cabrera Castillos, Frédéric Dadeau, and Jacques Julliand. Scenariobased testing from UML/OCL behavioral models: Application to POSIX compliance. International Journal on Software Tools for Technology Transfer (STTT), 13(5):431–448, October 2011. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=13&issue=5&spage=431.

Ishii:2011:IBS

[417] Daisuke Ishii, Kazunori Ueda, and Hiroshi Hosobe. An interval-based SAT modulo ODE solver for model checking nonlinear hybrid systems. International Journal on Software Tools for Technology Transfer (STTT), 13 (5):449–461, October 2011. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=13&issue=5&spage=449.

Freitas:2011:FMS

[418] Leo Freitas and John McDermott. Formal methods for security in the Xenon hypervisor. International Journal on Software Tools for Technology Transfer (STTT), 13(5):463–489, October 2011. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=13&issue=5&spage=463.

Jones:2011:PSS

[419] Neil D. Jones and Markus Müller-Olm. Preface to a special section on verification, model checking, and abstract interpretation. International Journal on Software Tools for Technology Transfer (STTT), 13 (6):491–493, November 2011. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=13&issue=6&spage=491.

Kidd:2011:FCR

[420] Nicholas Kidd, Thomas Reps, Julian Dolby, and Mandana Vaziri. Finding concurrency-related bugs using random isolation. International Journal on Software Tools for Technology Transfer (STTT), 13(6):495– 518, November 2011. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=13&issue=6&spage=495.

Taly:2011:SSL

[421] Ankur Taly, Sumit Gulwani, and Ashish Tiwari. Synthesizing switching logic using constraint solving. International Journal on Software Tools for Technology Transfer (STTT), 13 (6):519–535, November 2011. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=13&issue=6&spage=519.

Etessami:2011:AAM

[422] Kousha Etessami and Patrice Godefroid. An abort-aware model of transactional programming. International Journal on Software Tools for Technology Transfer (STTT), 13 (6):537–551, November 2011. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=1433-2779&
                                        volume=13&issue=6&spage=537.

Rutkowski:2011:APP

[423] Michal Rutkowski, Ranko Lazić, and Marcin Jurdziński. Average-price-per-reward games on hybrid automata with strong resets. International Journal on Software Tools for Technology Transfer (STTT), 13 (6):553–569, November 2011. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=13&issue=6&spage=553.

Godefroid:2011:LGM

[424] Patrice Godefroid and Nir Piterman. LTL generalized model checking revisited. International Journal on Software Tools for Technology Transfer (STTT), 13(6):571–584, November 2011. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=13&issue=6&spage=571.

Laviron:2011:SFN

[425] Vincent Laviron and Francesco Logozzo. SubPolyhedra: a family of numerical abstract domains for the (more) scalable inference of linear inequalities. International Journal on Software Tools for Technology Transfer (STTT), 13 (6):585–601, November 2011. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=13&issue=6&spage=585.

Sharygina:2012:ARA

[426] Natasha Sharygina, Stefano Tonetta, and Aliaksei Tsitovich. An abstraction refinement approach combining precise and approximated techniques. International Journal on Software Tools for Technology Transfer (STTT), 14(1):1–14, February 2012. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=14&issue=1&spage=1.

Ghamarian:2012:MAU

[427] Amir Hossein Ghamarian, Maarten de Mol, Arend Rensink, Eduardo Zambon, and Maria Zimakova. Modelling and analysis using GROOVE. International Journal on Software Tools for Technology Transfer (STTT), 14 (1):15–40, February 2012. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=14&issue=1&spage=15.

Katz:2012:CAP

[428] Shmuel Katz and David Faitelson.

The Common Aspect Proof Environment. International Journal on Software Tools for Technology Transfer (STTT), 14(1):41–52, February 2012. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=14&issue=1&spage=41.

Basu:2012:SAM

[429] Ananda Basu, Saddek Bensalem, Marius Bozga, Benoît Delahaye, and Axel Legay. Statistical abstraction and model-checking of large heterogeneous systems. International Journal on Software Tools for Technology Transfer (STTT), 14(1):53–72, February 2012. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=14&issue=1&spage=53.

Denise:2012:CBR

[430] Alain Denise, Marie-Claude Gaudel, Sandrine-Dominique Gouraud, Richard Lassaigne, Johan Oudinet, et al. Coverage-biased random exploration of large models and application to testing. International Journal on Software Tools for Technology Transfer (STTT), 14(1):73–93, February 2012. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=14&issue=1&spage=73.

Safe:2012:SFV

[431] Georgia Penido Safe, Claudionor Coelho, Luiz Filipe M. Vieira, Celina Gomes Do Val, Jose Augusto Nacif, et al. Selection of formal verification heuristics for parallel execution. International Journal on Software Tools for Technology Transfer (STTT), 14(1):95–108, February 2012. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=14&issue=1&spage=95.

Abdulla:2012:RMCa

[432] Parosh Aziz Abdulla. Regular model checking. International Journal on Software Tools for Technology Transfer (STTT), 14(2):109–118, April 2012. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=14&issue=2&spage=109.

Legay:2012:EOR

[433] Axel Legay. Extrapolating (omega- )regular model checking. International Journal on Software Tools for Technology Transfer (STTT), 14 (2):119–143, April 2012. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=14&issue=2&spage=119.

Bouajjani:2012:WTR

[434] Ahmed Bouajjani and Tayssir Touili.

Widening techniques for regular tree model checking. International Journal on Software Tools for Technology Transfer (STTT), 14(2):145–165, April 2012. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=14&issue=2&spage=145.

Bouajjani:2012:ART

[435] Ahmed Bouajjani, Peter Habermehl, Adam Rogalewicz, and Tomás Vojnar. Abstract regular (tree) model checking. International Journal on Software Tools for Technology Transfer (STTT), 14(2):167–191, April 2012. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=1433-2779&
                                        volume=14&issue=2&spage=167.

Boigelot:2012:DSR

[436] Bernard Boigelot. Domain-specific regular acceleration. International Journal on Software Tools for Technology Transfer (STTT), 14(2):193–206, April 2012. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=14&issue=2&spage=193.

Delzanno:2012:LRM

[437] Giorgio Delzanno and Ahmed Rezine. A lightweight regular model checking approach for parameterized systems. International Journal on Software Tools for Technology Transfer (STTT), 14(2):207–222, April 2012. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=14&issue=2&spage=207.

Abdulla:2012:RMCb

[438] Parosh Aziz Abdulla, Bengt Jonsson, Marcus Nilsson, Julien d’Orso, and Mayank Saksena. Regular model checking for LTL(MSO). International Journal on Software Tools for Technology Transfer (STTT), 14 (2):223–241, April 2012. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=14&issue=2&spage=223.

Sokolsky:2012:ISS

[439] Oleg Sokolsky, Klaus Havelund, and Insup Lee. Introduction to the special section on runtime verification. International Journal on Software Tools for Technology Transfer (STTT), 14(3):243–247, June 2012. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=14&issue=3&spage=243.

Meredith:2012:OMR

[440] Patrick O’Neil Meredith, Dongyun Jin, Dennis Griffith, Feng Chen, and Grigore Rosu. An overview of the MOP runtime verification framework. International Journal on Software Tools for Technology Transfer (STTT), 14 (3):249–289, June 2012. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=14&issue=3&spage=249.

Qadeer:2012:RVC

[441] Shaz Qadeer and Serdar Tasiran.

Runtime verification of concurrencyspecific correctness criteria. International Journal on Software Tools for Technology Transfer (STTT), 14 (3):291–305, June 2012. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=14&issue=3&spage=291.

Bodden:2012:CFH

[442] Eric Bodden and Laurie Hendren. The Clara framework for hybrid typestate analysis. International Journal on Software Tools for Technology Transfer (STTT), 14(3):307–326, June 2012. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=14&issue=3&spage=307.

Huang:2012:SMC

[443] Xiaowan Huang, Justin Seyster, Sean Callanan, Ketan Dixit, Radu Grosu, Scott A. Smolka, Scott D. Stoller, and Erez Zadok. Software monitoring with controllable overhead. International Journal on Software Tools for Technology Transfer (STTT), 14 (3):327–347, June 2012. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=14&issue=3&spage=327.

Falcone:2012:WCY

[444] Yliès Falcone, Jean-Claude Fernandez, and Laurent Mounier. What can you verify and enforce at runtime? International Journal on Software Tools for Technology Transfer (STTT), 14(3):349–382, June 2012. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=14&issue=3&spage=349.

Petrenko:2012:MBT

[445] Alexandre Petrenko, Adenilso Simao, and José Carlos Maldonado. Modelbased testing of software and systems: recent advances and challenges. International Journal on Software Tools for Technology Transfer (STTT), 14 (4):383–386, August 2012. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=14&issue=4&spage=383.

Veanes:2012:ASI

[446] Margus Veanes and Nikolaj Bjørner. Alternating simulation and IOCO. International Journal on Software Tools for Technology Transfer (STTT), 14 (4):387–405, August 2012. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=14&issue=4&spage=387.

Falcone:2012:MTP

[447] Yliès Falcone, Jean-Claude Fernandez, Thierry Jéron, Hervé Marchand, and Laurent Mounier. More testable properties. International Journal on Software Tools for Technology Transfer (STTT), 14(4):407–437, August 2012. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=14&issue=4&spage=407.

Gladisch:2012:MGQ

[448] Christoph D. Gladisch. Model generation for quantified formulas with application to test data generation. International Journal on Software Tools for Technology Transfer (STTT), 14 (4):439–459, August 2012. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=1433-2779&
                                        volume=14&issue=4&spage=439.

Vergilio:2012:MOO

[449] Silvia Regina Vergilio, Aurora Pozo, João Carlos Garcia Árias, Rafael da Veiga Cabral, and Tiago Nobre. Multi-objective optimization algorithms applied to the class integration and test order problem. International Journal on Software Tools for Technology Transfer (STTT), 14(4):461–475, August 2012. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=14&issue=4&spage=461.

Schaefer:2012:SDS

[450] Ina Schaefer, Rick Rabiser, Dave Clarke, Lorenzo Bettini, David Benavides, Goetz Botterweck, Animesh Pathak, Salvador Trujillo, and Karina Villela. Software diversity: state of the art and perspectives. International Journal on Software Tools for Technology Transfer (STTT), 14(5):477–495, October 2012. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=14&issue=5&spage=477.

Pleuss:2012:VVC

[451] Andreas Pleuss and Goetz Botterweck. Visualization of variability and configuration options. International Journal on Software Tools for Technology Transfer (STTT), 14(5):497–510, October 2012. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.

asp?genre=article&issn=1433-2779&
volume=14&issue=5&spage=497.

Jorges:2012:CBV

[452] Sven Jörges, Anna-Lena Lamprecht, Tiziana Margaria, Ina Schaefer, and Bernhard Steffen. A constraint-based variability modeling framework. International Journal on Software Tools for Technology Transfer (STTT), 14 (5):511–530, October 2012. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=14&issue=5&spage=511.

Tartler:2012:RRC

[453] Reinhard Tartler, Julio Sincero, Christian Dietrich, Wolfgang Schröder-Preikschat, and Daniel Lohmann. Revealing and repairing configuration inconsistencies in large-scale system software. International Journal on Software Tools for Technology Transfer (STTT), 14(5):531–551, October 2012. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=14&issue=5&spage=531.

Heymans:2012:CTA

[454] Patrick Heymans, Quentin Boucher, Andreas Classen, Arnaud Bourdoux, and Laurent Demonceau. A code tagging approach to software product line development: an application to satellite communication libraries. International Journal on Software Tools for Technology Transfer (STTT), 14 (5):553–566, October 2012. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=14&issue=5&spage=553.

Wong:2012:ATS

[455] Peter Y. H. Wong, Elvira Albert, Radu Muschevici, José Proença, Jan Schäfer, and Rudolf Schlatte. The ABS tool suite: modelling, executing and analysing distributed adaptable object-oriented systems. International Journal on Software Tools for Technology Transfer (STTT), 14 (5):567–588, October 2012. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=14&issue=5&spage=567.

Classen:2012:MCS

[456] Andreas Classen, Maxime Cordy, Patrick Heymans, Axel Legay, and Pierre-Yves Schobbens. Model checking software product lines with SNIP. International Journal on Software Tools for Technology Transfer (STTT), 14 (5):589–612, October 2012. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=14&issue=5&spage=589.

Heider:2012:FEP

[457] Wolfgang Heider, Rick Rabiser, and Paul Grünbacher. Facilitating the evolution of products in product line engineering by capturing and replaying configuration decisions. International Journal on Software Tools for Technology Transfer (STTT), 14 (5):613–630, October 2012. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=14&issue=5&spage=613.

Liu:2012:RSV

[458] Zhiming Liu and Abhik Roychoudhury. Relating software validation to technology trends. International Journal on Software Tools for Technology Transfer (STTT), 14 (6):631–638, November 2012. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=14&issue=6&spage=631.

Pan:2012:TAM

[459] Minxue Pan and Xuandong Li. Timing analysis of MSC specifications with asynchronous concatenation. International Journal on Software Tools for Technology Transfer (STTT), 14 (6):639–651, November 2012. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&
volume=14&issue=6&spage=639.

Chen:2012:FMV

[460] Chunqing Chen, Jun Sun, Yang Liu, Jin Song Dong, and Manchun Zheng. Formal modeling and validation of Stateflow diagrams. International Journal on Software Tools for Technology Transfer (STTT), 14 (6):653–671, November 2012. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=14&issue=6&spage=653.

Gomes:2012:CMB

[461] Adriano Gomes, Alexandre Mota, Augusto Sampaio, Felipe Ferri, and Edson Watanabe. Constructive modelbased analysis for safety assessment. International Journal on Software Tools for Technology Transfer (STTT), 14 (6):673–702, November 2012. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=14&issue=6&spage=673.

David:2012:CVR

[462] Alexandre David, Kim. G. Larsen, Axel Legay, Mikael H. Møller, Ulrik Nyman, Anders P. Ravn, Arne Skou, and Andrzej Wa̧sowski. Compositional verification of real-time systems using Ecdar. International Journal on Software Tools for Technology Transfer (STTT), 14(6):703–720, November 2012. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1433-2779&
                                        volume=14&issue=6&spage=703.

Lisper:2013:ATP

[463] Björn Lisper. The ALL-TIMES project: introduction and overview. International Journal on Software Tools for Technology Transfer (STTT), 15(1): 1–8, February 2013. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10009-012-
                                        0265-7.

Merriam:2013:MTM

[464] Nicholas Merriam, Peter Gliwa, and Ian Broster. Measurement and tracing methods for timing analysis. International Journal on Software Tools for Technology Transfer (STTT), 15(1):9–28, February 2013. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-012-0266-6.

Heckmann:2013:AET

[465] Reinhold Heckmann, Christian Ferdinand, Daniel Kästner, and Stefana Nenova. Architecture exploration and timing estimation during early design phases. International Journal on Software Tools for Technology Transfer (STTT), 15(1):29–39, February 2013. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-012-0248-8.

Schreiner:2013:CTB

[466] Dietmar Schreiner, Gergö Barany, Markus Schordan, and Jens Knoop. Comparison of type-based and aliasbased component recognition for embedded systems software. International Journal on Software Tools for Technology Transfer (STTT), 15(1): 41–52, February 2013. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10009-012-
0251-0.

Lisper:2013:PEA

[467] Björn Lisper, Andreas Ermedahl, Dietmar Schreiner, Jens Knoop, and Peter Gliwa. Practical experiences of applying source-level WCET flow analysis to industrial code. International Journal on Software Tools for Technology Transfer (STTT), 15(1):53–63, February 2013. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-012-0255-9.

Merriam:2013:EPI

[468] Nicholas Merriam and Björn Lisper. Estimation of productivity increase for timing analysis tool chains. International Journal on Software Tools for Technology Transfer (STTT), 15(1): 65–84, February 2013. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10009-012-
0262-x.

Abdulla:2013:TSV

[469] Parosh Aziz Abdulla and K. Rustan M. Leino. Tools for software verification. International Journal on Software Tools for Technology Transfer (STTT), 15(2):85–88, April 2013. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-013-0270-5.

Garavel:2013:CTC

[470] Hubert Garavel, Frédéric Lang, Radu Mateescu, and Wendelin Serwe. CADP 2011: a toolbox for the construction and analysis of distributed processes. International Journal on Software Tools for Technology Transfer (STTT), 15(2):89–107, April 2013. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-012-0244-z.

Tsay:2013:BSO

[471] Yih-Kuen Tsay, Ming-Hsien Tsai, Jinn-Shu Chang, Yi-Wen Chang, and Chi-Shiang Liu. Büchi Store: an open repository of ω-automata. International Journal on Software Tools for Technology Transfer (STTT), 15(2):109–123, April 2013. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-012-0268-4.

Marques:2013:MCW

[472] Abinoam P. Marques Jr., Anders P.

Ravn, Jirí Srba, and Saleem Vighio. Model-checking Web services business activity protocols. International Journal on Software Tools for Technology Transfer (STTT), 15(2):125–147, April 2013. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-012-0231-4.

Bensalem:2013:RED

[473] Saddek Bensalem, Axel Legay, and Marius Bozga. Rigorous embedded design: challenges and perspectives. International Journal on Software Tools for Technology Transfer (STTT), 15(3):149–154, June 2013. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-013-0271-4.

Lampka:2013:CBS

[474] Kai Lampka, Simon Perathoner, and Lothar Thiele. Component-based system design: analytic real-time interfaces for state-based component implementations. International Journal on Software Tools for Technology Transfer (STTT), 15(3):155–170, June 2013. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-012-0257-7.

Rox:2013:CPA

[475] Jonas Rox and Rolf Ernst. Compositional performance analysis with improved analysis techniques for obtaining viable end-to-end latencies in distributed embedded systems. International Journal on Software Tools for Technology Transfer (STTT), 15(3): 171–187, June 2013. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10009-012-
0260-z.

Craciunas:2013:TIR

[476] Silviu S. Craciunas, Christoph M. Kirsch, Hannes Payer, Harald Röck, and Ana Sokolova. Temporal isolation in real-time systems: the VBS approach. International Journal on Software Tools for Technology Transfer (STTT), 15(3):189–209, June 2013. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-012-0246-x.

Le:2013:TAB

[477] Thi Thieu Hoa Le, Luigi Palopoli, Roberto Passerone, and Yusi Ramadian. Timed-automata based schedulability analysis for distributed firm real-time systems: a case study. International Journal on Software Tools for Technology Transfer (STTT), 15 (3):211–228, June 2013. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL http://link.springer.

com/article/10.1007/s10009-012-
0245-y.

Assayad:2013:TEB

[478] Ismail Assayad, Alain Girault, and Hamoudi Kalla. Tradeoff exploration between reliability, power consumption, and execution time for embedded systems. International Journal on Software Tools for Technology Transfer (STTT), 15(3):229–245, June 2013. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-012-0263-9.

Maler:2013:MPA

[479] Oded Maler and Dejan Nicković. Monitoring properties of analog and mixedsignal circuits. International Journal on Software Tools for Technology Transfer (STTT), 15(3):247–268, June 2013. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-012-0247-9.

Houben:2013:MTS

[480] Fred Houben, Georgeta Igna, and Frits Vaandrager. Modeling task systems using parameterized partial orders. International Journal on Software Tools for Technology Transfer (STTT), 15(3):269–286, June 2013. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-012-0264-8.

Kowalewski:2013:MCA

[481] Stefan Kowalewski, Anna Philippou, and Jörg Brauer. Model checking and abstract interpretation as building blocks of advanced program analysis techniques. International Journal on Software Tools for Technology Transfer (STTT), 15(4):287–289, August 2013. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//link.springer.com/article/10.
1007/s10009-013-0280-3;   http://
link.springer.com/content/pdf/10.
1007/s10009-013-0280-3.pdf.

Gupta:2013:TP

[482] Ashutosh Kumar Gupta, Rupak Majumdar, and Andrey Rybalchenko. From tests to proofs. International Journal on Software Tools for Technology Transfer (STTT), 15(4):291–303, August 2013. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-012-0267-5.

Plaku:2013:FLS

[483] Erion Plaku, Lydia E. Kavraki, and Moshe Y. Vardi. Falsification of LTL safety properties in hybrid systems. International Journal on Software Tools for Technology Transfer (STTT), 15(4):305–320, August 2013. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-012-0233-2.

Kahlon:2013:SAC

[484] Vineet Kahlon, Sriram Sankaranarayanan, and Aarti Gupta. Static analysis for concurrent programs with applications to data race detection. International Journal on Software Tools for Technology Transfer (STTT), 15(4): 321–336, August 2013. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10009-013-
0274-1.

Nguyen:2013:SDI

[485] Viet Yen Nguyen and Theo C. Ruys. Selected dynamic issues in software model checking. International Journal on Software Tools for Technology Transfer (STTT), 15(4):337–362, August 2013. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-012-0261-y.

Bombino:2013:MDC

[486] Massimo Bombino and Patrizia Scandurra. A model-driven co-simulation environment for heterogeneous systems. International Journal on Software Tools for Technology Transfer (STTT), 15(4):363–374, August 2013. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-012-0230-5.

Abdelhalim:2013:IFC

[487] Islam Abdelhalim, Steve Schneider, and Helen Treharne. An integrated framework for checking the behaviour of fUML models using CSP. International Journal on Software Tools for Technology Transfer (STTT), 15(4): 375–396, August 2013. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10009-012-
0243-0.

Bodik:2013:APS

[488] Rastislav Bodik and Barbara Jobstmann. Algorithmic program synthesis: introduction. International Journal on Software Tools for Technology Transfer (STTT), 15(5–6):397–411, October 2013. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-013-0287-9.

Vechev:2013:AGS

[489] Martin Vechev, Eran Yahav, and Greta Yorsh. Abstraction-guided synthesis of synchronization. International Journal on Software Tools for Technology Transfer (STTT), 15(5–6):413–431, October 2013. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-012-0232-3.

Sohail:2013:SFT

[490] Saqib Sohail and Fabio Somenzi. Safety first: a two-stage algorithm for the synthesis of reactive systems. International Journal on Software Tools for Technology Transfer (STTT), 15(5–6):433–454, October 2013. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-012-0224-3.

Kuncak:2013:FSL

[491] Viktor Kuncak, Mikaël Mayer, Ruzica Piskac, and Philippe Suter. Functional synthesis for linear arithmetic and sets. International Journal on Software Tools for Technology Transfer (STTT), 15(5–6):455–474, October 2013. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-011-0217-7.

Solar-Lezama:2013:PS

[492] Armando Solar-Lezama. Program sketching. International Journal on Software Tools for Technology Transfer (STTT), 15(5–6):475–495, October 2013. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-012-0249-7.

Srivastava:2013:TBP

[493] Saurabh Srivastava, Sumit Gulwani, and Jeffrey S. Foster. Templatebased program verification and program synthesis. International Journal on Software Tools for Technology Transfer (STTT), 15(5–6):497–518, October 2013. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-012-0223-4.

Finkbeiner:2013:BS

[494] Bernd Finkbeiner and Sven Schewe.

Bounded synthesis. International Journal on Software Tools for Technology Transfer (STTT), 15(5–6):519–539, October 2013. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10009-012-
0228-z.

Filiot:2013:ESL

[495] Emmanuel Filiot, Naiyong Jin, and Jean-François Raskin. Exploiting structure in LTL synthesis. International Journal on Software Tools for Technology Transfer (STTT), 15(5–6):541– 561, October 2013. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s10009-012-
                                        0222-5.

Konighofer:2013:DFS

[496] Robert Könighofer, Georg Hofferek, and Roderick Bloem. Debugging formal specifications: a practical approach using model-based diagnosis and counterstrategies. International Journal on Software Tools for Technology Transfer (STTT), 15(5–6):563–583, October 2013. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-011-0221-y.

Godhal:2013:SAA

[497] Yashdeep Godhal, Krishnendu Chatterjee, and Thomas A. Henzinger. Synthesis of AMBA AHB from formal specification: a case study. International Journal on Software Tools for Technology Transfer (STTT), 15(5–6):585–601, October 2013. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-011-0207-9.

Lustig:2013:SCL

[498] Yoad Lustig and Moshe Y. Vardi. Synthesis from component libraries. International Journal on Software Tools for Technology Transfer (STTT), 15 (5–6):603–618, October 2013. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-012-0236-z.

Bensalem:2014:VVM

[499] Saddek Bensalem, Klaus Havelund, and Andrea Orlandini. Verification and validation meet planning and scheduling. International Journal on Software Tools for Technology Transfer (STTT), 16(1):1–12, February 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-013-0294-x.

Goldman:2014:LAT

[500] Robert P. Goldman, Michael J. S.

Pelican, and David J. Musliner. A loop acceleration technique to speed up verification of automatically generated plans. International Journal on Software Tools for Technology Transfer (STTT), 16(1):13–29, February 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-013-0284-z.

Crampton:2014:AWS

[501] Jason Crampton, Michael Huth, and Jim Huan-Pu Kuo. Authorized workflow schemas: deciding realizability through LTL(F) model checking. International Journal on Software Tools for Technology Transfer (STTT), 16(1):31–48, February 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-012-0269-3.

Razavi:2014:GET

[502] Niloofar Razavi, Azadeh Farzan, and Sheila A. McIlraith. Generating effective tests for concurrent programs via AI automated planning techniques. International Journal on Software Tools for Technology Transfer (STTT), 16(1):49–65, February 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-013-0277-y.

Snyder:2014:OSS

[503] Jason Snyder, Deepan Seeralan, Shereef Sayed, Jeffery Wilson, Carl B. Dietrich, Stephen H. Edwards, and Jeffrey H. Reed. Open source software-defined radio tools for education, research, and rapid prototyping. International Journal on Software Tools for Technology Transfer (STTT), 16(1):67–80, February 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-012-0241-2.

Muhlberg:2014:SOC

[504] Jan Tobias Mühlberg and Gerald Lüttgen. Symbolic object code analysis. International Journal on Software Tools for Technology Transfer (STTT), 16(1):81–102, February 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-012-0256-8.

Collavizza:2014:CBB

[505] Hélène Collavizza, Nguyen Le Vinh, Olivier Ponsini, Michel Rueher, and Antoine Rollet. Constraint-based BMC: a backjumping strategy. International Journal on Software Tools for Technology Transfer (STTT), 16(1):103– 121, February 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10009-012-
                                        0258-6.

Flanagan:2014:DAV

[506] Cormac Flanagan and Barbara König. Developments in automated verification techniques. International Journal on Software Tools for Technology Transfer (STTT), 16(2):123–125, April 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-013-0297-7.

Bouajjani:2014:BPA

[507] Ahmed Bouajjani and Michael Emmi.

Bounded phase analysis of messagepassing programs. International Journal on Software Tools for Technology Transfer (STTT), 16(2):127–146, April 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-013-0276-z.

Song:2014:PMC

[508] Fu Song and Tayssir Touili. Pushdown model checking for malware detection. International Journal on Software Tools for Technology Transfer (STTT), 16(2):147–173, April 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-013-0290-1.

Cox:2014:BTP

[509] Arlen Cox, Sriram Sankaranarayanan, and Bor-Yuh Evan Chang. A bit too precise? Verification of quantized digital filters. International Journal on Software Tools for Technology Transfer (STTT), 16(2):175–190, April 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-013-0279-9.

Jiang:2014:CLV

[510] Zhihao Jiang, Miroslav Pajic, Rajeev Alur, and Rahul Mangharam. Closed-loop verification of medical devices with model abstraction and refinement. International Journal on Software Tools for Technology Transfer (STTT), 16(2):191–213, April 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-013-0289-7.

Grabowski:2014:HSR

[511] Jens Grabowski, Ina Schieferdecker, and Andreas Ulrich. History, status, and recent trends of the testing and test control notation version 3 (TTCN-3). International Journal on Software Tools for Technology Transfer (STTT), 16(3):215–225, June 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-014-0302-9.

Makedonski:2014:QET

[512] Philip Makedonski, Jens Grabowski, and Florian Philipp. Quantifying the evolution of TTCN-3 as a language. International Journal on Software Tools for Technology Transfer (STTT), 16(3):227–246, June 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-013-0282-1.

Grossmann:2014:THS

[513] Juergen Grossmann. Testing hybrid systems with TTCN-3 embedded. International Journal on Software Tools for Technology Transfer (STTT), 16(3):247–267, June 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-013-0283-0.

Stepien:2014:IEI

[514] Bernard Stepien and Liam Peyton. Innovation and evolution in integrated web application testing with TTCN-3. International Journal on Software Tools for Technology Transfer (STTT), 16(3):269–283, June 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-013-0278-x.

Zeiss:2014:CTS

[515] Benjamin Zeiss, Andras Kovacs, Nikolay Pakulin, and Bogdan Stanca-Kaposta. A conformance test suite for TTCN-3 tools. International Journal on Software Tools for Technology Transfer (STTT), 16(3):285–294, June 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-013-0285-y.

Rings:2014:GIT

[516] Thomas Rings, Patrick Poglitsch, Stephan Schulz, Luca Serazio, and Theofanis Vassiliou-Gioles. A generic interoperability testing framework and a systematic development process for automated interoperability testing. International Journal on Software Tools for Technology Transfer (STTT), 16(3):295–313, June 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-013-0281-2.

Alnusair:2014:RBD

[517] Awny Alnusair, Tian Zhao, and Gongjun Yan. Rule-based detection of design patterns in program code. International Journal on Software Tools for Technology Transfer (STTT), 16(3):315–334, June 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-013-0292-z.

Margaria:2014:PVT

[518] Tiziana Margaria, Zongyan Qiu, and Hongli Yang. Program verification and testing technologies. International Journal on Software Tools for Technology Transfer (STTT), 16(4): 335–337, August 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10009-014-
                                        0327-0.

Fang:2014:FVS

[519] Huixing Fang, Jianqi Shi, Huibiao Zhu, Jian Guo, Kim Guldstrand Larsen, and Alexandre David. Formal verification and simulation for platform screen doors and collision avoidance in subway control systems. International Journal on Software Tools for Technology Transfer (STTT), 16(4):339–361, August 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-014-0318-1.

Gherghina:2014:EPV

[520] Cristian Gherghina, Cristina David, Shengchao Qin, and Wei-Ngan Chin. Expressive program verification via structured specifications. International Journal on Software Tools for Technology Transfer (STTT), 16(4):363– 380, August 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10009-014-
0306-5.

Ferreira:2014:AVF

[521] João F. Ferreira, Cristian Gherghina, Guanhua He, Shengchao Qin, and Wei-Ngan Chin. Automated verification of the FreeRTOS scheduler in Hip/Sleek. International Journal on Software Tools for Technology Transfer (STTT), 16(4):381–397, August 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-014-0307-4.

Jarraya:2014:QQA

[522] Yosr Jarraya and Mourad Debbabi.

Quantitative and qualitative analysis of SysML activity diagrams. International Journal on Software Tools for Technology Transfer (STTT), 16(4): 399–419, August 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10009-014-
0305-6.

Yuksel:2014:QMA

[523] Ender Yüksel, Hanne Riis Nielson, Flemming Nielson, Huibiao Zhu, and Heqing Huang. Quantitative modelling and analysis of a Chinese smart grid: a stochastic model checking case study. International Journal on Software Tools for Technology Transfer (STTT), 16(4):421–435, August 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-014-0311-8.

Guo:2014:MBT

[524] Hai-Feng Guo and Mahadevan Subramaniam. Model-based test generation using extended symbolic grammars. International Journal on Software Tools for Technology Transfer (STTT), 16(4):437–455, August 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-014-0316-3.

Howar:2014:RER

[525] Falk Howar, Malte Isberner, Maik Merten, Bernhard Steffen, Dirk Beyer, and Corina S. Pasareanu. Rigorous examination of reactive systems. International Journal on Software Tools for Technology Transfer (STTT), 16(5):457–464, October 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-014-0337-y.

Steffen:2014:PDB

[526] Bernhard Steffen, Malte Isberner, Stefan Naujokat, Tiziana Margaria, and Maren Geske. Property-driven benchmark generation: synthesizing programs of realistic structure. International Journal on Software Tools for Technology Transfer (STTT), 16(5):465–479, October 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-014-0336-z.

vandePol:2014:TBF

[527] Jaco van de Pol, Theo C. Ruys, and Steven te Brinke. Thoughtful bruteforce attack of the RERS 2012 and 2013 Challenges. International Journal on Software Tools for Technology Transfer (STTT), 16(5):481–491, October 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-014-0324-3.

Schordan:2014:CSA

[528] Markus Schordan and Adrian Prantl.

Combining static analysis and state transition graphs for verification of event-condition-action systems in the RERS 2012 and 2013 Challenges. International Journal on Software Tools for Technology Transfer (STTT), 16(5): 493–505, October 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10009-014-
0338-x.

Beyer:2014:BBS

[529] Dirk Beyer and Andreas Stahlbauer.

BDD-based software verification. International Journal on Software Tools for Technology Transfer (STTT), 16(5): 507–518, October 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10009-014-
0334-1.

Morse:2014:ASB

[530] Jeremy Morse, Lucas Cordeiro, Denis Nicole, and Bernd Fischer. Applying symbolic bounded model checking to the 2012 RERS Greybox Challenge. International Journal on Software Tools for Technology Transfer (STTT), 16(5):519–529, October 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-014-0335-0.

Bauer:2014:APB

[531] Oliver Bauer, Maren Geske, and Malte Isberner. Analyzing program behavior through active automata learning. International Journal on Software Tools for Technology Transfer (STTT), 16(5):531–542, October 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-014-0333-2.

Steffen:2014:TGC

[532] Bernhard Steffen, Falk Howar, Malte Isberner, Stefan Naujokat, and Tiziana Margaria. Tailored generation of concurrent benchmarks. International Journal on Software Tools for Technology Transfer (STTT), 16(5):543–558, October 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-014-0339-9.

Felderer:2014:TRB

[533] Michael Felderer and Ina Schieferdecker. A taxonomy of risk-based testing. International Journal on Software Tools for Technology Transfer (STTT), 16(5):559–568, October 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-014-0332-3.

Neubauer:2014:RBT

[534] Johannes Neubauer, Stephan Windmüller, and Bernhard Steffen. Risk-based testing via active continuous quality control. International Journal on Software Tools for Technology Transfer (STTT), 16(5):569–591, October 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-014-0321-6.

Carrozza:2014:DTP

[535] Gabriella Carrozza, Roberto Pietrantuono, and Stefano Russo. Dynamic test planning: a study in an industrial context. International Journal on Software Tools for Technology Transfer (STTT), 16(5):593–607, October 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-014-0319-0.

Felderer:2014:MCS

[536] Michael Felderer and Rudolf Ramler. A multiple case study on risk-based testing in industry. International Journal on Software Tools for Technology Transfer (STTT), 16(5):609–625, October 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-014-0328-z.

Erdogan:2014:ACU

[537] Gencer Erdogan, Yan Li, Ragnhild Kobro Runde, Fredrik Seehusen, and Ketil Stølen. Approaches for the combined use of risk analysis and testing: a systematic literature review. International Journal on Software Tools for Technology Transfer (STTT), 16(5):627–642, October 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-014-0330-5.

Fantechi:2014:FMR

[538] Alessandro Fantechi, Francesco Flammini, and Stefania Gnesi. Formal methods for railway control systems. International Journal on Software Tools for Technology Transfer (STTT), 16(6):643–646, November 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-014-0342-1.

Ferrari:2014:CDS

[539] Alessio Ferrari, Giorgio O. Spagnolo, Giacomo Martelli, and Simone Menabeni. From commercial documents to system requirements: an approach for the engineering of novel CBTC solutions. International Journal on Software Tools for Technology Transfer (STTT), 16(6):647–667, November 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-013-0298-6.

Marrone:2014:TMD

[540] Stefano Marrone, Francesco Flammini, Nicola Mazzocca, Roberto Nardone, and Valeria Vittorini. Towards modeldriven V&V assessment of railway control systems. International Journal on Software Tools for Technology Transfer (STTT), 16(6):669–683, November 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-014-0320-7.

James:2014:TMV

[541] Phillip James, Faron Moller, Hoang Nga Nguyen, Markus Roggenbach, Steve Schneider, and Helen Treharne. Techniques for modelling and verifying railway interlockings. International Journal on Software Tools for Technology Transfer (STTT), 16(6):685–711, November 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-014-0304-7.

Haxthausen:2014:AGF

[542] Anne E. Haxthausen. Automated generation of formal safety conditions from railway interlocking tables. International Journal on Software Tools for Technology Transfer (STTT), 16(6):713–726, November 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-013-0295-9.

Galler:2014:STD

[543] Stefan J. Galler and Bernhard K. Aichernig. Survey on test data generation tools. International Journal on Software Tools for Technology Transfer (STTT), 16(6):727–751, November 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-013-0272-3.

Quer:2014:MCE

[544] Stefano Quer. Model checking evaluation of airplane landing trajectories. International Journal on Software Tools for Technology Transfer (STTT), 16(6):753–773, November 2014. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-013-0273-2.

Nilsson:2015:AEI

[545] Agneta Nilsson, Laura M. Castro, Samuel Rivas, and Thomas Arts. Assessing the effects of introducing a new software development process: a methodological description. International Journal on Software Tools for Technology Transfer (STTT), 17(1):1–16, February 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-013-0275-0.

David:2015:RTS

[546] Alexandre David, Kim G. Larsen, Axel Legay, Ulrik Nyman, Louis-Marie Traonouez, and Andrzej Wasowski. Real-time specifications. International Journal on Software Tools for Technology Transfer (STTT), 17(1): 17–45, February 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10009-013-
                                        0286-x.

Chen:2015:TPL

[547] Hong Yi Chen, Shaked Flur, and Supratik Mukhopadhyay. Termination proofs for linear simple loops. International Journal on Software Tools for Technology Transfer (STTT), 17(1): 47–57, February 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10009-013-
                                        0288-8.

Shafique:2015:SRS

[548] Muhammad Shafique and Yvan Labiche. A systematic review of state-based test tools. International Journal on Software Tools for Technology Transfer (STTT), 17(1):59–76, February 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-013-0291-0.

Pulungan:2015:CMS

[549] Reza Pulungan and Holger Hermanns.

A construction and minimization service for continuous probability distributions. International Journal on Software Tools for Technology Transfer (STTT), 17(1):77–90, February 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-013-0296-8.

Lincke:2015:FPG

[550] Daniel Lincke, Sibylle Schupp, and Cezar Ionescu. Functional prototypes for generic C++ libraries: a transformational approach based on higher-order, typed signatures. International Journal on Software Tools for Technology Transfer (STTT), 17(1):91–105, February 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-014-0299-0.

Wong:2015:TAB

[551] Peter Y. H. Wong, Richard Bubel, Frank S. de Boer, Miguel Gómez-Zamalloa, Stijn de Gouw, Reiner Hähnle, Karl Meinke, and Muddassar Azam Sindhu. Testing abstract behavioral specifications. International Journal on Software Tools for Technology Transfer (STTT), 17(1):107– 119, February 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s10009-014-
                                        0301-x.

Falcone:2015:RVA

[552] Yliès Falcone and Lenore D. Zuck. Runtime verification: the application perspective. International Journal on Software Tools for Technology Transfer (STTT), 17(2):121–123, April 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-014-0360-z.

Halle:2015:PRM

[553] Sylvain Hallé, Jason Vallet, and Raphaël Tremblay-Lessard. On piggyback runtime monitoring of object-oriented programs. International Journal on Software Tools for Technology Transfer (STTT), 17(2):125–142, April 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-014-0326-1.

Havelund:2015:RBR

[554] Klaus Havelund. Rule-based runtime verification revisited. International Journal on Software Tools for Technology Transfer (STTT), 17(2): 143–170, April 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10009-014-
                                        0309-2.

Nouri:2015:SMC

[555] Ayoub Nouri, Saddek Bensalem, Marius Bozga, Benoit Delahaye, Cyrille Jegourel, and Axel Legay. Statistical model checking QoS properties of systems with SBIP. International Journal on Software Tools for Technology Transfer (STTT), 17(2):171–185, April 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-014-0313-6.

David:2015:SHR

[556] Alexandre David, Kim G. Larsen, Axel Legay, and Marius Mikucionis. Schedulability of Herschel revisited using statistical model checking. International Journal on Software Tools for Technology Transfer (STTT), 17(2):187–199, April 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-014-0331-4.

Salva:2015:AAA

[557] Sébastien Salva and Stassia R. Zafimiharisoa. APSET, an Android aPplication SEcurity Testing tool for detecting intent-based vulnerabilities. International Journal on Software Tools for Technology Transfer (STTT), 17(2): 201–221, April 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10009-014-
0303-8.

Wang:2015:MCF

[558] Farn Wang. Model-checking fair densetime systems with propositions and events. International Journal on Software Tools for Technology Transfer (STTT), 17(2):223–243, April 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-014-0312-7.

Felderer:2015:PMS

[559] Michael Felderer and Basel Katt. A process for mastering security evolution in the development lifecycle. International Journal on Software Tools for Technology Transfer (STTT), 17(3):245–250, June 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-015-0371-4.

Refsdal:2015:SRA

[560] Atle Refsdal, Bjørnar Solhaug, and Ketil Stølen. Security risk analysis of system changes exemplified within the oil and gas domain. International Journal on Software Tools for Technology Transfer (STTT), 17(3):251–266, June 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-014-0351-0.

Burger:2015:RSE

[561] Jens Bürger, Jan Jürjens, and Sven Wenzel. Restoring security of evolving software models using graph transformation. International Journal on Software Tools for Technology Transfer (STTT), 17(3):267–289, June 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-014-0364-8.

Vanoverberghe:2015:PIC

[562] Dries Vanoverberghe and Frank Piessens. Policy ignorant caller-side inline reference monitoring. International Journal on Software Tools for Technology Transfer (STTT), 17(3):291–303, June 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-014-0348-8.

Felderer:2015:SCS

[563] Michael Felderer and Elizabeta Fourneret.

A systematic classification of security regression testing approaches. International Journal on Software Tools for Technology Transfer (STTT), 17(3): 305–319, June 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10009-015-
0365-2.

Turner:2015:WQD

[564] Kenneth J. Turner and Paul S.

Lambert. Workflows for quantitative data analysis in the social sciences. International Journal on Software Tools for Technology Transfer (STTT), 17(3):321–338, June 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-014-0315-4.

Anielak:2015:ITC

[565] Grzegorz Anielak, Grzegorz Jakacki, and Slawomir Lasota. Incremental test case generation using bounded model checking: an application to automatic rating. International Journal on Software Tools for Technology Transfer (STTT), 17(3):339–349, June 2015. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//link.springer.com/article/10.
1007/s10009-014-0317-2;   http://
link.springer.com/content/pdf/10.
1007/s10009-014-0317-2.pdf.

David:2015:SMC

[566] Alexandre David, Kim G. Larsen, Axel Legay, Marius Mikucionis, Danny Bøgsted Poulsen, and Sean Sedwards. Statistical model checking for biological systems. International Journal on Software Tools for Technology Transfer (STTT), 17(3):351–367, June 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-014-0323-4.

Legay:2015:SMC

[567] Axel Legay and Mahesh Viswanathan. Statistical model checking: challenges and perspectives. International Journal on Software Tools for Technology Transfer (STTT), 17(4):369–376, August 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-015-0384-z.

Reijsbergen:2015:HTS

[568] Daniël Reijsbergen, Pieter-Tjerk de Boer, Werner Scheinhardt, and Boudewijn Haverkort. On hypothesis testing for statistical model checking. International Journal on Software Tools for Technology Transfer (STTT), 17(4): 377–395, August 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10009-014-
                                        0350-1.

David:2015:UST

[569] Alexandre David, Kim G. Larsen, Axel Legay, Marius Mikucionis, and Danny Bøgsted Poulsen. Uppaal SMC tutorial. International Journal on Software Tools for Technology Transfer (STTT), 17(4):397–415, August 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-014-0361-y.

Roohi:2015:SMC

[570] Nima Roohi and Mahesh Viswanathan.

Statistical model checking for unbounded until formulas. International Journal on Software Tools for Technology Transfer (STTT), 17(4):417– 427, August 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10009-015-
0368-z.

Hartmanns:2015:SSM

[571] Arnd Hartmanns and Mark Timmer.

Sound statistical model checking for MDP using partial order and confluence reduction. International Journal on Software Tools for Technology Transfer (STTT), 17(4):429–456, August 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-014-0349-7.

Lassaigne:2015:APV

[572] Richard Lassaigne and Sylvain Peyronnet. Approximate planning and verification for large Markov decision processes. International Journal on Software Tools for Technology Transfer (STTT), 17(4):457–467, August 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-014-0344-z.

DArgenio:2015:SSL

[573] Pedro D’Argenio, Axel Legay, Sean Sedwards, and Louis-Marie Traonouez. Smart sampling for lightweight verification of Markov decision processes. International Journal on Software Tools for Technology Transfer (STTT), 17(4):469–484, August 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-015-0383-0.

Ellen:2015:SMC

[574] Christian Ellen, Sebastian Gerwinn, and Martin Fränzle. Statistical model checking for stochastic hybrid systems involving nondeterminism over continuous domains. International Journal on Software Tools for Technology Transfer (STTT), 17(4):485–504, August 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-014-0329-y.

Ballarini:2015:AOT

[575] Paolo Ballarini. Analysing oscillatory trends of discrete-state stochastic processes through HASL statistical model checking. International Journal on Software Tools for Technology Transfer (STTT), 17(4):505–526, August 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-015-0370-5.

Zuliani:2015:SMC

[576] Paolo Zuliani. Statistical model checking for biological applications. International Journal on Software Tools for Technology Transfer (STTT), 17(4):

527–536, August 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10009-014-
0343-0.

Chakraborty:2015:MSM

[577] Souymodip Chakraborty, Joost-Pieter Katoen, Falak Sher, and Martin Strelec. Modelling and statistical model checking of a microgrid. International Journal on Software Tools for Technology Transfer (STTT), 17(4):537–554, August 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-014-0345-y.

Gnesi:2015:SSI

[578] Stefania Gnesi and Stan Jarzabek.

Special section on the 17th International Software Product Line Conference. International Journal on Software Tools for Technology Transfer (STTT), 17(5):555–557, October 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-015-0386-x.

Eichelberger:2015:MDS

[579] Holger Eichelberger and Klaus Schmid.

Mapping the design-space of textual variability modeling languages: a refined analysis. International Journal on Software Tools for Technology Transfer (STTT), 17(5):559–584, October 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-014-0362-x.

Filho:2015:GCM

[580] João Bosco Ferreira Filho and Olivier Barais. Generating counterexamples of model-based software product lines. International Journal on Software Tools for Technology Transfer (STTT), 17(5):585–600, October 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-014-0341-2.

Haber:2015:SSD

[581] Arne Haber and Katrin Hölldobler. Systematic synthesis of delta modeling languages. International Journal on Software Tools for Technology Transfer (STTT), 17(5):601–626, October 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-015-0387-9.

Rubin:2015:CPV

[582] Julia Rubin and Krzysztof Czarnecki. Cloned product variants: from adhoc to managed software product lines. International Journal on Software Tools for Technology Transfer (STTT), 17(5):627–646, October 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-014-0347-9.

Huisman:2015:V

[583] Marieke Huisman and Vladimir Klebanov. VerifyThis 2012. International Journal on Software Tools for Technology Transfer (STTT), 17(6):647– 657, November 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s10009-015-
0396-8.

Jacobs:2015:SVC

[584] Bart Jacobs, Jan Smans, and Frank Piessens. Solving the VerifyThis 2012 challenges with VeriFast. International Journal on Software Tools for Technology Transfer (STTT), 17(6):659– 676, November 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10009-014-
0310-9.

Ernst:2015:KOV

[585] Gidon Ernst and Jörg Pfähler. KIV:

overview and VerifyThis competition. International Journal on Software Tools for Technology Transfer (STTT), 17(6):677–694, November 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-014-0308-3.

Hoang:2015:SG

[586] Duc Hoang, Yannick Moy, and Angela Wallenburg. SPARK 2014 and GNATprove. International Journal on Software Tools for Technology Transfer (STTT), 17(6):695–707, November 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-014-0322-5.

Bobot:2015:LVW

[587] François Bobot and Jean-Christophe Filliâtre. Let’s verify this with Why3. International Journal on Software Tools for Technology Transfer (STTT), 17(6):709–727, November 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-014-0314-5.

Bruns:2015:ILV

[588] Daniel Bruns and Wojciech Mostowski. Implementation-level verification of algorithms with KeY. International Journal on Software Tools for Technology Transfer (STTT), 17(6):729– 744, November 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10009-013-
                                        0293-y.

Tschannen:2015:AMS

[589] Julian Tschannen and Carlo A. Furia. AutoProof meets some verification challenges. International Journal on Software Tools for Technology Transfer (STTT), 17(6):745–755, November 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-014-0300-y.

Blom:2015:WEM

[590] Stefan Blom and Marieke Huisman. Witnessing the elimination of magic wands. International Journal on Software Tools for Technology Transfer (STTT), 17(6):757– 781, November 2015. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10009-015-0372-3;   http://
                                        link.springer.com/content/pdf/10.
                                        1007/s10009-015-0372-3.pdf.

Pezze:2016:MDG

[591] Mauro Pezzé and Jochen Wuttke.

Model-driven generation of runtime checks for system properties. International Journal on Software Tools for Technology Transfer (STTT), 18(1): 1–19, February 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10009-014-
0325-2.

Hendriks:2016:BSL

[592] Martijn Hendriks and Twan Basten.

A blueprint for system-level performance modeling of software-intensive embedded systems. International Journal on Software Tools for Technology Transfer (STTT), 18(1):21–40, February 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-014-0340-3.

Wehrle:2016:DPR

[593] Martin Wehrle and Sebastian Kupferschmid. Downward pattern refinement for timed automata. International Journal on Software Tools for Technology Transfer (STTT), 18 (1):41–56, February 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//link.springer.com/article/10.
1007/s10009-014-0346-x;   http://
link.springer.com/content/pdf/10.
1007/s10009-014-0346-x.pdf.

Inoue:2016:GSL

[594] Takeru Inoue, Hiroaki Iwashita, and Jun Kawahara. Graphillion: software library for very large sets of labeled graphs. International Journal on Software Tools for Technology Transfer (STTT), 18(1):57–66, February 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10009-014-0352-z;   http://
                                        link.springer.com/content/pdf/10.
                                        1007/s10009-014-0352-z.pdf.

Quesel:2016:HMP

[595] Jan-David Quesel, Stefan Mitsch, and Sarah Loos. How to model and prove hybrid systems with KeYmaera: a tutorial on safety. International Journal on Software Tools for Technology Transfer (STTT), 18(1):67–91, February 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10009-015-0367-0. See cor-

rection [837].

Osaiweran:2016:EEL

[596] Ammar Osaiweran and Mathijs Schuts. Evaluating the effect of a lightweight formal technique in industry. International Journal on Software Tools for Technology Transfer (STTT), 18 (1):93–108, February 2016. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10009-015-0374-1;   http://
                                        link.springer.com/content/pdf/10.
                                        1007/s10009-015-0374-1.pdf.

Kutsuna:2016:ARM

[597] Takuro Kutsuna and Yoshinao Ishii. Abstraction and refinement of mathematical functions toward SMT-based testcase generation. International Journal on Software Tools for Technology Transfer (STTT), 18(1):109–120, February 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-015-0389-7.

Abraham:2016:SRA

[598] Erika Ábrahám and Klaus Havelund.

Some recent advances in automated analysis. International Journal on Software Tools for Technology Transfer (STTT), 18(2):121–128, April 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-015-0403-0.

Lowe:2016:CDF

[599] Gavin Lowe. Concurrent depth-first search algorithms based on Tarjan’s algorithm. International Journal on Software Tools for Technology Transfer (STTT), 18(2):129–147, April 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-015-0382-1.

Gibson-Robinson:2016:FPR

[600] Thomas Gibson-Robinson and Philip Armstrong. FDR3: a parallel refinement checker for CSP. International Journal on Software Tools for Technology Transfer (STTT), 18(2):149–167, April 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-015-0377-y.

Wijs:2016:MCF

[601] Anton Wijs and Dragan Bosnacki.

Many-core on-the-fly model checking of safety properties using GPUs. International Journal on Software Tools for Technology Transfer (STTT), 18 (2):169–185, April 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//link.springer.com/article/10.
1007/s10009-015-0379-9;   http://
link.springer.com/content/pdf/10.
1007/s10009-015-0379-9.pdf.

Armando:2016:SSB

[602] Alessandro Armando and Roberto Carbone. SATMC: a SAT-based model checker for security protocols, business processes, and security APIs. International Journal on Software Tools for Technology Transfer (STTT), 18(2):187–204, April 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-015-0385-y.

Decker:2016:MMT

[603] Normann Decker, Martin Leucker, and Daniel Thoma. Monitoring modulo theories. International Journal on Software Tools for Technology Transfer (STTT), 18(2):205–225, April 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-015-0380-3.

vonEssen:2016:PVS

[604] Christian von Essen and Dimitra Giannakopoulou. Probabilistic verification and synthesis of the next generation airborne collision avoidance system. International Journal on Software Tools for Technology Transfer (STTT), 18(2):227–243, April 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-015-0388-8.

Yenigun:2016:ATG

[605] Hüsnü Yenigün, Cemal Yilmaz, and Andreas Ulrich. Advances in test generation for testing software and systems. International Journal on Software Tools for Technology Transfer (STTT), 18(3):245–249, June 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-015-0404-z.

Kushik:2016:AEN

[606] Natalia Kushik, Khaled El-Fakih, Nina Yevtushenko, and Ana R. Cavalli. On adaptive experiments for nondeterministic finite state machines. International Journal on Software Tools for Technology Transfer (STTT), 18(3): 251–264, June 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10009-014-
0357-7.

Huang:2016:CMB

[607] Wen ling Huang and Jan Peleska. Complete model-based equivalence class testing. International Journal on Software Tools for Technology Transfer (STTT), 18(3):265–283, June 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-014-0356-8.

Faria:2016:TCT

[608] João Pascoal Faria and Ana C. R.

Paiva. A toolset for conformance testing against UML sequence diagrams based on event-driven colored Petri nets. International Journal on Software Tools for Technology Transfer (STTT), 18(3):285–304, June 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-014-0354-x.

deLeon:2016:MBT

[609] Hernán Ponce de León, Stefan Haar, and Delphine Longuet. Model-based testing for concurrent systems: unfoldingbased test selection. International Journal on Software Tools for Technology Transfer (STTT), 18(3):305–318, June 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-014-0353-y.

Schrammel:2016:GTC

[610] Peter Schrammel, Tom Melham, and Daniel Kroening. Generating test case chains for reactive systems. International Journal on Software Tools for Technology Transfer (STTT), 18 (3):319–334, June 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//link.springer.com/article/10.
1007/s10009-014-0358-6;   http://
link.springer.com/content/pdf/10.
1007/s10009-014-0358-6.pdf.

Enoiu:2016:ATG

[611] Eduard P. Enoiu, Adnan Causević, Thomas J. Ostrand, Elaine J. Weyuker, Daniel Sundmark, and Paul Pettersson. Automated test generation using model checking: an industrial evaluation. International Journal on Software Tools for Technology Transfer (STTT), 18(3):335–353, June 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.com/article/
10.1007/s10009-014-0355-9.

Bartocci:2016:PSI

[612] Ezio Bartocci and C. R. Ramakrishnan. Preface of the special issue on model checking of software. International Journal on Software Tools for Technology Transfer (STTT), 18 (4):355–357, August 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//link.springer.com/article/10.
1007/s10009-016-0414-5;   http://
link.springer.com/content/pdf/10.
1007/s10009-016-0414-5.pdf.

Lopes:2016:AEC

[613] Nuno P. Lopes and José Monteiro. Automatic equivalence checking of programs with uninterpreted functions and integer arithmetic. International Journal on Software Tools for Technology Transfer (STTT), 18(4):359–374, August 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10009-015-0366-1;   http://link.
springer.com/article/10.1007/s10009-
015-0366-1.

Sethi:2016:MCU

[614] Divjyot Sethi, Muralidhar Talupur, and Sharad Malik. Model checking unbounded concurrent lists. International Journal on Software Tools for Technology Transfer (STTT), 18(4): 375–391, August 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10009-015-0369-y;   http://link.

                                        springer.com/article/10.1007/s10009-
                                        015-0369-y.

Adhikari:2016:VQR

[615] Kiran Adhikari, James Street, Chao Wang, Yang Liu, and Shaojie Zhang. Verifying a quantitative relaxation of linearizability via refinement. International Journal on Software Tools for Technology Transfer (STTT), 18(4): 393–407, August 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s10009-015-0373-2;   http://link.
                                        springer.com/article/10.1007/s10009-
                                        015-0373-2.

Jensen:2016:EMC

[616] Jonas Finnemann Jensen, Kim Guldstrand Larsen, Jirí Srba, and Lars Kaerlund Oestergaard. Efficient modelchecking of weighted CTL with upperbound constraints. International Journal on Software Tools for Technology Transfer (STTT), 18(4):409–426, August 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s10009-014-0359-5;   http://link.
                                        springer.com/article/10.1007/s10009-
                                        014-0359-5.

Laarman:2016:GBP

[617] Alfons Laarman, Elwin Pater, Jaco van de Pol, and Henri Hansen. Guardbased partial-order reduction. International Journal on Software Tools for Technology Transfer (STTT), 18(4): 427–448, August 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/

s10009-014-0363-9;   http://link.
springer.com/article/10.1007/s10009-
014-0363-9.

Bogomolov:2016:GSH

[618] Sergiy Bogomolov, Alexandre Donzé, Goran Frehse, Radu Grosu, Taylor T. Johnson, Hamed Ladan, Andreas Podelski, and Martin Wehrle. Guided search for hybrid systems based on coarse-grained space abstractions. International Journal on Software Tools for Technology Transfer (STTT), 18 (4):449–467, August 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//link.springer.com/article/10.
1007/s10009-015-0393-y;   http://
link.springer.com/content/pdf/10.
1007/s10009-015-0393-y.pdf.

Abdulla:2016:PV

[619] Parosh A. Abdulla and Giorgio Delzanno. Parameterized verification. International Journal on Software Tools for Technology Transfer (STTT), 18(5): 469–473, October 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10009-016-0424-3;   http://link.
springer.com/article/10.1007/s10009-
016-0424-3.

Delzanno:2016:UVP

[620] Giorgio Delzanno. A unified view of parameterized verification of abstract models of broadcast communication. International Journal on Software Tools for Technology Transfer (STTT), 18(5): 475–493, October 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/

                                        s10009-016-0412-7;   http://link.
                                        springer.com/article/10.1007/s10009-
                                        016-0412-7.

Abdulla:2016:PVT

[621] Parosh Abdulla, Frédéric Haziza, and Lukás Holík. Parameterized verification through view abstraction. International Journal on Software Tools for Technology Transfer (STTT), 18(5): 495–516, October 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s10009-015-0406-x;   http://link.
                                        springer.com/article/10.1007/s10009-
                                        015-0406-x.

Ganjei:2016:CDS

[622] Zeinab Ganjei, Ahmed Rezine, Petru Eles, and Zebo Peng. Counting dynamically synchronizing processes. International Journal on Software Tools for Technology Transfer (STTT), 18(5): 517–534, October 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s10009-015-0411-0;   http://link.
                                        springer.com/article/10.1007/s10009-
                                        015-0411-0.

Montali:2016:SDA

[623] Marco Montali and Diego Calvanese. Soundness of data-aware, case-centric processes. International Journal on Software Tools for Technology Transfer (STTT), 18(5):535–558, October 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s10009-016-0417-2;   http://link.

springer.com/article/10.1007/s10009-
016-0417-2.

Ranise:2016:PMC

[624] Silvio Ranise, Anh Truong, and Riccardo Traverso. Parameterized model checking for security policy analysis. International Journal on Software Tools for Technology Transfer (STTT), 18(5): 559–573, October 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10009-015-0410-1;   http://link.
springer.com/article/10.1007/s10009-
015-0410-1.

Lamprecht:2016:SW

[625] Anna-Lena Lamprecht and Kenneth J.

Turner. Scientific workflows. International Journal on Software Tools for Technology Transfer (STTT), 18(6): 575–580, November 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10009-016-0428-z;   http://link.
springer.com/article/10.1007/s10009-
016-0428-z.

Menager:2016:URI

[626] Hervé Ménager, Matús Kalas, Kristoffer Rapacki, and Jon Ison. Using registries to integrate bioinformatics tools and services into workbench environments. International Journal on Software Tools for Technology Transfer (STTT), 18 (6):581–586, November 2016. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//link.springer.com/article/10.
1007/s10009-015-0392-z;   http://
link.springer.com/content/pdf/10.
1007/s10009-015-0392-z.pdf.

Sinnott:2016:SCB

[627] R. O. Sinnott and W. Voorsluys. A scalable cloud-based system for dataintensive spatial analysis. International Journal on Software Tools for Technology Transfer (STTT), 18(6):587– 605, November 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s10009-015-0398-6;   http://link.
                                        springer.com/article/10.1007/s10009-
                                        015-0398-6.

Bolt:2016:SWP

[628] Alfredo Bolt, Massimiliano de Leoni, and Wil M. P. van der Aalst. Scientific workflows for process mining: building blocks, scenarios, and implementation. International Journal on Software Tools for Technology Transfer (STTT), 18(6):607–628, November 2016. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10009-015-0399-5;   http://
                                        link.springer.com/content/pdf/10.
                                        1007/s10009-015-0399-5.pdf.

Lamprecht:2016:SWJ

[629] Anna-Lena Lamprecht, Bernhard Steffen, and Tiziana Margaria. Scientific workflows with the jABC framework. International Journal on Software Tools for Technology Transfer (STTT), 18(6): 629–651, November 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s10009-016-0427-0;   http://link.
                                        springer.com/article/10.1007/s10009-
                                        016-0427-0.

Amighi:2016:PCC

[630] Afshin Amighi, Pedro de Carvalho Gomes, Dilian Gurov, and Marieke Huisman. Provably correct control flow graphs from Java bytecode programs with exceptions. International Journal on Software Tools for Technology Transfer (STTT), 18(6):653–684, November 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10009-015-0375-0;   http://link.
springer.com/article/10.1007/s10009-
015-0375-0.

Gupta:2016:ECV

[631] Amar Kumar Gupta and Guy Edward Gallasch. Equivalence class verification of the contract net protocol-extension. International Journal on Software Tools for Technology Transfer (STTT), 18(6): 685–706, November 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10009-015-0376-z;   http://link.
springer.com/article/10.1007/s10009-
015-0376-z.

Steffen:2017:PST

[632] Bernhard Steffen. The physics of software tools: SWOT analysis and vision. International Journal on Software Tools for Technology Transfer (STTT), 19(1):1–7, February 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10009-016-0446-x;   http://link.
springer.com/article/10.1007/s10009-
016-0446-x.

Lomuscio:2017:MOS

[633] Alessio Lomuscio, Hongyang Qu, and Franco Raimondi. MCMAS: an opensource model checker for the verification of multi-agent systems. International Journal on Software Tools for Technology Transfer (STTT), 19 (1):9–30, February 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//link.springer.com/article/10.
1007/s10009-015-0378-x;   http://
link.springer.com/content/pdf/10.
1007/s10009-015-0378-x.pdf.

Rivera:2017:CGE

[634] Víctor Rivera, Néstor Cataño, Tim Wahls, and Camilo Rueda. Code generation for Event-B. International Journal on Software Tools for Technology Transfer (STTT), 19(1):31–52, February 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10009-015-0381-2;   http://link.
springer.com/article/10.1007/s10009-
015-0381-2.

Damasceno:2017:TRT

[635] Adriana C. Damasceno, Patricia D. L. Machado, and Wilkerson L. Andrade. Testing real-time systems from compositional symbolic specifications. International Journal on Software Tools for Technology Transfer (STTT), 19(1): 53–71, February 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10009-015-0390-1;   http://link.
springer.com/article/10.1007/s10009-
015-0390-1.

Ye:2017:MCS

[636] Kangfeng Ye and Jim Woodcock. Model checking of state-rich formalism by linking to CSP ∥ B. International Journal on Software Tools for Technology Transfer (STTT), 19(1):73–96, February 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10009-015-0402-1;   http://link.
springer.com/article/10.1007/s10009-
015-0402-1.

Gadelha:2017:HLB

[637] Mikhail Y. R. Gadelha, Hussama I. Ismail, and Lucas C. Cordeiro. Handling loops in bounded model checking of C programs via k-induction. International Journal on Software Tools for Technology Transfer (STTT), 19(1): 97–114, February 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10009-015-0407-9;   http://link.
springer.com/article/10.1007/s10009-
015-0407-9.

Zech:2017:MBR

[638] Philipp Zech, Philipp Kalb, Michael Felderer, Colin Atkinson, and Ruth Breu. Model-based regression testing by OCL. International Journal on Software Tools for Technology Transfer (STTT), 19(1):115–131, February 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10009-015-0408-8;   http://link.
springer.com/article/10.1007/s10009-
015-0408-8.

Boniol:2017:LGC

[639] Frédéric Boniol, Virginie Wiels, Yamine Aït-Ameur, and Klaus-Dieter Schewe. The landing gear case study: challenges and experiments. International Journal on Software Tools for Technology Transfer (STTT), 19(2):133–140, April 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10009-016-0431-4;   http://
                                        link.springer.com/content/pdf/10.
                                        1007/s10009-016-0431-4.pdf.

Su:2017:ALG

[640] Wen Su and Jean-Raymond Abrial. Aircraft landing gear system: approaches with Event-B to the modeling of an industrial system. International Journal on Software Tools for Technology Transfer (STTT), 19(2): 141–166, April 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s10009-015-0400-3;   http://link.
                                        springer.com/article/10.1007/s10009-
                                        015-0400-3.

Mammar:2017:MLG

[641] Amel Mammar and Régine Laleau. Modeling a landing gear system in Event-B. International Journal on Software Tools for Technology Transfer (STTT), 19(2):167–186, April 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s10009-015-0391-0;   http://link.
                                        springer.com/article/10.1007/s10009-
                                        015-0391-0.

Ladenberger:2017:VAL

[642] Lukas Ladenberger, Dominik Hansen, Harald Wiegard, Jens Bendisposto, and Michael Leuschel. Validation of the ABZ landing gear system using ProB. International Journal on Software Tools for Technology Transfer (STTT), 19(2): 187–203, April 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10009-015-0395-9;   http://link.
springer.com/article/10.1007/s10009-
015-0395-9.

Banach:2017:LGS

[643] Richard Banach. The landing gear system in multi-machine Hybrid Event-B. International Journal on Software Tools for Technology Transfer (STTT), 19(2):205–228, April 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//link.springer.com/article/10.
1007/s10009-015-0409-7;   http://
link.springer.com/content/pdf/10.
1007/s10009-015-0409-7.pdf.

Teodorov:2017:EDR

[644] Ciprian Teodorov, Philippe Dhaussy, and Luka Le Roux. Environment-driven reachability for timed systems. International Journal on Software Tools for Technology Transfer (STTT), 19(2): 229–245, April 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10009-015-0401-2;   http://link.
springer.com/article/10.1007/s10009-
015-0401-2.

Arcaini:2017:RDP

[645] Paolo Arcaini, Angelo Gargantini, and Elvinia Riccobene. Rigorous development process of a safety-critical system: from ASM models to Java code. International Journal on Software Tools for Technology Transfer (STTT), 19(2): 247–269, April 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s10009-015-0394-x;   http://link.
                                        springer.com/article/10.1007/s10009-
                                        015-0394-x.

Herbold:2017:MBT

[646] Steffen Herbold and Andreas Hoffmann. Model-based testing as a service. International Journal on Software Tools for Technology Transfer (STTT), 19(3):271–279, June 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL https://link.

                                        springer.com/content/pdf/10.1007/
                                        s10009-017-0449-2.pdf.

Hillah:2017:AIS

[647] Lom Messan Hillah, Ariele-Paolo Maesano, Fabio De Rosa, Fabrice Kordon, Pierre-Henri Wuillemin, Riccardo Fontanelli, Sergio Di Bona, Davide Guerri, and Libero Maesano. Automation and intelligent scheduling of distributed system functional testing. International Journal on Software Tools for Technology Transfer (STTT), 19(3): 281–308, June 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Herbold:2017:CUB

[648] Steffen Herbold, Patrick Harms, and Jens Grabowski. Combining usagebased and model-based testing for service-oriented architectures in the industrial practice. International Journal on Software Tools for Technology Transfer (STTT), 19(3):309–324, June 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Barcelona:2017:PEU

[649] M. A. Barcelona, L. García-Borgoñón, and G. López-Nicolás. Practical experiences in the usage of MIDAS in the logistics domain. International Journal on Software Tools for Technology Transfer (STTT), 19(3):325–339, June 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Falcone:2017:FAR

[650] Yliès Falcone and Mohamad Jaber.

Fully automated runtime enforcement of component-based systems with formal and sound recovery. International Journal on Software Tools for Technology Transfer (STTT), 19(3):341–365, June 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Jacobs:2017:FRS

[651] Swen Jacobs, Roderick Bloem, Romain Brenguier, Rüdiger Ehlers, Timotheus Hell, Robert Könighofer, Guillermo A. Pérez, Jean-François Raskin, Leonid Ryzhyk, Ocan Sankur, Martina Seidl, Leander Tentrup, and Adam Walker. The first reactive synthesis competition (SYNTCOMP 2014). International Journal on Software Tools for Technology Transfer (STTT), 19(3):367–390, June 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Gudemann:2017:PSI

[652] Matthias Güdemann and Manuel Núñez. Preface of the special issue on formal methods in industrial critical systems. International Journal on Software Tools for Technology Transfer (STTT), 19(4):391–393, August 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-017-0455-
                                        4.

Aman:2017:VCS

[653] Bogdan Aman and Gabriel Ciobanu. Verification of critical systems described in real-time TiMo. International Journal on Software Tools for Technology Transfer (STTT), 19(4):395–408, August 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-016-0439-
                                        9.

Aichernig:2017:RTT

[654] Bernhard K. Aichernig, Klaus Hörmaier, Florian Lorber, Dejan Nicković, and Stefan Tiran. Require, test, and trace IT. International Journal on Software Tools for Technology Transfer (STTT), 19(4):409–426, August 2017. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10009-016-0444-z;   https:
                                        //link.springer.com/content/pdf/
                                        10.1007/s10009-016-0444-z.pdf.

Damouche:2017:INA

[655] Nasrine Damouche, Matthieu Martel, and Alexandre Chapoutot. Improving the numerical accuracy of programs by automatic transformation. International Journal on Software Tools for Technology Transfer (STTT), 19(4):427–448, August 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-016-0435-
0.

Katz:2017:SCI

[656] Gal Katz and Doron Peled. Synthesizing, correcting and improving code, using model checking-based genetic programming. International Journal on Software Tools for Technology Transfer (STTT), 19(4):449–464, August 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-016-0418-
1.

Farah:2017:CCM

[657] Zoubeyr Farah, Yamine Ait-Ameur, Meriem Ouederni, and Kamel Tari. A correct-by-construction model for asynchronously communicating systems. International Journal on Software Tools for Technology Transfer (STTT), 19(4):465–485, August 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-016-0421-
6.

Hendriks:2017:AET

[658] Martijn Hendriks, Jacques Verriet, Twan Basten, Bart Theelen, Marco Brassé, and Lou Somers. Analyzing execution traces: critical-path analysis and distance analysis. International Journal on Software Tools for Technology Transfer (STTT), 19(4):487–510, August 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-016-0436-
                                        z.

Piterman:2017:AVP

[659] Nir Piterman. Advances in verification presented in TACAS’13. International Journal on Software Tools for Technology Transfer (STTT), 19 (5):511–515, October 2017. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10009-017-0460-7;   https:
                                        //link.springer.com/content/pdf/
                                        10.1007/s10009-017-0460-7.pdf.

Fedyukovich:2017:FSB

[660] Grigory Fedyukovich, Ondrej Sery, and Natasha Sharygina. Flexible SAT-based framework for incremental bounded upgrade checking. International Journal on Software Tools for Technology Transfer (STTT), 19(5):517–534, October 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-015-0405-
                                        y.

Dillig:2017:SCC

[661] Isil Dillig, Thomas Dillig, Boyang Li, Ken McMillan, and Mooly Sagiv. Synthesis of circular compositional program proofs via abduction. International Journal on Software Tools for Technology Transfer (STTT), 19(5):535–547, October 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-015-0397-
                                        7.

Abdulla:2017:ISV

[662] Parosh Aziz Abdulla, Frédéric Haziza, Lukás Holík, Bengt Jonsson, and Ahmed Rezine. An integrated specification and verification technique for highly concurrent data structures for highly concurrent data structures. International Journal on Software Tools for Technology Transfer (STTT), 19(5):549–563, October 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-016-0415-

4. See correction [836].

Ganty:2017:UPS

[663] Pierre Ganty, Radu Iosif, and Filip Konecný. Underapproximation of procedure summaries for integer programs. International Journal on Software Tools for Technology Transfer (STTT), 19(5):565–584, October 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-016-0420-
7.

Dimovski:2017:EFB

[664] Aleksandar S. Dimovski, Ahmad Salim Al-Sibahi, Claus Brabrand, and Andrzej Wasowski. Efficient familybased model checking via variability abstractions. International Journal on Software Tools for Technology Transfer (STTT), 19(5):585–603, October 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-016-0425-

Riener:2017:MFY

[665] Heinz Riener, Finn Haedicke, Stefan Frehse, Mathias Soeken, Daniel Große, Rolf Drechsler, and Goerschwin Fey. metaSMT: focus on your application and not on solver integration. International Journal on Software Tools for Technology Transfer (STTT), 19(5):605–621, October 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-016-0426-
                                        1.

Abe:2017:GMC

[666] Tatsuya Abe and Toshiyuki Maeda. A general model checking framework for various memory consistency models. International Journal on Software Tools for Technology Transfer (STTT), 19(5):623–647, October 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-016-0429-
                                        y.

Baier:2017:SAT

[667] Christel Baier and Cesare Tinelli. Some advances in tools and algorithms for the construction and analysis of systems. International Journal on Software Tools for Technology Transfer (STTT), 19(6):649–652, November 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-017-0471-
                                        4.

Renault:2017:VPE

[668] E. Renault, A. Duret-Lutz, F. Kordon, and D. Poitrenaud. Variations on parallel explicit emptiness checks for generalized Büchi automata. International Journal on Software Tools for Technology Transfer (STTT), 19(6):653–673, November 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-016-0422-
5.

vanDijk:2017:SMC

[669] Tom van Dijk and Jaco van de Pol. Sylvan: multi-core framework for decision diagrams. International Journal on Software Tools for Technology Transfer (STTT), 19 (6):675–696, November 2017. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL https:

//link.springer.com/article/10.
1007/s10009-016-0433-2;   https:
//link.springer.com/content/pdf/
10.1007/s10009-016-0433-2.pdf.

Furia:2017:AAA

[670] Carlo A. Furia, Martin Nordio, Nadia Polikarpova, and Julian Tschannen. AutoProof: auto-active functional verification of object-oriented programs. International Journal on Software Tools for Technology Transfer (STTT), 19(6):697–716, November 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-016-0419-
0.

Jeannin:2017:FVH

[671] Jean-Baptiste Jeannin, Khalil Ghorbal, Yanni Kouskoulas, Aurora Schmidt, Ryan Gardner, Stefan Mitsch, and André Platzer. A formally verified hybrid system for safe advisories in the next-generation airborne collision avoidance system. International Journal on Software Tools for Technology Transfer (STTT), 19(6):717–741, November 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-016-0434-
1.

Parizi:2017:RAA

[672] Reza Meimandi Parizi, Abdul Azim Abdul Ghani, Sai Peck Lee, and Saif Ur Rehman Khan. RAMBUTANS: automatic AOP-specific test generation tool. International Journal on Software Tools for Technology Transfer (STTT), 19(6):743–761, November 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-016-0432-
3.

Huisman:2017:V

[673] Marieke Huisman, Vladimir Klebanov, Rosemary Monahan, and Michael Tautschnig. VerifyThis 2015. International Journal on Software Tools for Technology Transfer (STTT), 19 (6):763–771, November 2017. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL https:

//link.springer.com/article/10.
1007/s10009-016-0438-x;   https:
//link.springer.com/content/pdf/
10.1007/s10009-016-0438-x.pdf.

Jaber:2018:HLM

[674] Mohamad Jaber, Yliès Falcone, Kinan Dak-Al-Bab, John Abou-Jaoudeh, and Mostafa El-Katerji. A high-level modeling language for the efficient design, implementation, and testing of Android applications. International Journal on Software Tools for Technology Transfer (STTT), 20(1):1–18, February 2018. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-016-0441-
2.

Boukhari:2018:RUR

[675] Ilyès Boukhari, Stéphane Jean, Idir Ait-Sadoune, and Ladjel Bellatreche. The role of user requirements in data repository design. International Journal on Software Tools for Technology Transfer (STTT), 20(1):19–34, February 2018. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-016-0443-
0.

Ulyantsev:2018:EFS

[676] Vladimir Ulyantsev, Igor Buzhinsky, and Anatoly Shalyto. Exact finitestate machine identification from scenarios and temporal properties. International Journal on Software Tools for Technology Transfer (STTT), 20(1): 35–55, February 2018. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10009-016-
0442-1.

Holmes:2018:TTS

[677] Josie Holmes, Alex Groce, Jervis Pinto, Pranjal Mittal, Pooria Azimi, Kevin Kellar, and James O’Brien. TSTL: the template scripting testing language. International Journal on Software Tools for Technology Transfer (STTT), 20(1):57–78, February 2018. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-016-0445-
                                        y.

Hoxha:2018:MPT

[678] Bardh Hoxha, Adel Dokhanchi, and Georgios Fainekos. Mining parametric temporal logic properties in model-based design for cyber-physical systems. International Journal on Software Tools for Technology Transfer (STTT), 20(1):79–93, February 2018. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-017-0447-
                                        4.

Pantelic:2018:SEP

[679] Vera Pantelic, Steven Postma, Mark Lawford, Monika Jaskolka, Bennett Mackenzie, Alexandre Korobkine, Marc Bender, Jeff Ong, Gordon Marks, and Alan Wassyng. Software engineering practices and Simulink: bridging the gap. International Journal on Software Tools for Technology Transfer (STTT), 20(1):95–117, February 2018. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-017-0450-
                                        9.

Mateescu:2018:RAI

[680] Radu Mateescu. Recent advances in interactive and automated analysis. International Journal on Software Tools for Technology Transfer (STTT), 20(2): 119–123, April 2018. CODEN ????

ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10009-017-
0477-y.

Faithfull:2018:C

[681] Alexander Faithfull, Jesper Bengtson, Enrico Tassi, and Carst Tankink. Coqoon. International Journal on Software Tools for Technology Transfer (STTT), 20(2):125–137, April 2018. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-017-0457-
2.

Grov:2018:TTG

[682] Gudmund Grov and Yuhui Lin. The Tinker tool for graphical tactic development. International Journal on Software Tools for Technology Transfer (STTT), 20(2):139–155, April 2018. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL https:

//link.springer.com/article/10.
1007/s10009-017-0452-7;   https:
//link.springer.com/content/pdf/
10.1007/s10009-017-0452-7.pdf.

vanDijk:2018:MCS

[683] Tom van Dijk and Jaco van de Pol.

Multi-core symbolic bisimulation minimisation. International Journal on Software Tools for Technology Transfer (STTT), 20(2):157–177, April 2018. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL https:

//link.springer.com/article/10.
1007/s10009-017-0468-z;   https:
//link.springer.com/content/pdf/
10.1007/s10009-017-0468-z.pdf.

Klein:2018:APM

[684] Joachim Klein, Christel Baier, Philipp Chrszon, Marcus Daum, Clemens Dubslaff, Sascha Klüppelholz, Steffen Märcker, and David Müller. Advances in probabilistic model checking with PRISM: variable reordering, quantiles and weak deterministic Büchi automata. International Journal on Software Tools for Technology Transfer (STTT), 20(2):179–194, April 2018. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-017-0456-
3.

Kwiatkowska:2018:PGV

[685] Marta Kwiatkowska, David Parker, and Clemens Wiltsche. PRISM-games: verification and strategy synthesis for stochastic multi-player games with multiple objectives. International Journal on Software Tools for Technology Transfer (STTT), 20(2):195–210, April 2018. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL https:

//link.springer.com/article/10.
1007/s10009-017-0476-z;   https:
//link.springer.com/content/pdf/
10.1007/s10009-017-0476-z.pdf.

Tran-Jorgensen:2018:ATV

[686] Peter W. V. Tran-Jørgensen, Peter Gorm Larsen, and Gary T. Leavens. Automated translation of VDM to JML-annotated Java. International Journal on Software Tools for Technology Transfer (STTT), 20(2):211–235, April 2018. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/

article/10.1007/s10009-017-0448-
3.

terBeek:2018:FMT

[687] Maurice H. ter Beek, Stefania Gnesi, and Alexander Knapp. Formal methods for transport systems. International Journal on Software Tools for Technology Transfer (STTT), 20(3):237–241, June 2018. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-018-0487-
4.

Vanit-Anunchai:2018:MST

[688] Somsak Vanit-Anunchai. Modelling and simulating a Thai railway signalling system using Coloured Petri Nets. International Journal on Software Tools for Technology Transfer (STTT), 20 (3):243–262, June 2018. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10009-018-
0482-9.

Mazzanti:2018:TFM

[689] Franco Mazzanti, Alessio Ferrari, and Giorgio O. Spagnolo. Towards formal methods diversity in railways: an experience report with seven frameworks. International Journal on Software Tools for Technology Transfer (STTT), 20 (3):263–288, June 2018. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10009-018-
0488-3.

Ciancia:2018:STM

[690] Vincenzo Ciancia, Stephen Gilmore, Gianluca Grilletti, Diego Latella, Michele Loreti, and Mieke Massink.

Spatio-temporal model checking of vehicular movement in public transport systems. International Journal on Software Tools for Technology Transfer (STTT), 20(3):289–311, June 2018. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-018-0483-
8.

Cabodi:2018:SGD

[691] G. Cabodi, P. E. Camurati, C. Loiacono, M. Palena, P. Pasini, D. Patti, and S. Quer. To split or to group: from divide-and-conquer to sub-task sharing for verifying multiple properties in model checking. International Journal on Software Tools for Technology Transfer (STTT), 20(3):313–325, June 2018. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-017-0451-
8.

Naujokat:2018:CSD

[692] Stefan Naujokat, Michael Lybecait, Dawid Kopetzki, and Bernhard Steffen. CINCO: a simplicity-driven approach to full generation of domain-specific graphical modeling tools. International Journal on Software Tools for Technology Transfer (STTT), 20(3):327–354, June 2018. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-017-0453-
6.

terBeek:2018:FMA

[693] Maurice H. ter Beek, Stefania Gnesi, and Alexander Knapp. Formal methods and automated verification of critical systems. International Journal on Software Tools for Technology Transfer (STTT), 20(4):355–358, August 2018. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-018-0494-
5.

Leupolz:2018:QQA

[694] Johannes Leupolz, Alexander Knapp, Axel Habermaier, and Wolfgang Reif. Qualitative and quantitative analysis of safety-critical systems with. International Journal on Software Tools for Technology Transfer (STTT), 20(4): 359–377, August 2018. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10009-017-
0464-3.

Pedro:2018:RVA

[695] André de Matos Pedro, Jorge Sousa Pinto, David Pereira, and Luís Miguel Pinho. Runtime verification of autopilot systems using a fragment of MTL-∫. International Journal on Software Tools for Technology Transfer (STTT), 20(4):379–395, August 2018. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-017-0470-
5.

Chadli:2018:HLF

[696] Mounir Chadli, Jin H. Kim, Kim G.

Larsen, Axel Legay, Stefan Naujokat, Bernhard Steffen, and Louis-Marie Traonouez. High-level frameworks for the specification and verification of scheduling problems. International Journal on Software Tools for Technology Transfer (STTT), 20(4):397–422, August 2018. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-017-0466-
                                        1.

Ge:2018:IFV

[697] Ning Ge, Eric Jenn, Nicolas Breton, and Yoann Fonteneau. Integrated formal verification of safetycritical software. International Journal on Software Tools for Technology Transfer (STTT), 20(4):423–440, August 2018. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-017-0475-
                                        0.

Huang:2018:MBT

[698] Wen ling Huang and Jan Peleska. Model-based testing strategies and their (in)dependence on syntactic model representations. International Journal on Software Tools for Technology Transfer (STTT), 20(4):441–465, August 2018. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-017-0479-
                                        9.

Bride:2018:ASC

[699] Hadrien Bride, Olga Kouchnarenko, Fabien Peureux, and Guillaume Voiron. Assessing SMT and CLP approaches for workflow nets verification. International Journal on Software Tools for Technology Transfer (STTT), 20(4):467–491, August 2018. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/

article/10.1007/s10009-018-0486-
5.

Bosnacki:2018:MCR

[700] Dragan Bosnacki and Anton Wijs.

Model checking: recent improvements and applications. International Journal on Software Tools for Technology Transfer (STTT), 20(5): 493–497, October 2018. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL https:

//link.springer.com/article/10.
1007/s10009-018-0501-x;   https:
//link.springer.com/content/pdf/
10.1007/s10009-018-0501-x.pdf.

Gallardo:2018:IRB

[701] María del Mar Gallardo, Pedro Merino, Laura Panizo, and Alberto Salmerón. Integrating river basin DSSs with model checking. International Journal on Software Tools for Technology Transfer (STTT), 20(5):499–514, October 2018. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-017-0478-
x.

Edelkamp:2018:CSP

[702] Stefan Edelkamp and Christoph Greulich.

A case study of planning for smart factories. International Journal on Software Tools for Technology Transfer (STTT), 20(5):515–528, October 2018. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL https:

//link.springer.com/article/10.
1007/s10009-018-0498-1;   https:
//link.springer.com/content/pdf/
10.1007/s10009-018-0498-1.pdf.

Jensen:2018:DCS

[703] Peter Gjøl Jensen, Kim Guldstrand Larsen, and Jirí Srba. Discrete and continuous strategies for timed-arc Petri net games. International Journal on Software Tools for Technology Transfer (STTT), 20(5):529–546, October 2018. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-017-0473-
                                        2.

Khamespanah:2018:MAR

[704] Ehsan Khamespanah, Marjan Sirjani, Kirill Mechitov, and Gul Agha. Modeling and analyzing real-time wireless sensor and actuator networks using actors and model checking. International Journal on Software Tools for Technology Transfer (STTT), 20(5):547–561, October 2018. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-017-0480-
                                        3.

Mateescu:2018:FMC

[705] Radu Mateescu and José Ignacio Requeno. On-the-fly model checking for extended action-based probabilistic operators. International Journal on Software Tools for Technology Transfer (STTT), 20(5):563–587, October 2018. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-018-0499-
                                        0.

Valmari:2018:FTS

[706] Antti Valmari and Walter Vogler. Fair testing and stubborn sets. International Journal on Software Tools for Technology Transfer (STTT), 20(5):589–610, October 2018. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-017-0481-
2.

Huisman:2018:SQT

[707] Marieke Huisman and Julia Rubin. Software quality tools and techniques presented in FASE’17. International Journal on Software Tools for Technology Transfer (STTT), 20 (6):611–613, November 2018. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL https:

//link.springer.com/article/10.
1007/s10009-018-0504-7;   https:
//link.springer.com/content/pdf/
10.1007/s10009-018-0504-7.pdf.

Muller:2018:TCC

[708] Andreas Müller, Stefan Mitsch, Werner Retschitzegger, Wieland Schwinger, and André Platzer. Tactical contract composition for hybrid system component verification. International Journal on Software Tools for Technology Transfer (STTT), 20 (6):615–643, November 2018. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL https:

//link.springer.com/article/10.
1007/s10009-018-0502-9;   https:
//link.springer.com/content/pdf/
10.1007/s10009-018-0502-9.pdf.

Cheng:2018:SAM

[709] Zheng Cheng and Massimo Tisi. Slicing ATL model transformations for scalable deductive verification and fault localization. International Journal on Software Tools for Technology Transfer (STTT), 20(6):645–663, November 2018. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-018-0491-
8.

Uva:2018:AWJ

[710] Marcelo Uva, Pablo Ponzio, Germán Regis, Nazareno Aguirre, and Marcelo F. Frias. Automated workarounds from Java program specifications based on SAT solving. International Journal on Software Tools for Technology Transfer (STTT), 20(6):665–688, November 2018. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-018-0503-
8.

Wang:2018:LPM

[711] Jingyi Wang, Jun Sun, Qixia Yuan, and Jun Pang. Learning probabilistic models for model checking: an evolutionary approach and an empirical study. International Journal on Software Tools for Technology Transfer (STTT), 20(6):689–704, November 2018. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-018-0492-
7.

Schneider:2018:ARA

[712] Sven Schneider, Leen Lambers, and Fernando Orejas. Automated reasoning for attributed graph properties. International Journal on Software Tools for Technology Transfer (STTT), 20(6):705–737, November 2018. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-018-0496-
3.

Asavoae:2018:SFS

[713] Irina Mariuca Asavoae, Mihail Asavoae, and Adrián Riesco. Slicing from formal semantics: Chisel — a tool for generic program slicing. International Journal on Software Tools for Technology Transfer (STTT), 20(6):739–769, November 2018. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-018-0500-
y.

Beyer:2019:RBR

[714] Dirk Beyer, Stefan Löwe, and Philipp Wendler. Reliable benchmarking: requirements and solutions. International Journal on Software Tools for Technology Transfer (STTT), 21(1): 1–29, February 2019. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10009-017-

0469-y. See correction [885].

Bartocci:2019:FIC

[715] Ezio Bartocci, Yliès Falcone, Borzoo Bonakdarpour, Christian Colombo, Normann Decker, Klaus Havelund, Yogi Joshi, Felix Klaedtke, Reed Milewicz, Giles Reger, Grigore Rosu, Julien Signoles, Daniel Thoma, Eugen Zalinescu, and Yi Zhang. First international Competition on Runtime Verification: rules, benchmarks, tools, and final results of CRV 2014. International Journal on Software Tools for Technology Transfer (STTT), 21(1):31–70, February 2019.

CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10009-017-0454-5;   https:
                                        //link.springer.com/content/pdf/
                                        10.1007/s10009-017-0454-5.pdf.

Fellner:2019:GPP

[716] Andreas Fellner and Bruno Woltzenlogel Paleo. Greedy pebbling for proof space compression. International Journal on Software Tools for Technology Transfer (STTT), 21(1):71–86, February 2019. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10009-017-0459-0;   https:
                                        //link.springer.com/content/pdf/
                                        10.1007/s10009-017-0459-0.pdf.

Bak:2019:HAV

[717] Stanley Bak, Omar Ali Beg, Sergiy Bogomolov, Taylor T. Johnson, Luan Viet Nguyen, and Christian Schilling. Hybrid automata: from verification to implementation. International Journal on Software Tools for Technology Transfer (STTT), 21(1):87–104, February 2019. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-017-0458-
                                        1.

Dashevskyi:2019:TFR

[718] Stanislav Dashevskyi, Daniel Ricardo dos Santos, Fabio Massacci, and Antonino Sabetta. TestREx: a framework for repeatable exploits. International Journal on Software Tools for Technology Transfer (STTT), 21(1): 105–119, February 2019. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10009-017-
0474-1.

Routhier:2019:QAR

[719] Maxime Routhier and Richard St-Denis.

A qualitative assessment of α Rby in the perspective of the supervisory control theory. International Journal on Software Tools for Technology Transfer (STTT), 21(2):121–141, April 2019. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-017-0461-
6.

Zaraket:2019:HLM

[720] Fadi A. Zaraket, Mohamad Jaber, Mohamad Noureddine, and Yliès Falcone. From high-level modeling toward efficient and trustworthy circuits. International Journal on Software Tools for Technology Transfer (STTT), 21(2): 143–163, April 2019. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10009-017-
0462-5.

Fu:2019:FMA

[721] Chunyan Fu and Kougen Zheng. Formal modeling and analysis of ad hoc Zone Routing Protocol in Event-B. International Journal on Software Tools for Technology Transfer (STTT), 21(2): 165–181, April 2019. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10009-017-
0463-4.

Perez:2019:FSI

[722] Gervasio Pérez and Sergio Yovine. Formal specification and implementation of an automated pattern-based parallelcode generation framework. International Journal on Software Tools for Technology Transfer (STTT), 21(2): 183–202, April 2019. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s10009-017-
                                        0465-2.

Andre:2019:WDA

[723] Étienne André. What’s decidable about parametric timed automata? International Journal on Software Tools for Technology Transfer (STTT), 21(2): 203–219, April 2019. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s10009-017-
                                        0467-0.

Zech:2019:KBS

[724] Philipp Zech, Michael Felderer, and Ruth Breu. Knowledge-based security testing of web applications by logic programming. International Journal on Software Tools for Technology Transfer (STTT), 21(2):221–246, April 2019. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10009-017-0472-3;   https:
                                        //link.springer.com/content/pdf/
                                        10.1007/s10009-017-0472-3.pdf.

Erdogmus:2019:ISP

[725] Hakan Erdogmus and Klaus Havelund. Introduction to Selected Papers from SPIN 2017. International Journal on Software Tools for Technology Transfer (STTT), 21(3):247–248, June 2019. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL https:

//link.springer.com/article/10.
1007/s10009-019-00515-5;   https:
//link.springer.com/content/pdf/
10.1007/s10009-019-00515-5.pdf.

Hua:2019:EED

[726] Jinru Hua, Yushan Zhang, Yuqun Zhang, and Sarfraz Khurshid. EdSketch: execution-driven sketching for Java. International Journal on Software Tools for Technology Transfer (STTT), 21(3):249–265, June 2019. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10009-019-
00512-8.

Ratiu:2019:IES

[727] Daniel Ratiu and Andreas Ulrich. An integrated environment for spin-based c code checking. International Journal on Software Tools for Technology Transfer (STTT), 21(3):267–286, June 2019. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-019-00510-
w.

Kokologiannakis:2019:SMC

[728] Michalis Kokologiannakis and Konstantinos Sagonas. Stateless model checking of the Linux kernel’s read-copy update (RCU). International Journal on Software Tools for Technology Transfer (STTT), 21(3):287–306, June 2019. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL https:

//link.springer.com/article/10.
1007/s10009-019-00514-6;   https:

//link.springer.com/content/pdf/
10.1007/s10009-019-00514-6.pdf.

Bloemen:2019:MCG

[729] Vincent Bloemen, Alexandre Duret-Lutz, and Jaco van de Pol. Model checking with generalized Rabin and Finless automata. International Journal on Software Tools for Technology Transfer (STTT), 21(3):307–324, June 2019. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL https:

//link.springer.com/article/10.
1007/s10009-019-00508-4;   https:
//link.springer.com/content/pdf/
10.1007/s10009-019-00508-4.pdf.

Fearnley:2019:OAS

[730] John Fearnley, Sanjay Jain, Bart de Keijzer, Sven Schewe, Frank Stephan, and Dominik Wojtczak. An ordered approach to solving parity games in quasi-polynomial time and quasilinear space. International Journal on Software Tools for Technology Transfer (STTT), 21(3):325–349, June 2019. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-019-00509-
3.

Nikravan:2019:RBA

[731] Esmaeel Nikravan and Saeed Parsa. A reasoning-based approach to dynamic domain reduction in test data generation. International Journal on Software Tools for Technology Transfer (STTT), 21(3):351–364, June 2019. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-018-0493-
6.

Parizek:2019:FDC

[732] Pavel Parízek and Ondrej Lhoták. Fast detection of concurrency errors by state space traversal with randomization and early backtracking. International Journal on Software Tools for Technology Transfer (STTT), 21(4):365–400, August 2019. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-018-0484-
7.

Gurung:2019:PRA

[733] Amit Gurung, Rajarshi Ray, Ezio Bartocci, Sergiy Bogomolov, and Radu Grosu. Parallel reachability analysis of hybrid systems in XSpeed. International Journal on Software Tools for Technology Transfer (STTT), 21(4):401–423, August 2019. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-018-0485-
6.

Botella:2019:CTS

[734] Julien Botella, Jean-François Capuron, Frédéric Dadeau, Elizabeta Fourneret, Bruno Legeard, and Florence Schadle. Complementary test selection criteria for model-based testing of security components. International Journal on Software Tools for Technology Transfer (STTT), 21(4):425–448, August 2019. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-018-0489-
2.

Busard:2019:CAM

[735] Simon Busard, Charles Pecheur, Hongyang Qu, and Franco Raimondi. Comparing approaches for model-checking strategies under imperfect information and fairness constraints. International Journal on Software Tools for Technology Transfer (STTT), 21(4):449–469, August 2019. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-018-0505-
6.

Tennyson:2019:ASC

[736] Matthew F. Tennyson. ASAP: a Source Code Authorship Program. International Journal on Software Tools for Technology Transfer (STTT), 21(4): 471–484, August 2019. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10009-019-
00517-3.

Hentschel:2019:SED

[737] Martin Hentschel, Richard Bubel, and Reiner Hähnle. The Symbolic Execution Debugger (SED): a platform for interactive symbolic execution, debugging, verification and more. International Journal on Software Tools for Technology Transfer (STTT), 21(5):485–513, October 2019. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-018-0490-
9.

Becker:2019:SPE

[738] Martin Becker, Ravindra Metta, R. Venkatesh, and Samarjit Chakraborty. Scalable and precise estimation and debugging of the worst-case execution time for analysis-friendly processors: a comeback of model checking. International Journal on Software Tools for Technology Transfer (STTT), 21(5):515–543, October 2019. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-018-0497-
2.

Herdt:2019:CSB

[739] Vladimir Herdt, Hoang M. Le, Daniel Große, and Rolf Drechsler. Combining sequentialization-based verification of multi-threaded C programs with symbolic Partial Order Reduction. International Journal on Software Tools for Technology Transfer (STTT), 21(5):545–565, October 2019. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-019-00507-
5.

Gibson-Robinson:2019:SRC

[740] Thomas Gibson-Robinson and Gavin Lowe. Symmetry reduction in CSP model checking. International Journal on Software Tools for Technology Transfer (STTT), 21(5):567–605, October 2019. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10009-019-
00516-4;   https://link.springer.
com/content/pdf/10.1007/s10009-
019-00516-4.pdf.

terBeek:2019:QVM

[741] Maurice H. ter Beek and Axel Legay.

Quantitative variability modelling and analysis. International Journal on Software Tools for Technology Transfer (STTT), 21(6):607–612, December 2019. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-019-00535-
                                        1.

Luthmann:2019:SSP

[742] Lars Luthmann, Timo Gerecht, and Malte Lochau. Sampling strategies for product lines with unbounded parametric real-time constraints. International Journal on Software Tools for Technology Transfer (STTT), 21(6):613– 633, December 2019. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s10009-019-
                                        00532-4.

Cordy:2019:VAR

[743] Maxime Cordy and Axel Legay. Verification and abstraction of real-time variability-intensive systems. International Journal on Software Tools for Technology Transfer (STTT), 21(6): 635–649, December 2019. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s10009-019-
                                        00537-z.

Herrmann:2019:CIP

[744] Linda Herrmann, Martin Küttler, Tobias Stumpf, Christel Baier, Hermann Härtig, and Sascha Klüppelholz. Configuration of inter-process communication with probabilistic model checking. International Journal on Software Tools for Technology Transfer (STTT), 21(6):651–666, December 2019. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/

article/10.1007/s10009-019-00536-
0.

Fahrenberg:2019:QPF

[745] Uli Fahrenberg and Axel Legay. Quantitative properties of featured automata. International Journal on Software Tools for Technology Transfer (STTT), 21(6):667–677, December 2019. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-019-00538-
y.

Basile:2019:ASC

[746] Davide Basile. Applying supervisory control synthesis to priced featured automata and energy problems. International Journal on Software Tools for Technology Transfer (STTT), 21(6):679–689, December 2019. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-019-00533-
3.

Damiani:2019:ARD

[747] Ferruccio Damiani, Michael Lienhardt, and Luca Paolini. Automatic refactoring of delta-oriented SPLs to remove-free form and replacefree form. International Journal on Software Tools for Technology Transfer (STTT), 21(6):691–707, December 2019. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-019-00534-

Russo:2020:MBS

[748] Alessandra Russo and Andy Schürr. Model-based software quality assurance tools and techniques presented at FASE 2018. International Journal on Software Tools for Technology Transfer (STTT), 22(1):1–2, February 2020. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-019-00541-
3.

Gioulekas:2020:CCM

[749] Fotios Gioulekas, Peter Poplavko, Panagiotis Katsaros, Saddek Bensalem, and Pedro Palomo. Correct-by-construction model-based design of reactive streaming software for multi-core embedded systems. International Journal on Software Tools for Technology Transfer (STTT), 22(1):3–32, February 2020. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-019-00521-

7. See correction [750].

Gioulekas:2020:CCC

[750] Fotios Gioulekas, Peter Poplavko, Panagiotis Katsaros, Saddek Bensalem, and Pedro Palomo. Correction to: Correctby-construction model-based design of reactive streaming software for multicore embedded systems. International Journal on Software Tools for Technology Transfer (STTT), 22(1):33–34, February 2020. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10009-019-
00523-5;   https://link.springer.

com/content/pdf/10.1007/s10009-
019-00523-5.pdf. See [749].

Dimovski:2020:CFB

[751] Aleksandar S. Dimovski. CTL^{⋆} familybased model checking using variability abstractions and modal transition systems. International Journal on Software Tools for Technology Transfer (STTT), 22(1):35–55, February 2020. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-019-00528-
0.

Semerath:2020:DGM

[752] Oszkár Semeráth, Rebeka Farkas, Gábor Bergmann, and Dániel Varró. Diversity of graph models and graph generators in mutation testing. International Journal on Software Tools for Technology Transfer (STTT), 22 (1):57–78, February 2020. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL https:

//link.springer.com/article/10.
1007/s10009-019-00530-6;   https:
//link.springer.com/content/pdf/
10.1007/s10009-019-00530-6.pdf.

Bur:2020:DGQ

[753] Márton Búr, Gábor Szilágyi, András Vörös, and Dániel Varró. Distributed graph queries over models@run.time for runtime monitoring of cyber-physical systems. International Journal on Software Tools for Technology Transfer (STTT), 22(1):79–102, February 2020. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL https:

//link.springer.com/article/10.
1007/s10009-019-00531-5;   https:

                                        //link.springer.com/content/pdf/
                                        10.1007/s10009-019-00531-5.pdf.

Gallardo:2020:ISI

[754] María del Mar Gallardo and Pedro Merino. Introduction to the special issue devoted to SPIN 2018. International Journal on Software Tools for Technology Transfer (STTT), 22 (2):103–104, April 2020. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10009-020-00550-7;   https:
                                        //link.springer.com/content/pdf/
                                        10.1007/s10009-020-00550-7.pdf.

Panizo:2020:MBT

[755] Laura Panizo, Almudena Díaz, and Bruno García. Model-based testing of apps in real network scenarios. International Journal on Software Tools for Technology Transfer (STTT), 22(2): 105–114, April 2020. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s10009-019-
                                        00518-2.

Chalupa:2020:JFM

[756] Marek Chalupa, Jan Strejcek, and Martina Vitovská. Joint forces for memory safety checking revisited. International Journal on Software Tools for Technology Transfer (STTT), 22(2):115–133, April 2020. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-019-00526-
                                        2.

Lange:2020:ISM

[757] Tim Lange, Martin R. Neuhäußer, Thomas Noll, and Joost-Pieter Katoen. IC3 software model checking. International Journal on Software Tools for Technology Transfer (STTT), 22(2): 135–161, April 2020. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10009-019-
00547-x.

Berthomieu:2020:CPN

[758] Bernard Berthomieu, Didier Le Botlan, and Silvano Dal Zilio. Counting Petri net markings from reduction equations. International Journal on Software Tools for Technology Transfer (STTT), 22(2):163–181, April 2020. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-019-00519-
1.

Allawi:2020:GPS

[759] Hamzeh M. Allawi, Waref Al Manaseer, and Mohammad Al Shraideh. A greedy particle swarm optimization (GPSO) algorithm for testing real-world smart card applications. International Journal on Software Tools for Technology Transfer (STTT), 22(2):183–194, April 2020. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-018-00506-
y.

Buonamici:2020:SLM

[760] Fabrizio Banci Buonamici, Gina Belmonte, Vincenzo Ciancia, Diego Latella, and Mieke Massink. Spatial logics and model checking for medical imaging. International Journal on Software Tools for Technology Transfer (STTT), 22(2):195–217, April 2020. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-019-00511-
                                        9.

Mkaouar:2020:FAA

[761] Hana Mkaouar, Bechir Zalila, Jérôme Hugues, and Mohamed Jmaiel. A formal approach to AADL model-based software engineering. International Journal on Software Tools for Technology Transfer (STTT), 22(2):219–247, April 2020. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-019-00513-
                                        7.

Butler:2020:ISS

[762] Michael Butler, Thai Son Hoang, Alexander Raschke, and Klaus Reichl. Introduction to special section on the ABZ 2018 case study: Hybrid ERTMS/ ETCS Level 3. International Journal on Software Tools for Technology Transfer (STTT), 22(3):249–255, June 2020. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10009-020-00562-3;   https:
                                        //link.springer.com/content/pdf/
                                        10.1007/s10009-020-00562-3.pdf.

Abrial:2020:ACS

[763] Jean-Raymond Abrial. The ABZ-2018 case study with Event-B. International Journal on Software Tools for Technology Transfer (STTT), 22(3):257–264, June 2020. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-019-00525-
                                        3.

Arcaini:2020:VHE

[764] Paolo Arcaini, Jan Kofron, and Pavel Jezek. Validation of the Hybrid ERTMS/ETCS Level 3 using Spin. International Journal on Software Tools for Technology Transfer (STTT), 22 (3):265–279, June 2020. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10009-019-
00539-x.

Cunha:2020:VHE

[765] Alcino Cunha and Nuno Macedo. Validating the Hybrid ERTMS/ETCS Level 3 concept with Electrum. International Journal on Software Tools for Technology Transfer (STTT), 22(3):281–296, June 2020. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-019-00540-
4.

Dghaym:2020:FHE

[766] Dana Dghaym, Mohammadsadegh Dalvandi, Michael Poppleton, and Colin Snook. Formalising the Hybrid ERTMS Level 3 specification in iUML-B and Event-B. International Journal on Software Tools for Technology Transfer (STTT), 22(3):297–313, June 2020. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL https:

//link.springer.com/article/10.
1007/s10009-019-00548-w;   https:
//link.springer.com/content/pdf/
10.1007/s10009-019-00548-w.pdf.

Hansen:2020:VRL

[767] Dominik Hansen, Michael Leuschel, Philipp Körner, Sebastian Krings, Thomas Naulin, Nader Nayeri, David Schneider, and Frank Skowron. Validation and real-life demonstration of ETCS hybrid level 3 principles using a formal B model. International Journal on Software Tools for Technology Transfer (STTT), 22(3):315–332, June 2020. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL https:

//link.springer.com/article/10.
1007/s10009-020-00551-6;   https:
//link.springer.com/content/pdf/
10.1007/s10009-020-00551-6.pdf.

Mammar:2020:FRB

[768] Amel Mammar, Marc Frappier, Steve Jeffrey Tueno Fotso, and Régine Laleau. A formal refinement-based analysis of the hybrid ERTMS/ETCS Level 3 standard. International Journal on Software Tools for Technology Transfer (STTT), 22(3):333–347, June 2020. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-019-00543-
1.

Fotso:2020:MHE

[769] Steve Jeffrey Tueno Fotso, Marc Frappier, Régine Laleau, and Amel Mammar. Modeling the hybrid ERTMS/ ETCS Level 3 standard using a formal requirements engineering approach. International Journal on Software Tools for Technology Transfer (STTT), 22 (3):349–363, June 2020. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10009-019-
00542-2.

Parsai:2020:CMC

[770] Ali Parsai and Serge Demeyer. Comparing mutation coverage against branch coverage in an industrial setting. International Journal on Software Tools for Technology Transfer (STTT), 22 (4):365–388, August 2020. CODEN

????  ISSN ????  URL https:

//link.springer.com/article/10.
1007/s10009-020-00567-y.

DeNicola:2020:REC

[771] Rocco De Nicola, Stefan Jähnichen, and Martin Wirsing. Rigorous engineering of collective adaptive systems: special section. International Journal on Software Tools for Technology Transfer (STTT), 22(4):389–397, August 2020. CODEN ???? ISSN ????

URL https://link.springer.com/
article/10.1007/s10009-020-00565-
0.

Abeywickrama:2020:SAE

[772] Dhaminda B. Abeywickrama, Nicola Bicocchi, and Franco Zambonelli. The SOTA approach to engineering collective adaptive systems. International Journal on Software Tools for Technology Transfer (STTT), 22(4):399–415, August 2020. CODEN ???? ISSN ????

URL https://link.springer.com/
article/10.1007/s10009-020-00554-
3.

BenMahfoudh:2020:LBC

[773] Houssem Ben Mahfoudh, Giovanna Di Marzo Serugendo, and Nabil Abdennadher. Learning-based coordination model for spontaneous selfcomposition of reliable services in a distributed system. International Journal on Software Tools for Technology Transfer (STTT), 22(4):417–436, August 2020. CODEN ???? ISSN ????

URL https://link.springer.com/

                                        article/10.1007/s10009-020-00557-
                                        0.

DeNicola:2020:DFD

[774] Rocco De Nicola, Alessandro Maggi, and Joseph Sifakis. The DReAM framework for dynamic reconfigurable architecture modelling: theory and applications. International Journal on Software Tools for Technology Transfer (STTT), 22(4):437–455, August 2020. CODEN ???? ISSN ????

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-020-00555-
                                        2.

Gabor:2020:SCP

[775] Thomas Gabor, Andreas Sedlmeier, and Claudia Linnhoff-Popien. The scenario coevolution paradigm: adaptive quality assurance for adaptive systems. International Journal on Software Tools for Technology Transfer (STTT), 22(4):457–476, August 2020. CODEN ???? ISSN ????

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-020-00560-
                                        5.

Alrahman:2020:DAC

[776] Yehia Abd Alrahman and Giulio Garbi. A distributed API for coordinating AbC programs. International Journal on Software Tools for Technology Transfer (STTT), 22(4):477–496, August 2020. CODEN ???? ISSN ????

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-020-00553-
                                        4.

Bures:2020:LFD

[777] Tomas Bures, Ilias Gerostathopoulos, and Jan Kofron. A language and framework for dynamic component ensembles in smart systems. International Journal on Software Tools for Technology Transfer (STTT), 22(4):497–509, August 2020. CODEN ???? ISSN ????

URL https://link.springer.com/
article/10.1007/s10009-020-00558-
z.

AlAli:2020:TAC

[778] Rima Al Ali, Tomas Bures, and Jiri Vinarek. Toward autonomically composable and context-dependent access control specification through ensembles. International Journal on Software Tools for Technology Transfer (STTT), 22(4):511–522, August 2020. CODEN ???? ISSN ????

URL https://link.springer.com/
article/10.1007/s10009-020-00556-
1.

Ceska:2020:ARF

[779] Milan Ceska, Vojtech Havlena, and Tomás Vojnar. Approximate reduction of finite automata for high-speed network intrusion detection. International Journal on Software Tools for Technology Transfer (STTT), 22(5):523–539, October 2020. CODEN ???? ISSN ????

URL https://link.springer.com/
article/10.1007/s10009-019-00520-
8.

Amparore:2020:VOM

[780] Elvio G. Amparore, Susanna Donatelli, and Gianfranco Ciardo. Variable order metrics for decision diagrams in system verification. International Journal on Software Tools for Technology Transfer (STTT), 22(5):541–562, October 2020. CODEN ???? ISSN ????

URL https://link.springer.com/
article/10.1007/s10009-019-00522-
6.

Biondi:2020:ISI

[781] Fabrizio Biondi, Thomas Given-Wilson, and Axel Legay. Introduction to the special issue for SPIN 2019. International Journal on Software Tools for Technology Transfer (STTT), 22(5):563–564, October 2020. CODEN ???? ISSN ????

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-020-00580-
                                        1.

Metzler:2020:EST

[782] Patrick Metzler, Neeraj Suri, and Georg Weissenbacher. Extracting safe thread schedules from incomplete model checking results. International Journal on Software Tools for Technology Transfer (STTT), 22(5):565–581, October 2020. CODEN ???? ISSN ????

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-020-00575-
                                        y.

DeFrancisco:2020:SMC

[783] Richard DeFrancisco, Shenghsun Cho, and Scott A. Smolka. Swarm model checking on the GPU. International Journal on Software Tools for Technology Transfer (STTT), 22(5):583–599, October 2020. CODEN ???? ISSN ????

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-020-00576-
                                        x.

Usman:2020:SLL

[784] Muhammad Usman, Wenxi Wang, and Sarfraz Khurshid. A study of learning likely data structure properties using machine learning models. International Journal on Software Tools for Technology Transfer (STTT), 22(5):601–615, October 2020. CODEN ???? ISSN ????

URL https://link.springer.com/
article/10.1007/s10009-020-00577-
w.

Yousefi:2020:VFV

[785] Farnaz Yousefi, Ehsan Khamespanah, and Ali Movaghar. VeriVANca framework: verification of VANETs by property-based message passing of actors in Rebeca with inheritance. International Journal on Software Tools for Technology Transfer (STTT), 22 (5):617–633, October 2020. CODEN

????  ISSN ????  URL https:

//link.springer.com/article/10.
1007/s10009-020-00579-8.

Enevoldsen:2020:DGA

[786] Søren Enevoldsen, Kim G. Larsen, and Jirí Srba. Dependency graphs with applications to verification. International Journal on Software Tools for Technology Transfer (STTT), 22(5):635–654, October 2020. CODEN ???? ISSN ????

URL https://link.springer.com/
article/10.1007/s10009-020-00578-
9.

Li:2020:RFS

[787] Yi Li, Wenyuan Wu, and Yong Feng. On ranking functions for single-path linearconstraint loops. International Journal on Software Tools for Technology Transfer (STTT), 22(6):655–666, December 2020. CODEN ???? ISSN ????

URL https://link.springer.com/
article/10.1007/s10009-019-00549-
9.

Couto:2020:ECI

[788] Luis Diogo Couto, Peter W. V. Tran-Jørgensen, and Peter Gorm Larsen. Enabling continuous integration in a formal methods setting. International Journal on Software Tools for Technology Transfer (STTT), 22(6):667–683, December 2020. CODEN ???? ISSN ????

URL https://link.springer.com/
article/10.1007/s10009-019-00546-
y.

Beyer:2020:TCA

[789] Dirk Beyer and Marieke Huisman. Tools for the construction and analysis of systems. International Journal on Software Tools for Technology Transfer (STTT), 22(6):685–687, December 2020. CODEN ???? ISSN ????

URL https://link.springer.com/
article/10.1007/s10009-020-00581-
0.

Dragomir:2020:RCR

[790] Iulia Dragomir, Viorel Preoteasa, and Stavros Tripakis. The Refinement Calculus of Reactive Systems Toolset. International Journal on Software Tools for Technology Transfer (STTT), 22(6):689–708, December 2020. CODEN ???? ISSN ????

URL https://link.springer.com/
article/10.1007/s10009-020-00561-
4.

Summers:2020:ADV

[791] Alexander J. Summers and Peter Müller. Automating deductive verification for weak-memory programs (extended version). International Journal on Software Tools for Technology Transfer (STTT), 22(6):709–728, December 2020. CODEN ???? ISSN ????

URL https://link.springer.com/
article/10.1007/s10009-020-00559-
y.

Finkbeiner:2020:EMH

[792] Bernd Finkbeiner, Christopher Hahn, and Leander Tentrup. Efficient monitoring of hyperproperties using prefix trees. International Journal on Software Tools for Technology Transfer (STTT), 22(6):729–740, December 2020. CODEN ???? ISSN ????

URL https://link.springer.com/
article/10.1007/s10009-020-00552-
5.

Nickovic:2020:AQQ

[793] Dejan Nicković, Olivier Lebeltel, and Dogan Ulus. AMT 2.0: qualitative and quantitative trace analysis with extended signal temporal logic. International Journal on Software Tools for Technology Transfer (STTT), 22(6):741–758, December 2020. CODEN ???? ISSN ????

URL https://link.springer.com/
article/10.1007/s10009-020-00582-
z.

Budde:2020:ESM

[794] Carlos E. Budde, Pedro R. D’Argenio, and Sean Sedwards. An efficient statistical model checker for nondeterminism and rare events. International Journal on Software Tools for Technology Transfer (STTT), 22(6):759–780, December 2020. CODEN ???? ISSN ????

URL https://link.springer.com/
article/10.1007/s10009-020-00563-
2.

Giantamidis:2021:LMM

[795] Georgios Giantamidis, Stavros Tripakis, and Stylianos Basagiannis. Learning Moore machines from input– output traces. International Journal on Software Tools for Technology Transfer (STTT), 23(1):1–29, February 2021. CODEN ???? ISSN ????

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-019-00544-
                                        0.

Alvin:2021:SGU

[796] Chris Alvin, Brian Peterson, and Supratik Mukhopadhyay. Static generation of UML sequence diagrams. International Journal on Software Tools for Technology Transfer (STTT), 23(1):31– 53, February 2021. CODEN ???? ISSN

                                        ???? URL https://link.springer.
                                        com/article/10.1007/s10009-019-
                                        00545-z.

Yuan:2021:DML

[797] Yue Yuan, Yi Li, and Wenchang Shi. Detecting multiphase linear ranking functions for single-path linearconstraint loops. International Journal on Software Tools for Technology Transfer (STTT), 23(1):55–67, February 2021. CODEN ???? ISSN ????

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-019-00527-
                                        1.

Kammuller:2021:MCC

[798] Florian Kammüller, Axel Legay, and Stefano Schivo. Masterminding change by combining secure system design with security risk assessment. International Journal on Software Tools for Technology Transfer (STTT), 23(1):69–70, February 2021. CODEN ???? ISSN ????

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-020-00595-
                                        8.

Ferrara:2021:SAD

[799] Pietro Ferrara, Amit Kr Mandal, and Fausto Spoto. Static analysis for discovering IoT vulnerabilities. International Journal on Software Tools for Technology Transfer (STTT), 23(1):71–88, February 2021. CODEN ???? ISSN ????

URL https://link.springer.com/
article/10.1007/s10009-020-00592-
x.

Hansen:2021:APL

[800] René Rydhof Hansen, Kim Guldstrand Larsen, and Danny Bøgsted Poulsen. ADTLang: a programming language approach to attack defense trees. International Journal on Software Tools for Technology Transfer (STTT), 23(1):89–104, February 2021. CODEN ???? ISSN ????

URL https://link.springer.com/
article/10.1007/s10009-020-00593-
w.

Shakhov:2021:GBT

[801] Vladimir Shakhov and Insoo Koo.

Graph-based technique for survivability assessment and optimization of IoT applications. International Journal on Software Tools for Technology Transfer (STTT), 23(1):105–114, February 2021. CODEN ???? ISSN ????

URL https://link.springer.com/
article/10.1007/s10009-020-00594-
9.

Alhawi:2021:VRC

[802] Omar M. Alhawi, Herbert Rocha, and Eddie Batista. Verification and refutation of C programs based on k-induction and invariant inference. International Journal on Software Tools for Technology Transfer (STTT), 23(2):115–135, April 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/

                                        article/10.1007/s10009-020-00564-
                                        1.

Khoury:2021:ABM

[803] Raphaël Khoury, Sylvain Hallé, and Yannick Lebrun. Automata-based monitoring for LTL-FO^{+}. International Journal on Software Tools for Technology Transfer (STTT), 23(2):137–154, April 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-020-00566-
                                        z.

Leucker:2021:P

[804] Martin Leucker and Christian Colombo. Preface. International Journal on Software Tools for Technology Transfer (STTT), 23(2):155–156, April 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-021-00610-
                                        6.

Gorostiaga:2021:SRV

[805] Felipe Gorostiaga and César Sánchez. Stream runtime verification of realtime event streams with the Striver language. International Journal on Software Tools for Technology Transfer (STTT), 23(2):157–183, April 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-021-00605-
                                        3.

Schneider:2021:SOF

[806] Joshua Schneider, David Basin, and Dmitriy Traytel. Scalable online firstorder monitoring. International Journal on Software Tools for Technology Transfer (STTT), 23(2):185–208, April 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-021-00607-
1.

Reger:2021:PTS

[807] Giles Reger and David Rydeheard.

From parametric trace slicing to rule systems. International Journal on Software Tools for Technology Transfer (STTT), 23(2):209–228, April 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-021-00608-
0.

Bouyer:2021:DTA

[808] Patricia Bouyer, Léo Henry, and Nicolas Markey. Diagnosing timed automata using timed markings. International Journal on Software Tools for Technology Transfer (STTT), 23(2):229–253, April 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-021-00606-
2.

Falcone:2021:TCR

[809] Yliès Falcone, Srdan Krstić, and Dmitriy Traytel. A taxonomy for classifying runtime verification tools. International Journal on Software Tools for Technology Transfer (STTT), 23(2): 255–284, April 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10009-021-
00609-z.

Hahnle:2021:AMA

[810] Reiner Hähnle and Wil van der Aalst. Automated model analysis tools and techniques presented at FASE 2019. International Journal on Software Tools for Technology Transfer (STTT), 23(3):285–287, June 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-020-00589-
                                        6.

Boronat:2021:IER

[811] Artur Boronat. Incremental execution of rule-based model transformation. International Journal on Software Tools for Technology Transfer (STTT), 23 (3):289–311, June 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s10009-020-
                                        00583-y.

Beyer:2021:CVB

[812] Dirk Beyer and Marie-Christine Jakobs. Cooperative verifier-based testing with CoVeriTest. International Journal on Software Tools for Technology Transfer (STTT), 23(3):313–333, June 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-020-00587-
                                        8.

Fritsche:2021:AUI

[813] Lars Fritsche, Jens Kosiol, and Gabriele Taentzer. Avoiding unnecessary information loss: correct and efficient model synchronization based on triple graph grammars. International Journal on Software Tools for Technology Transfer (STTT), 23(3):335–368, June 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-020-00588-
7.

Schneider:2021:LBI

[814] Sven Schneider, Leen Lambers, and Fernando Orejas. A logic-based incremental approach to graph repair featuring delta preservation. International Journal on Software Tools for Technology Transfer (STTT), 23(3):369–410, June 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-020-00584-
x.

Schneider:2021:FTT

[815] Sven Schneider, Maria Maximova, and Holger Giese. Formal testing of timed graph transformation systems using metric temporal graph logic. International Journal on Software Tools for Technology Transfer (STTT), 23(3): 411–488, June 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10009-020-
00585-w.

Dubrulle:2021:PDF

[816] Paul Dubrulle, Nikolai Kosmatov, and Arnault Lapitre. PolyGraph: a data flow model with frequency arithmetic. International Journal on Software Tools for Technology Transfer (STTT), 23 (3):489–517, June 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10009-020-

00586-9. See correction [817].

Dubrulle:2021:CPD

[817] Paul Dubrulle, Nikolai Kosmatov, and Arnault Lapitre. Correction to: PolyGraph: a data flow model with frequency arithmetic. International Journal on Software Tools for Technology Transfer (STTT), 23(3):519, June 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-020-00590-

z. See [816].

Frohme:2021:CLM

[818] Markus Frohme and Bernhard Steffen. Compositional learning of mutually recursive procedural systems. International Journal on Software Tools for Technology Transfer (STTT), 23(4): 521–543, August 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s10009-021-
                                        00634-y.

Finkbeiner:2021:ISI

[819] Bernd Finkbeiner and Leonardo Mariani. Introduction to the special issue of the 19th International Conference on Runtime Verification. International Journal on Software Tools for Technology Transfer (STTT), 23(4):545–546, August 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-021-00621-
                                        3.

Havelund:2021:EFO

[820] Klaus Havelund and Doron Peled. An extension of first-order LTL with rules with application to runtime verification. International Journal on Software Tools for Technology Transfer (STTT), 23(4):547–563, August 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-021-00626-
y.

Nickovic:2021:SDT

[821] Dejan Nicković, Xin Qin, and Jyotirmoy Deshmukh. Specifying and detecting temporal patterns with shape expressions. International Journal on Software Tools for Technology Transfer (STTT), 23(4):565–577, August 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-021-00627-
x.

Kauffman:2021:WCW

[822] Sean Kauffman, Klaus Havelund, and Sebastian Fischmeister. What can we monitor over unreliable channels? International Journal on Software Tools for Technology Transfer (STTT), 23(4):579–600, August 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-021-00625-
z.

Aceto:2021:CCS

[823] Luca Aceto, Ian Cassar, and Anna Ingólfsdóttir. Comparing controlled system synthesis and suppression enforcement. International Journal on Software Tools for Technology Transfer (STTT), 23(4):601–614, August 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-021-00624-
                                        0.

Bortolussi:2021:NPM

[824] Luca Bortolussi, Francesca Cairoli, and Scott D. Stoller. Neural predictive monitoring and a comparison of frequentist and Bayesian approaches. International Journal on Software Tools for Technology Transfer (STTT), 23(4):615–640, August 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-021-00623-
                                        1.

Tizpaz-Niari:2021:QES

[825] Saeid Tizpaz-Niari, Pavol Cerný, and Ashutosh Trivedi. Quantitative estimation of side-channel leaks with neural networks. International Journal on Software Tools for Technology Transfer (STTT), 23(4):641–654, August 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-021-00622-
                                        2.

Kopetzki:2021:TLL

[826] Dawid Kopetzki, Michael Lybecait, and Bernhard Steffen. Towards language-to-language transformation. International Journal on Software Tools for Technology Transfer (STTT), 23(5):655–677, October 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-021-00630-
                                        2.

Bliudze:2021:MTR

[827] Simon Bliudze, Panagiotis Katsaros, and Martin Wirsing. On methods and tools for rigorous system design. International Journal on Software Tools for Technology Transfer (STTT), 23(5):679–684, October 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-021-00632-
0.

Grompanopoulos:2021:SVU

[828] Christos Grompanopoulos, Antonios Gouglidis, and Anastasia Mavridou. Specifying and verifying usage control models and policies in TLA^{+}. International Journal on Software Tools for Technology Transfer (STTT), 23(5): 685–700, October 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10009-020-
00600-0.

ElBallouli:2021:PDR

[829] Rim El Ballouli, Saddek Bensalem, and Joseph Sifakis. Programming dynamic reconfigurable systems. International Journal on Software Tools for Technology Transfer (STTT), 23(5):701–719, October 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-020-00596-
7.

Han:2021:MBO

[830] Pujie Han, Zhengjun Zhai, and Ulrik Nyman. Model-based optimization of ARINC-653 partition scheduling. International Journal on Software Tools for Technology Transfer (STTT), 23(5):721–740, October 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-020-00597-
6.

Blom:2021:CPP

[831] S. Blom, S. Darabi, and M. Safari. Correct program parallelisations. International Journal on Software Tools for Technology Transfer (STTT), 23(5):741–763, October 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-020-00601-
z.

Lekidis:2021:ECI

[832] Alexios Lekidis and Panagiotis Katsaros. Energy characterization of IoT systems through design aspect monitoring. International Journal on Software Tools for Technology Transfer (STTT), 23(5):765–781, October 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-020-00598-
5.

Bartocci:2021:CAF

[833] Ezio Bartocci, Niveditha Manjunath, and Dejan Nicković. CPSDebug: Automatic failure explanation in CPS models. International Journal on Software Tools for Technology Transfer (STTT), 23(5):783–796, October 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/

article/10.1007/s10009-020-00599-
4.

Bertrand:2021:VRC

[834] Nathalie Bertrand, Igor Konnov, and Josef Widder. Verification of randomized consensus algorithms under roundrigid adversaries. International Journal on Software Tools for Technology Transfer (STTT), 23(5):797–821, October 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-020-00603-

x. See correction [835].

Bertrand:2021:CVR

[835] Nathalie Bertrand, Igor Konnov, and Josef Widder. Correction to: Verification of randomized consensus algorithms under round-rigid adversaries. International Journal on Software Tools for Technology Transfer (STTT), 23(5):823, October 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-021-00612-

4. See [834].

Abdulla:2021:CIS

[836] Parosh Aziz Abdulla, Frédéric Haziza, and Ahmed Rezine. Correction to: An integrated specification and verification technique for highly concurrent data structures. International Journal on Software Tools for Technology Transfer (STTT), 23(5):825, October 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-021-00629-

9. See [662].

Quesel:2021:CHM

[837] Jan-David Quesel, Stefan Mitsch, and André Platzer. Correction to: How to model and prove hybrid systems with KeYmaera: a tutorial on safety. International Journal on Software Tools for Technology Transfer (STTT), 23(5):827, October 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-021-00643-

x. See [595].

Beyer:2021:TCS

[838] Dirk Beyer and Marieke Huisman. TOOLympics I: Competition on software testing. International Journal on Software Tools for Technology Transfer (STTT), 23(6):829–832, December 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-021-00611-
                                        5.

Beyer:2021:FIC

[839] Dirk Beyer. First international competition on software testing. International Journal on Software Tools for Technology Transfer (STTT), 23(6): 833–846, December 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s10009-021-
                                        00613-3.

Jakobs:2021:CIV

[840] Marie-Christine Jakobs. CoVeriTest: interleaving value and predicate analysis for test-case generation. International Journal on Software Tools for Technology Transfer (STTT), 23(6): 847–851, December 2021. CODEN ????

ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10009-020-
00572-1.

Ruland:2021:CTM

[841] Sebastian Ruland, Malte Lochau, and Andy Schürr. CPA/Tiger-MGP: testgoal set partitioning for efficient multigoal test-suite generation. International Journal on Software Tools for Technology Transfer (STTT), 23(6):853– 856, December 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10009-020-
00574-z.

Gadelha:2021:EAT

[842] Mikhail R. Gadelha, Rafael S. Menezes, and Lucas C. Cordeiro. ESBMC 6.1: automated test case generation using bounded model checking. International Journal on Software Tools for Technology Transfer (STTT), 23(6):857– 861, December 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10009-020-
00571-2.

Lemieux:2021:FTF

[843] Caroline Lemieux and Koushik Sen.

FairFuzz-TC: a fuzzer targeting rare branches. International Journal on Software Tools for Technology Transfer (STTT), 23(6):863–866, December 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-020-00569-
w.

Cadar:2021:KSE

[844] Cristian Cadar and Martin Nowack. KLEE symbolic execution engine in 2019. International Journal on Software Tools for Technology Transfer (STTT), 23(6):867–870, December 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-020-00570-
                                        3.

Lemberger:2021:PRT

[845] Thomas Lemberger. Plain random test generation with PRTest. International Journal on Software Tools for Technology Transfer (STTT), 23(6):871– 873, December 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s10009-020-
                                        00568-x.

Chalupa:2021:SGT

[846] Marek Chalupa, Martina Vitovská, and Jan Strejcek. Symbiotic 6: generating test cases by slicing and symbolic execution. International Journal on Software Tools for Technology Transfer (STTT), 23(6):875–877, December 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-020-00573-
                                        0.

Beyer:2021:TIC

[847] Dirk Beyer, Marieke Huisman, and Bernhard Steffen. TOOLympics II: competitions on formal methods. International Journal on Software Tools for Technology Transfer (STTT), 23(6): 879–881, December 2021. CODEN ????

ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10009-021-
00631-1.

Dross:2021:VPV

[848] Claire Dross, Carlo A. Furia, and Peter Müller. VerifyThis 2019: a program verification competition. International Journal on Software Tools for Technology Transfer (STTT), 23(6): 883–893, December 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10009-021-
00619-x.

Sighireanu:2021:SCC

[849] Mihaela Sighireanu. SL-COMP: competition of solvers for separation logic. International Journal on Software Tools for Technology Transfer (STTT), 23(6):895–903, December 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-021-00628-
w.

Middeldorp:2021:CRE

[850] Aart Middeldorp, Julian Nagele, and Kiraku Shintani. CoCo 2019: report on the eighth confluence competition. International Journal on Software Tools for Technology Transfer (STTT), 23(6):905–916, December 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-021-00620-

Howar:2021:RCT

[851] Falk Howar, Marc Jasper, and Bernhard Steffen. The RERS challenge: towards controllable and scalable benchmark synthesis. International Journal on Software Tools for Technology Transfer (STTT), 23(6):917–930, December 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-021-00617-
                                        z.

Kordon:2021:SEM

[852] Fabrice Kordon, Lom Messan Hillah, and Emmanuel Paviot-Adet. Study of the efficiency of model checking techniques using results of the MCC from 2015 to 2019. International Journal on Software Tools for Technology Transfer (STTT), 23(6):931–952, December 2021. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-021-00615-
                                        1.

Charvat:2022:UPS

[853] Lukás Charvát, Ales Smrcka, and Tomás Vojnar. Utilizing parametric systems for detection of pipeline hazards. International Journal on Software Tools for Technology Transfer (STTT), 24(1):1–28, February 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-020-00591-
                                        y.

Vojnar:2022:TAC

[854] Tomás Vojnar and Lijun Zhang. Tools and algorithms for the construction and analysis of systems: a special issue for TACAS 2019. International Journal on Software Tools for Technology Transfer (STTT), 24(1):29–31, February 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-021-00642-
y.

Stoilkovska:2022:VSS

[855] Ilina Stoilkovska, Igor Konnov, and Florian Zuleger. Verifying safety of synchronous fault-tolerant algorithms by bounded model checking. International Journal on Software Tools for Technology Transfer (STTT), 24(1): 33–48, February 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10009-021-
00637-9.

Enevoldsen:2022:EAD

[856] Søren Enevoldsen, Kim Guldstrand Larsen, and Jirí Srba. Extended abstract dependency graphs. International Journal on Software Tools for Technology Transfer (STTT), 24(1):49–65, February 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-021-00638-
8.

Vukmirovic:2022:EBP

[857] Petar Vukmirović, Jasmin Blanchette, and Stephan Schulz. Extending a brainiac prover to lambda-free higherorder logic. International Journal on Software Tools for Technology Transfer (STTT), 24(1):67–87, February 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-021-00639-
                                        7.

Babar:2022:CBD

[858] Junaid Babar, Gianfranco Ciardo, and Andrew Miner. CESRBDDs: binary decision diagrams with complemented edges and edge-specified reductions. International Journal on Software Tools for Technology Transfer (STTT), 24(1):89–109, February 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-021-00640-
                                        0.

Blicha:2022:ULA

[859] Martin Blicha, Antti E. J. Hyvärinen, and Natasha Sharygina. Using linear algebra in decomposition of Farkas interpolants. International Journal on Software Tools for Technology Transfer (STTT), 24(1):111–125, February 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-021-00641-
                                        z.

Vernotte:2022:DSL

[860] Alexandre Vernotte, Aymeric Cretin, and Fabien Peureux. A domainspecific language to design false data injection tests for air traffic control systems. International Journal on Software Tools for Technology Transfer (STTT), 24(2):127–158, April 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-021-00604-
                                        4.

El-Hokayem:2022:BRV

[861] Antoine El-Hokayem and Yliès Falcone.

Bringing runtime verification home: a case study on the hierarchical monitoring of smart homes using decentralized specifications. International Journal on Software Tools for Technology Transfer (STTT), 24(2):159–181, April 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-021-00614-
2.

Dumas:2022:MLP

[862] Marlon Dumas, Luciano García-Bañuelos, and Maksym Yerokhin. Multilevel privacy analysis of business processes: the Pleak toolset. International Journal on Software Tools for Technology Transfer (STTT), 24(2):183–203, April 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-021-00636-
w.

Arras:2022:SLT

[863] Paul-Antoine Arras, Anastasios Andronidis, and Cristian Cadar. SaBRe: load-time selective binary rewriting. International Journal on Software Tools for Technology Transfer (STTT), 24(2): 205–223, April 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10009-021-
00644-w.

Louadah:2022:ICD

[864] Hassna Louadah and Yvan Labiche.

Interface control document modeling with Citrus (avionics systems interfaces). International Journal on Software Tools for Technology Transfer (STTT), 24(2):225–245, April 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-022-00645-
3.

Bettini:2022:SSM

[865] Lorenzo Bettini, Davide Di Ruscio, and Alfonso Pierantonio. Supporting safe metamodel evolution with edelta. International Journal on Software Tools for Technology Transfer (STTT), 24(2):247–260, April 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-022-00646-
2.

Strandberg:2022:STR

[866] Per Erik Strandberg, Wasif Afzal, and Daniel Sundmark. Software test results exploration and visualization with continuous integration and nightly testing. International Journal on Software Tools for Technology Transfer (STTT), 24(2):261–285, April 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-022-00647-
1.

Lowe:2022:PVS

[867] Gavin Lowe. Parameterized verification of systems with component identities, using view abstraction. International Journal on Software Tools for Technology Transfer (STTT), 24(2):287–324, April 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-022-00648-
0.

terBeek:2022:FMT

[868] Maurice H. ter Beek, Kim G. Larsen, and Tim A. C. Willemse. Formal methods and tools for industrial critical systems. International Journal on Software Tools for Technology Transfer (STTT), 24(3):325–330, June 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-022-00660-
4.

Weik:2022:DMA

[869] Norman Weik, Matthias Volk, and Nils Nießen. DFT modeling approach for operational risk assessment of railway infrastructure. International Journal on Software Tools for Technology Transfer (STTT), 24(3):331–350, June 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-022-00652-
4.

Basile:2022:EEE

[870] Davide Basile, Maurice H. ter Beek, and Axel Legay. Exploring the ERTMS/ETCS full moving block specification: an experience with formal methods. International Journal on Software Tools for Technology Transfer (STTT), 24(3):351–370, June 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-022-00653-
3.

Kouskoulas:2022:EWS

[871] Yanni Kouskoulas, T. J. Machado, and Joshua Brulé. Envelopes and waves: safe multivehicle collision avoidance for horizontal non-deterministic turns. International Journal on Software Tools for Technology Transfer (STTT), 24 (3):371–394, June 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s10009-022-
                                        00654-2.

Gu:2022:VSS

[872] Rong Gu, Peter G. Jensen, and Kristina Lundqvist. Verifiable strategy synthesis for multiple autonomous agents: a scalable approach. International Journal on Software Tools for Technology Transfer (STTT), 24(3):395–414, June 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-022-00657-
                                        z.

Binder:2022:FMV

[873] Benjamin Binder, Mihail Asavoae, and Mathieu Jan. Formal modeling and verification for amplification timing anomalies in the superscalar Tri-Core architecture. International Journal on Software Tools for Technology Transfer (STTT), 24(3):415–440, June 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-022-00655-
                                        1.

Bunte:2022:FVO

[874] Olav Bunte, Louis C. M. van Gool, and Tim A. C. Willemse. Formal verification of OIL component specifications using mCRL2. International Journal on Software Tools for Technology Transfer (STTT), 24(3):441–472, June 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-022-00658-
y.

Huang:2022:TLQ

[875] Samuel Huang and Rance Cleaveland.

Temporal-logic query checking over finite data streams. International Journal on Software Tools for Technology Transfer (STTT), 24(3):473–492, June 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-022-00656-
0.

Rockai:2022:DIS

[876] Petr Rockai and Jirí Barnat. DivSIM, an interactive simulator for LLVM bitcode. International Journal on Software Tools for Technology Transfer (STTT), 24(3):493–510, June 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-022-00659-
x.

Habibi:2022:GTW

[877] Elahe Habibi and Seyed-Hasan Mirian-Hosseinabadi. Generating test as a web service (TaaWS) through a methodbased attribute grammar. International Journal on Software Tools for Technology Transfer (STTT), 24(4):511–527, August 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-022-00649-
                                        z.

Salva:2022:LBM

[878] Sebastien Salva and Elliott Blot. Learning of behavioural models and dependency graphs for communicating systems with CkTailv2. International Journal on Software Tools for Technology Transfer (STTT), 24(4):529–548, August 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-022-00651-
                                        5.

Verbeek:2022:LSV

[879] H. M. W. Verbeek. The log skeleton visualizer in ProM 6.9. International Journal on Software Tools for Technology Transfer (STTT), 24(4):549–561, August 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-021-00618-
                                        y.

Back:2022:DAE

[880] Christoffer Olling Back, Tijs Slaats, Thomas Troels Hildebrandt, and Morten Marquard. DisCoveR: accurate and efficient discovery of declarative process models. International Journal on Software Tools for Technology Transfer (STTT), 24(4):563–587, August 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-021-00616-
                                        0.

Hensel:2022:PMC

[881] Christian Hensel, Sebastian Junges, Joost-Pieter Katoen, Tim Quatmann, and Matthias Volk. The probabilistic model checker Storm. International Journal on Software Tools for Technology Transfer (STTT), 24(4):589–610, August 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-021-00633-
z.

Legay:2022:TAC

[882] Axel Legay and Tiziana Margaria. Tools and algorithms for the construction and analysis of systems: a special issue for TACAS 2017. International Journal on Software Tools for Technology Transfer (STTT), 24(4):611–612, August 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-022-00664-
0.

Antonino:2022:AVC

[883] Pedro Antonino, Thomas Gibson-Robinson, and A. W. Roscoe. Approximate verification of concurrent systems using token structures and invariants. International Journal on Software Tools for Technology Transfer (STTT), 24(4):613–633, August 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-022-00650-
6.

Esparza:2022:LTL

[884] Javier Esparza, Jan Křetínský, Jean-François Raskin, and Salomon Sickert.

From linear temporal logic and limitdeterministic Büchi automata to deterministic parity automata. International Journal on Software Tools for Technology Transfer (STTT), 24(4):635–659, August 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-022-00663-
1.

Beyer:2022:CRB

[885] Dirk Beyer, Stefan Löwe, and Philipp Wendler. Correction to: Reliable benchmarking: requirements and solutions. International Journal on Software Tools for Technology Transfer (STTT), 24 (4):661, August 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10009-020-

00602-y. See [714].

Biere:2022:TAC

[886] Armin Biere and David Parker. Tools and algorithms for the construction and analysis of systems: a special issue for TACAS 2020. International Journal on Software Tools for Technology Transfer (STTT), 24(5):663–665, October 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-022-00677-
9.

Frenkel:2022:AGR

[887] Hadar Frenkel, Orna Grumberg, Corina S. Păsăreanu, and Sarai Sheinvald. Assume, guarantee or repair: a regular framework for non regular properties. International Journal on Software Tools for Technology Transfer (STTT), 24(5):667–689, October 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-022-00669-
9.

Frohn:2022:CML

[888] Florian Frohn and Carsten Fuhs. A calculus for modular loop acceleration and non-termination proofs. International Journal on Software Tools for Technology Transfer (STTT), 24(5):691–715, October 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-022-00670-
2.

Becker:2022:CPA

[889] Benedikt Becker, Nicolas Jeannerod, Claude Marché, Yann Régis-Gianas, Mihaela Sighireanu, and Ralf Treinen. The CoLiS platform for the analysis of maintainer scripts in Debian software packages. International Journal on Software Tools for Technology Transfer (STTT), 24(5):717–733, October 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-022-00671-
1.

Neele:2022:POR

[890] Thomas Neele, Tim A. C. Willemse, Wieger Wesselink, and Antti Valmari. Partial-order reduction for parity games and parameterised Boolean equation systems. International Journal on Software Tools for Technology Transfer (STTT), 24(5):735–756, October 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/

                                        article/10.1007/s10009-022-00672-
                                        0.

Hamers:2022:DPR

[891] Ruben Hamers, Erik Horlings, and Sung-Shik Jongmans. The Discourje project: run-time verification of communication protocols in Clojure. International Journal on Software Tools for Technology Transfer (STTT), 24(5):757–782, October 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-022-00674-
                                        y.

Hiep:2022:VOL

[892] Hans-Dieter A. Hiep, Olaf Maathuis, Jinting Bian, Frank S. de Boer, and Stijn de Gouw. Verifying Open-JDK’s LinkedList using KeY (extended paper). International Journal on Software Tools for Technology Transfer (STTT), 24(5):783–802, October 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-022-00679-
                                        7.

Badings:2022:SBV

[893] Thom Badings, Murat Cubuktepe, Nils Jansen, Sebastian Junges, Joost-Pieter Katoen, and Ufuk Topcu. Scenariobased verification of uncertain parametric MDPs. International Journal on Software Tools for Technology Transfer (STTT), 24(5):803–819, October 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-022-00673-
                                        z.

Budde:2022:ANM

[894] Carlos E. Budde, Pedro R. D’Argenio, Raúl E. Monti, and Mariëlle Stoelinga. Analysis of non-Markovian repairable fault trees through rare event simulation. International Journal on Software Tools for Technology Transfer (STTT), 24(5):821–841, October 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-022-00675-
x.

Chakraborty:2022:FPI

[895] Supratik Chakraborty, Ashutosh Gupta, and Divyesh Unadkat. Full-program induction: verifying array programs sans loop invariants. International Journal on Software Tools for Technology Transfer (STTT), 24(5):843–888, October 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-022-00676-
w.

Mohammed:2022:GTA

[896] Mawal A. Mohammed, Jameleddine Hassine, and Mohammad Alshayeb. GS-Detector: a tool for automatic detection of bad smells in GRL goal models. International Journal on Software Tools for Technology Transfer (STTT), 24(6):889–910, December 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-022-00662-
2.

Lee:2022:MFA

[897] Jaehun Lee, Kyungmin Bae, Peter Csaba Ölveczky, Sharon Kim, and Minseok Kang. Modeling and formal analysis of virtually synchronous cyber-physical systems in AADL. International Journal on Software Tools for Technology Transfer (STTT), 24(6): 911–948, December 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10009-022-
00665-z.

Taylor:2022:AFV

[898] Ramsay G. Taylor, Michael Foster, and Siobhán North. An automated framework for verifying or refuting trace properties of extended finite state machines. International Journal on Software Tools for Technology Transfer (STTT), 24(6):949–972, December 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-022-00666-
y.

Lafuente:2022:FMT

[899] Alberto Lluch Lafuente and Anastasia Mavridou. Formal methods and tools for industrial critical systems. International Journal on Software Tools for Technology Transfer (STTT), 24(6):973–976, December 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-022-00687-
7.

Lourenco:2022:AFA

[900] Cláudio Belo Lourenço, Denis Cousineau, Florian Faissole, Claude Marché, David Mentré, and Hiroaki Inoue. Automated formal analysis of temporal properties of Ladder programs. International Journal on Software Tools for Technology Transfer (STTT), 24(6):977–997, December 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-022-00680-
0.

Hansen:2022:VSC

[901] Simon Thrane Hansen, Casper Thule, Cláudio Gomes, Jaco van de Pol, Maurizio Palmieri, Emin Oguz Inci, Frederik Madsen, Jesús Alfonso, José Ángel Castellanos, and José Manuel Rodriguez. Verification and synthesis of co-simulation algorithms subject to algebraic loops and adaptive steps. International Journal on Software Tools for Technology Transfer (STTT), 24(6):999–1024, December 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-022-00686-
8.

Kiviriga:2022:RRA

[902] Andrej Kiviriga, Kim Guldstrand Larsen, and Ulrik Nyman. Randomized reachability analysis in UP-PAAL: fast error detection in timed systems. International Journal on Software Tools for Technology Transfer (STTT), 24(6):1025–1042, December 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-022-00681-
z.

Schmidt:2022:SSV

[903] Joshua Schmidt and Michael Leuschel.

SMT solving for the validation of B and Event-B models. International Journal on Software Tools for Technology Transfer (STTT), 24(6):1043–1077, December 2022. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-022-00682-
                                        y.

Berti:2023:OPA

[904] Alessandro Berti and Wil M. P. van der Aalst. OC-PM: analyzing object-centric event logs and process models. International Journal on Software Tools for Technology Transfer (STTT), 25(1):1–17, February 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-022-00668-
                                        w.

Lehmann:2023:BDB

[905] Sascha Lehmann and Sibylle Schupp. Bounded DBM-based clock state construction for timed automata in Uppaal. International Journal on Software Tools for Technology Transfer (STTT), 25(1):19–47, February 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-022-00667-
                                        x.

DeSanctis:2023:TTJ

[906] Martina De Sanctis, Amleto Di Salle, Ludovico Iovino, and Maria Teresa Rossi. A technology transfer journey to a model-driven access control system. International Journal on Software Tools for Technology Transfer (STTT), 25(1):49–74, February 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-023-00697-
z.

Laarman:2023:ISI

[907] Alfons Laarman and Ana Sokolova.

Introduction to the special issue for SPIN 2021. International Journal on Software Tools for Technology Transfer (STTT), 25(1):75–76, February 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-023-00698-
y.

Kirszenberg:2023:GFL

[908] Alexandre Kirszenberg, Antoine Martin, Hugo Moreau, and Etienne Renault. Go2Pins: a framework for the LTL verification of Go programs (extended version). International Journal on Software Tools for Technology Transfer (STTT), 25(1):77–94, February 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-022-00692-
w.

Amat:2023:LPR

[909] Nicolas Amat, Silvano Dal Zilio, and Didier Le Botlan. Leveraging polyhedral reductions for solving Petri net reachability problems. International Journal on Software Tools for Technology Transfer (STTT), 25(1):95–114, February 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-022-00694-

Nxumalo:2023:EAM

[910] Madoda Nxumalo, Nils Timm, and Stefan Gruner. An evaluation of approaches to model checking real-time task schedulability analysis. International Journal on Software Tools for Technology Transfer (STTT), 25(1):115–128, February 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-022-00693-
                                        9.

Jensen:2023:TAC

[911] Peter Gjøl Jensen and Thomas Neele. Tools and algorithms for the construction and analysis of systems: a special issue on tool papers for TACAS 2021. International Journal on Software Tools for Technology Transfer (STTT), 25(2):129–131, April 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-023-00699-
                                        x.

Kaufmann:2023:IAV

[912] Daniela Kaufmann and Armin Biere. Improving AMulet2 for verifying multiplier circuits using SAT solving and computer algebra. International Journal on Software Tools for Technology Transfer (STTT), 25(2):133–144, April 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-022-00688-
                                        6.

Sotoudeh:2023:STA

[913] Matthew Sotoudeh, Zhe Tao, and Aditya V. Thakur. SyReNN: a tool for analyzing deep neural networks. International Journal on Software Tools for Technology Transfer (STTT), 25(2): 145–165, April 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10009-023-
00695-1.

Tan:2023:VPR

[914] Yong Kiam Tan, Marijn J. H. Heule, and Magnus O. Myreen. Verified propagation redundancy and compositional UN-SAT checking in CakeML. International Journal on Software Tools for Technology Transfer (STTT), 25(2):167–184, April 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-022-00690-
y.

Abbasi:2023:CRS

[915] Rosa Abbasi, Jonas Schiffl, Eva Darulova, Mattias Ulbrich, and Wolfgang Ahrendt. Combining ruleand SMT-based reasoning for verifying floating-point Java programs in KeY. International Journal on Software Tools for Technology Transfer (STTT), 25(2):185–204, April 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-022-00691-
x.

Biewer:2023:RR

[916] Sebastian Biewer, Bernd Finkbeiner, Holger Hermanns, Maximilian A. Köhl, Yannik Schnitzer, and Maximilian Schwenger. On the road with RTLola. International Journal on Software Tools for Technology Transfer (STTT), 25(2):205–218, April 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-022-00689-
                                        5.

Scott:2023:ASS

[917] Joseph Scott, Aina Niemetz, Mathias Preiner, Saeed Nejati, and Vijay Ganesh. Algorithm selection for SMT. International Journal on Software Tools for Technology Transfer (STTT), 25(2):219–239, April 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-023-00696-

0. See publisher correction [952].

Jansen:2023:EPL

[918] Nils Jansen, Gerrit Nolte, and Bernhard Steffen. Explanation Paradigms Leveraging Analytic Intuition (ExPLAIn). International Journal on Software Tools for Technology Transfer (STTT), 25 (3):241–247, June 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s10009-023-
                                        00715-0.

Jungermann:2023:AEC

[919] Florian Jüngermann, Jan Kretínský, and Maximilian Weininger. Algebraically explainable controllers: decision trees and support vector machines join forces. International Journal on Software Tools for Technology Transfer (STTT), 25(3):249–266, June 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/

article/10.1007/s10009-023-00716-
z.

Gossen:2023:AAR

[920] Frederik Gossen and Bernhard Steffen. Algebraic aggregation of random forests: towards explainability and rapid evaluation. International Journal on Software Tools for Technology Transfer (STTT), 25(3):267–285, June 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-021-00635-
x.

Murtovi:2023:FGT

[921] Alnis Murtovi, Alexander Bainczyk, Gerrit Nolte, Maximilian Schlüter, and Bernhard Steffen. Forest GUMP: a tool for verification and explanation. International Journal on Software Tools for Technology Transfer (STTT), 25 (3):287–299, June 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10009-023-
00702-5.

Schluter:2023:TRU

[922] Maximilian Schlüter, Gerrit Nolte, Alnis Murtovi, and Bernhard Steffen. Towards rigorous understanding of neural networks via semantics-preserving transformations. International Journal on Software Tools for Technology Transfer (STTT), 25(3):301–327, June 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-023-00700-

Brix:2023:FTY

[923] Christopher Brix, Mark Niklas Müller, Stanley Bak, Taylor T. Johnson, and Changliu Liu. First three years of the international verification of neural networks competition (VNN-COMP). International Journal on Software Tools for Technology Transfer (STTT), 25 (3):329–339, June 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10009-023-
00703-4.

Khmelnitsky:2023:ARN

[924] Igor Khmelnitsky, Daniel Neider, Rajarshi Roy, Xuan Xie, Benoît Barbot, Benedikt Bollig, Alain Finkel, Serge Haddad, Martin Leucker, and Lina Ye. Analysis of recurrent neural networks via property-directed verification of surrogate models. International Journal on Software Tools for Technology Transfer (STTT), 25(3):341–354, June 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-022-00684-
w.

Nolte:2023:PTA

[925] Gerrit Nolte, Maximilian Schlüter, Alnis Murtovi, and Bernhard Steffen. The power of typed affine decision structures: a case study. International Journal on Software Tools for Technology Transfer (STTT), 25(3):355–374, June 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-023-00701-
6.

Badings:2023:DMU

[926] Thom Badings, Thiago D. Simão, Marnix Suilen, and Nils Jansen. Decision-making under uncertainty: beyond probabilities. International Journal on Software Tools for Technology Transfer (STTT), 25(3):375–391, June 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-023-00704-
3.

Usman:2023:OSC

[927] Muhammad Usman, Youcheng Sun, Divya Gopinath, Rishi Dange, Luca Manolache, and Corina S. Păsăreanu. An overview of structural coverage metrics for testing neural networks. International Journal on Software Tools for Technology Transfer (STTT), 25 (3):393–405, June 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10009-022-
00683-x.

Gros:2023:ANN

[928] Timo P. Gros, Holger Hermanns, Jörg Hoffmann, Michaela Klauck, and Marcel Steinmetz. Analyzing neural network behavior through deep statistical model checking. International Journal on Software Tools for Technology Transfer (STTT), 25(3):407–426, June 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-022-00685-
9.

Deshmukh:2023:ISI

[929] Jyotirmoy Deshmukh and Dejan Nicković. Introduction to the special issue on runtime verification. International Journal on Software Tools for Technology Transfer (STTT), 25(4):427–429, August 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-023-00721-
2.

Ho:2023:PAU

[930] Vivian M. Ho, Chris Alvin, Jimmie D. Lawson, Supratik Mukhopadhyay, and Brian Peterson. Program analysis using empirical abstraction. International Journal on Software Tools for Technology Transfer (STTT), 25(4):431–452, August 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-023-00709-
y.

Soueidi:2023:EEB

[931] Chukri Soueidi, Marius Monnier, and Yliès Falcone. Efficient and expressive bytecode-level instrumentation for Java programs. International Journal on Software Tools for Technology Transfer (STTT), 25(4):453–479, August 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-023-00708-
z.

Shafiei:2023:CRV

[932] Nastaran Shafiei, Klaus Havelund, and Peter Mehlitz. Concurrent runtime verification of data rich events. International Journal on Software Tools for Technology Transfer (STTT), 25(4): 481–501, August 2023. CODEN ????

ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10009-023-
00713-2.

Nenzi:2023:MLT

[933] Laura Nenzi, Ezio Bartocci, Luca Bortolussi, Simone Silvetti, and Michele Loreti. MoonLight: a lightweight tool for monitoring spatio-temporal properties. International Journal on Software Tools for Technology Transfer (STTT), 25(4):503–517, August 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-023-00710-
5.

Feng:2023:ISI

[934] Lu Feng and Dana Fisman. Introduction to the special issue on runtime verification. International Journal on Software Tools for Technology Transfer (STTT), 25(4):519–520, August 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-023-00722-
1.

Fernando:2023:DDM

[935] Vimuth Fernando, Keyur Joshi, Jacob Laurel, and Sasa Misailovic. Diamont: dynamic monitoring of uncertainty for distributed asynchronous programs. International Journal on Software Tools for Technology Transfer (STTT), 25(4):521–539, August 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-023-00717-
y.

Momtaz:2023:PMD

[936] Anik Momtaz, Niraj Basnet, Houssam Abbas, and Borzoo Bonakdarpour. Predicate monitoring in distributed cyber-physical systems. International Journal on Software Tools for Technology Transfer (STTT), 25(4):541–556, August 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-023-00718-
                                        x.

Mamouras:2023:CFA

[937] Konstantinos Mamouras, Agnishom Chattopadhyay, and Zhifu Wang. A compositional framework for algebraic quantitative online monitoring over continuous-time signals. International Journal on Software Tools for Technology Transfer (STTT), 25(4):557–573, August 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-023-00719-
                                        w.

Kueffner:2023:UAM

[938] Konstantin Kueffner, Anna Lukina, Christian Schilling, and Thomas A. Henzinger. Into the unknown: active monitoring of neural networks (extended version). International Journal on Software Tools for Technology Transfer (STTT), 25(4):575–592, August 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-023-00711-
                                        4.

Baumeister:2023:MVG

[939] Jan Baumeister, Johann C. Dauer, Bernd Finkbeiner, and Sebastian Schirmer. Monitoring with verified guarantees. International Journal on Software Tools for Technology Transfer (STTT), 25(4):593–616, August 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-023-00712-
3.

Wirsing:2023:REC

[940] Martin Wirsing, Stefan Jähnichen, and Rocco De Nicola. Rigorous engineering of collective adaptive systems — 2nd special section. International Journal on Software Tools for Technology Transfer (STTT), 25(5–6):617–624, December 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-023-00734-
x.

Bozga:2023:CDC

[941] Marius Bozga and Joseph Sifakis. Correct by design coordination of autonomous driving systems. International Journal on Software Tools for Technology Transfer (STTT), 25(5–6): 625–639, December 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10009-023-
00723-0.

Basile:2023:TSS

[942] Davide Basile, Maurice H. ter Beek, Laura Bussi, and Vincenzo Ciancia. A toolchain for strategy synthesis with spatial properties. International Journal on Software Tools for Technology Transfer (STTT), 25(5–6):641–658, December 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-023-00730-
                                        1.

Fettke:2023:CTI

[943] Peter Fettke and Wolfgang Reisig. A causal, time-independent synchronization pattern for collective adaptive systems. International Journal on Software Tools for Technology Transfer (STTT), 25(5–6):659–673, December 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-023-00733-
                                        y.

DeNicola:2023:MFB

[944] Rocco De Nicola, Luca Di Stefano, Omar Inverso, and Serenella Valiani. Modelling flocks of birds and colonies of ants from the bottom up. International Journal on Software Tools for Technology Transfer (STTT), 25(5–6):675–691, December 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-023-00731-
                                        0.

Monica:2023:KAI

[945] Stefania Monica, Federico Bergenti, and Franco Zambonelli. A kinetic approach to investigate the collective dynamics of multi-agent systems. International Journal on Software Tools for Technology Transfer (STTT), 25(5–6):693–705, December 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/

article/10.1007/s10009-023-00724-
z.

Yifeng:2023:MAC

[946] Chen Yifeng and J. W. Sanders. A modal approach to conscious social agents. International Journal on Software Tools for Technology Transfer (STTT), 25(5–6):707–716, December 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-023-00732-
z.

Topfer:2023:MLA

[947] Michal Töpfer, Milad Abdullah, Tomáš Bureš, Petr Hnětynka, and Martin Kruliš. Machine-learning abstractions for component-based self-optimizing systems. International Journal on Software Tools for Technology Transfer (STTT), 25(5–6):717–731, December 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-023-00726-
x.

Bures:2023:GAR

[948] Tomáš Bureš, Petr Hnětynka, Martin Kruliš, František Plášil, Danylo Khalyeyev, Sebastian Hahner, Stephan Seifermann, Maximilian Walter, and Robert Heinrich. Generating adaptation rule-specific neural networks. International Journal on Software Tools for Technology Transfer (STTT), 25(5–6): 733–746, December 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10009-023-
00725-y.

Bettini:2023:CPM

[949] Lorenzo Bettini, Khalid Bourr, Rosario Pugliese, and Francesco Tiezzi. Coordinating and programming multiple ROSbased robots with X-KLAIM. International Journal on Software Tools for Technology Transfer (STTT), 25(5–6): 747–764, December 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s10009-023-
                                        00727-w.

Alrahman:2023:LSV

[950] Yehia Abd Alrahman, Shaun Azzopardi, Luca Di Stefano, and Nir Piterman. Language support for verifying reconfigurable interacting systems. International Journal on Software Tools for Technology Transfer (STTT), 25(5–6):765–784, December 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-023-00729-
                                        8.

Murgia:2023:CPA

[951] Maurizio Murgia, Riccardo Pinciroli, Catia Trubiani, and Emilio Tuosto. Comparing performance abstractions for collective adaptive systems. International Journal on Software Tools for Technology Transfer (STTT), 25(5–6):785–798, December 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-023-00728-
                                        9.

Scott:2023:PCA

[952] Joseph Scott, Aina Niemetz, Mathias Preiner, Saeed Nejati, and Vijay Ganesh. Publisher correction: Algorithm selection for SMT. International Journal on Software Tools for Technology Transfer (STTT), 25(5–6):799– 800, December 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10009-023-

00714-1. See [917].

Angulo:2024:PCK

[953] Guisella Angulo, Daniel San Martín, Fabiano Ferrari, Ignacio García-Rodríguez de Guzmán, Ricardo Perez-Castillo, and Valter Vieira de Camargo. A process for creating KDM2PSM transformation engines. International Journal on Software Tools for Technology Transfer (STTT), 26(1):1–20, February 2024. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-024-00735-
4.

Kosinska:2024:KRS

[954] Joanna Kosińska, Grzegorz Brotoń, and Maciej Tobiasz. Knowledge representation of the state of a cloudnative application. International Journal on Software Tools for Technology Transfer (STTT), 26(1):21–32, February 2024. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-023-00705-
2.

Havelund:2024:PEM

[955] Klaus Havelund and Gerard J. Holzmann. Programming event monitors. International Journal on Software Tools for Technology Transfer (STTT), 26(1):33–47, February 2024. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-023-00706-
                                        1.

Schupp:2024:AHS

[956] Stefan Schupp, Erika Ábrahám, Md Tawhid Bin Waez, Thomas Rambow, and Zeng Qiu. On the applicability of hybrid systems safety verification tools from the automotive perspective. International Journal on Software Tools for Technology Transfer (STTT), 26(1):49–78, February 2024. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-023-00707-
                                        0.

Yamaguchi:2024:RRR

[957] Tomoya Yamaguchi, Bardh Hoxha, and Dejan Nicković. RTAMT — runtime robustness monitors with application to CPS and robotics. International Journal on Software Tools for Technology Transfer (STTT), 26(1):79–99, February 2024. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-023-00720-
                                        3.

Hendriks:2024:VTA

[958] Martijn Hendriks, Jacques Verriet, and Twan Basten. Visualization, transformation, and analysis of execution traces with the eclipse TRACE4CPS trace tool. International Journal on Software Tools for Technology Transfer (STTT), 26(1):101–126, February 2024. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-024-00736-
3.

Groote:2024:FMI

[959] Jan Friso Groote and Marieke Huisman. Formal methods for industrial critical systems. International Journal on Software Tools for Technology Transfer (STTT), 26(2):127–129, April 2024. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-024-00744-
3.

Cassez:2024:DVS

[960] Franck Cassez, Joanne Fuller, and Horacio Mijail Antón Quiles. Deductive verification of smart contracts with Dafny. International Journal on Software Tools for Technology Transfer (STTT), 26(2):131–145, April 2024. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-024-00738-
1.

Vu:2024:GID

[961] Fabian Vu, Christopher Happe, and Michael Leuschel. Generating interactive documents for domain-specific validation of formal models. International Journal on Software Tools for Technology Transfer (STTT), 26(2):147–168, April 2024. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-024-00739-
0.

MatosPedro:2024:MST

[962] André Matos Pedro, Tomás Silva, Tiago Sequeira, João Lourenço, João Costa Seco, and Carla Ferreira. Monitoring of spatio-temporal properties with nonlinear SAT solvers. International Journal on Software Tools for Technology Transfer (STTT), 26(2):169–188, April 2024. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-024-00740-
                                        7.

Ducoffe:2024:CAS

[963] Mélanie Ducoffe, Christophe Gabreau, Ileana Ober, Iulian Ober, and Eric Guillaume Vidot. Certification of avionic software based on machine learning: the case for formal monotony analysis. International Journal on Software Tools for Technology Transfer (STTT), 26(2):189–205, April 2024. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-024-00741-
                                        6.

Hampus:2024:FVD

[964] Anton Hampus and Mattias Nyberg. Formally verifying decompositions of stochastic specifications. International Journal on Software Tools for Technology Transfer (STTT), 26(2):207–228, April 2024. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10009-024-00742-
                                        5.

Adelt:2024:RFM

[965] Julius Adelt, Julian Gebker, and Paula Herber. Reusable formal models for concurrency and communication in custom real-time operating systems. International Journal on Software Tools for Technology Transfer (STTT), 26(2):229–245, April 2024. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

URL https://link.springer.com/
article/10.1007/s10009-024-00743-